A Complete Bibliography of Proceedings of the ACM on Programming Languages (PACMPL)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 May 2023 Version 1.08

Title word cross-reference

1 [XTZ^{+}23]. ▷◁ [RHdSOZ23]. d [HL20]. ι [KLL20]. λ [FKR20, KyKS22, SC18]. λ_{s} [SMC21]. λω [VS20]. O(1) [AZMV23]. ω [AKK^{+}21]. Π [BCRA18]. Σ [BCRA18].

-calculi [KyKS22]. -calculus [FKR20]. -Minimality [XTZ^{+}23]. -regular [AKK^{+}21].

.NET [PVV^{+}17].

2KB [KSS20].

3C [MKM^{+}22]. 3CPS [QRS22].

72 [Ing20].

abductive [ZDDJ21]. ableC [KKCV17]. Abridging [YMJ17]. absorbing [LP20]. Abstract [BGMW20, BGG^{+}20, Lem23, BCD22, CDG22, Cra19, FSSW22, FS22, JTD21, KMD^{+}22, KyKS22, KPE18, KE19, KMP19, LYU^{+}22, MJ20, RMH21, SPV18, VPD19, WDR18, WCR19, ZYT^{+}19]. Abstracting [BPPS19, DLNV17, MM19, SCJG21]. Abstraction [SL23, ZM19, BCDG22, CKA18, CdV20, GJJ^{+}20, JDT22, LYSM22, LRS^{+}20, MV20, WDS18, YSHZ21]. Abstraction-safe [ZM19]. Abstractions [CKN^{+}23, FNB^{+}21, KSW18]. academia [TRWS22]. accelerate [CB17]. Accelerating [LS23, JSXH20]. access [BP19]. account [vdRS22]. accuracy [BCK^{+}21, FJM19, MSRH20, SSH^{+}19]. accuracy-aware [MSRH20, SSH^{+}19]. Achieving [HLK^{+}20]. acquire [AAJN18, DD22, YM21]. across [XCB^{+}18]. Actor [BFSK20, BBB^{+}17, CFD^{+}17, SEP^{+}20]. actor-based [BBB^{+}17]. Actris [HBK20]. ad [JME20, SV23]. adaptive [WCHRP17]. Adding [ABF20]. ADEV [LHSM23]. adjunctions [GHHW18]. Admissible [PT23a]. ADO [HKSS21]. advanced [CAMS20, GLTS22]. adverbs [LW22a]. Adversarial [YAY20, ABG^{+}21]. Aeneas [HP22]. affected [MNT20]. Affine [DKPS23, BKPS18, ST21]. after [MKV18]. again [NJA20]. against [ML21, VREV19]. Agda [VMA19]. age [BCG^{+}20]. Agnostic [LCT^{+}23, XTZ^{+}23]. Agreement [WJS23, JWJ^{+}21]. Agreement-Based [WJS23, JWJ^{+}21]. ahead [MV20]. ahead-of-time [MV20]. aided [JGW^{+}20, NWP^{+}18, RGSNT17, WLH^{+}17]. Aiming [HKGK20]. air [OD18]. AL [CR19]. alarms [ZGSN17]. Algebra [AKL^{+}23, BHP^{+}19, CKA18, CA22, GHHW18, KKC^{+}17, PC22, SHW^{+}20, SFH^{+}20, ZdAG22, ZH18]. Algebraic [PvdR23, Ahm18, BPPS18, BPPS19, BAS^{+}18, FNB^{+}21, Ham17, KSG22, KKPS21, MRX^{+}19, Mor19, NPWW22, Par20, STI20, WBTR20, YD22, ZN21]. Algebras [KW21, PvdR23, YvGK18]. Algebro [GHMM23]. Algebro-geometric [GHMM23]. algorithm [GRB20, MS19, XK21]. Algorithmic [LY18, LSC^{+}22]. Algorithms [FKP23, GHMM23, WBdG^{+}23, Ell17b, LSZ22, ÖN22, WNC^{+}19]. alias [CCP18b, CKP22, GM21, SAB17]. alias-aware [SAB17]. Aliasing [EBP^{+}23, BWB^{+}21, JDKD20]. Alignment [AKL^{+}23]. alive [MT17]. allocation [SdSCQ19]. almost [DFD21, HFCG19, MMKK18, SV21]. almost-sure [DFD21, HFCG19, MMKK18]. Alone [KNNJ18]. Alpaca [MCL17]. already [RAT17]. Ambiguity [PCHB23]. Amortized [BKK^{+}23, LVSZ23, CLD20, KH21, RGGH21]. analyses [SBEV18]. Analysis [FAM23, LRY23, ADN22, AD17, AMS20, ABD21, BGC20, CDG22, COHY17, CCP18b, CKP22, Cho22, CLD20, DHP18, GM21, GS17, HH20, JJO18, JLO20, JJCO17, KH21, LYSM22, LY18, LTMS18, LH22, LX19, MP21a, Mil20, QGG19, RBG^{+}18, RGGH21, RKS18, SWYZ22, SGL^{+}21, SAB17, SNCM19, TLM^{+}21, TCJ20, WWC17, WKH20, WCR19, YSHZ21, YHW20, ZGSN17, ZYT^{+}19, ZPZX21]. analytic [PVSK18]. Analytical [BKPS18]. analytics [HLV21]. analyzers [RH22]. Analyzing [QRS22, INN21, MH21, RH22, SSK17]. ancestors [FMG^{+}20]. anchor [FF20]. Andersen [MP21a]. Android [BFV18, SYW^{+}21]. angelic [MNB^{+}22]. AnICA [RH22]. annotations [CC20]. Answer [SU23]. Answer-Effect [SU23]. Anti [CKN^{+}23]. Anti-unification [CKN^{+}23]. any [MM19, PXA^{+}21]. AnyDSL [LBH^{+}18]. AOT [Ser21]. API [BFSK20, FSD21, JGW^{+}20, YNIC19]. APIfix [GRS^{+}21]. APIs [CHJ^{+}19, GTT21, RG23]. APL [HK20]. app [FZSN20]. Application [LRY23, BGC18, CNRG19, GHMM20, JLO20, JR21, SBM20, WSH^{+}19, ZR21]. application-specific [BGC18, JLO20]. Applications [KZA^{+}23, AMT17, AFH^{+}18, GAGG^{+}18, Ham17, LMM17a, MDSM19, OBL^{+}18, WDLC18]. applicative [MLMD19]. Applying [DNFK22, BSZL^{+}18]. Approach [ADH^{+}23, LL23, SSS^{+}23, SL23, ACN18, BBRM17, CR20, HH20, HBP19, LAGN21, PES20, PS18, WR18, WCH^{+}19]. Approaches [BGC23]. ApproxHPVM [SSH^{+}19]. Approximate [ETG19, FJM19, SFRM20, USM22]. approximation [DR19, LN22, WAPJ22]. approximations [MPV18, MMS^{+}22]. apps [SYW^{+}21]. Architecture [FSX^{+}23, JPR22, RWNV20, Rei17]. architectures [DJR18]. argument [RAJ^{+}17]. Arithmetic [KKZ23, FK18, PUW^{+}21, SMC19, TML^{+}22]. ARM [PFD^{+}18, Rei17]. Arm [FSX^{+}23]. ARMv8 [PFD^{+}18, RWV19]. Aroma [LYB^{+}19]. array [HHY^{+}21, OCDR17, PJD^{+}21, SFVJ19]. array-processing [SFVJ19]. arrays [NcS17]. art [HP19]. ask [RvAP^{+}20]. ASP [BGC23]. aspects [Cho22]. Asphalion [VREV19]. assembly [CGL22, SCK^{+}20]. Assertion [HHT20]. Assertion-based [HHT20]. assertions [LZY^{+}20]. assessing [TBB^{+}19]. assessment [ACM^{+}18]. assets [CAMS20, HVH20]. assignments [LSSO18, SLO19]. assisted [CRMA20, SVR17]. association [SZD^{+}17]. assumptions [CTW21]. asymptotic [HLL20]. asymptotically [KJK^{+}22]. Asynchronous [AP21, AZMT18, CY19, FLMD19, GPP^{+}21, GTT21, GMOO21]. atomic [PFD^{+}18, ZZX^{+}20]. Attacks [SADC23]. attention [LWNN19]. attention-based [LWNN19]. auditing [ZPG^{+}17]. Augmented [ASMS20, DAY20]. autodiff [PJD^{+}21]. autogenerating [CR19]. Automata [DTSLT23, GMOO21, LKM23, AM17, THL^{+}20, WDS17]. Automated [DELT17, KTST23, MCF^{+}22, PWSD20, TLKC18, ZMSD23, ACM^{+}18, CTR^{+}20, HP19, SYW^{+}21, ZSL^{+}22]. Automatic [DPM23, GTT21, KH21, LSSO18, LHSM23, MP21b, PHXD19, SC20, SLO19, AAM^{+}21, Ell18, Ell21, KKCV17, KJK^{+}22, LYU^{+}22, PXA^{+}21]. Automatically [COHY17, KKSL22, VDvG^{+}21, BSPC19, BBZ18, DRS21, LSA18]. Automating [AD17, KLV23, Jab20, WKH20].

Automation [MK23]. AutoPandas

[BLF^{+}19]. avoidance [CSM^{+}17]. avoiding [OD18, PBC^{+}21]. Aware [DWH23, LVSZ23, CPY20, CEEW21, FG18, FBG20, GRXB19, HKVR21, HKSS21, MSRH20, NSGH22, OMO19, PWZS21, PXA^{+}21, SKS^{+}19, SC20, SSH^{+}19, SZZ^{+}19, SHTZ^{+}19, SAB17, WZS20]. axiomatic [KH18, PFD^{+}18]. axiomatizations [KLMM22].

babble [CKN^{+}23]. Back

[KCB^{+}23, MSBO23, BSLBG22, JDT22]. back-translation [JDT22]. backed [BLF^{+}19]. Backend [BBP23]. Backpropagation [BMP20]. backpropagator [WZD^{+}19]. backward [KZK^{+}18]. Based [ADH^{+}23, BGC23, GHMM23, KO23, WJS23, ASMS20, BKN19, BKKM21, BBB^{+}17, BSLBG22, CNRG19, CB18, CKP22, ESCL22, FDvdHP22, HHT20, HBK20, JWJ^{+}21, JBK22a, JLO20, JKT21, KNNR17, LHP19, LLSS22, LWNN19, LZY^{+}20, MKV18, Mil20, MJ22, MKLR20, MDBL20, NGR^{+}22, PCAGG21, PS18, RRG^{+}21, RFH^{+}22, SHdM20, SM20, SCZW20, SBEV18, VMS^{+}22, WMM20, WBS^{+}22, BGC20]. basis [KH18]. batch [SLDN17]. Bayesian [DKPS23, ŚKV^{+}18, ŚKG18]. BDA [ZYT^{+}19]. be [CAMS20, GJK^{+}21, JO22]. Beast [RHdSOZ23]. before [MKV18]. behavior [BKPS18]. behind [SHdM20]. beliefs [KNNR17]. benefits [SGDL20]. bent [RST20]. Berry [BM20]. Better [CKN^{+}23, KMP19, Mai17, SV21]. between [LMÖ^{+}22, WDR18]. Beyond [Ram22]. BFF [ZSL^{+}22]. bias [WCBG17]. bidirected [KP22, LZR21, LSZ22]. Bidirectional [MKC18, ZGHH23, DK19, KH18, LCOC20, YMDW21, ZSM20]. Big [ZMSD23, LRV^{+}22]. bijections [YF18]. bijective [MFP^{+}18]. binaries [DAY20]. binary [CP22, DSPC^{+}19, MOP21, ZYT^{+}19]. binder [Chl21]. Binders [BPPS20, CR20]. binding [AAC^{+}18, QRS22]. Bindings [BGPT19]. bioinformatics [SNB^{+}19]. BioScript [OLC^{+}18]. Bisimilarity [JW23]. Bisimulation [MV19]. Bit [LKM23]. bitfield [ZSL^{+}22]. bitfield-manipulating [ZSL^{+}22]. Blade [VDvG^{+}21]. blame [LKS^{+}20, LGFD21, SSC^{+}17, WMW18]. block [RFH^{+}22]. block-based [RFH^{+}22]. blockchain [HTLS22]. blocking [FF20, PVV^{+}17]. blocks [CB17]. blows [BBTK^{+}17]. Bonsai [CB18]. Boolean [MvdP20, MvdP21, PC22]. borrows [JDKD20]. both [FKR20]. Bottom [MNB^{+}22, BPP20, Lee21]. Bottom-up [MNB^{+}22, BPP20, Lee21]. bound [DR19]. boundary [FSSW21]. Bounded [BGM^{+}23, CPKG17, WJS23, AJRG21, ABM^{+}21, BMTZ21, BMTZ22, BGPT19, KESJ18, MJ21]. bounded-memory [ABM^{+}21]. Bounds [FCS^{+}23, AGLK18, HKGK20]. Bowtie [RHdSOZ23]. boxes [BSLBG22]. bracketed [SDB19]. brainer [DMS17]. breaking [GRS^{+}21, MNT20]. Bridging [LMÖ^{+}22, WDR18]. broken [AZMT18, GJK^{+}21, HFS22]. browser [KSP22]. bug [DRS21, LWNN19, PS18]. Bugs [PP22, BSPC19, BFSK20, CSD^{+}21, LRV^{+}22, MN18, RS20, SYW^{+}21, SJL19, WZS19]. Build [MMJ18, SMTH20, SCMS20, WSH^{+}19]. building [DFPG21, WLH^{+}17]. built [DTM^{+}18]. built-in [DTM^{+}18]. Bullet [CLO^{+}23]. bunch [FDvdHP22]. bunched [FDvdHP22]. bureaucracy [Chl21]. bytecode [CNRG19, WMLK18, XK21]. Byzantine [VREV19, WBdG^{+}23].

C [BDCN20, BBG^{+}20, COHY17, CNH20, ESDH21, EBP^{+}23, GLBP22, KKCV17, KLSV18, LSM^{+}22, LBR20, MKM^{+}22, MRdAP18, PMS^{+}23, SCK^{+}20, Str20, WCMH19, WXWS20, WZSK22, WCH^{+}19, XL21, ZCH23]. C-assembly [SCK^{+}20]. C-like [COHY17]. C/C [KLSV18]. C11 [DD22, ML21, YM21]. C11-style [ML21]. C4 [LyXK^{+}22]. CAAT [HMdL22]. cache [BKPS18, BZKT21, GCY^{+}20, XCB^{+}18]. caching [BC18, WSG^{+}20]. CakeML [GLPS^{+}20, ONK^{+}17]. Calculating [PH21]. calculation [BH22]. Calculi [AGK23, KyKS22, VRC22]. Calculus [BKK^{+}23, BDS23, GSF^{+}23, YTO23, ABH^{+}21, BGM19, BMP20, CdV20, DdVMY19, FKR20, FKT21, Ham17, KLL20, Mor19, Tej20, ZK22]. Call [HH19, AG22, BBBK17, BCD22, FKR20, JO22, KMLD20, PWSD20, ST21]. Call-by-need [HH19]. call-by-push-value [KMLD20]. call-by-value [AG22, FKR20, HH19, ST21]. call-site [JO22]. callback [GAGG^{+}18]. callbacks [AGR^{+}20]. calling [DAJE20]. calls [ONK^{+}17]. CAMP [CPY20]. Can [CAMS20, RFH^{+}22, CSD^{+}21, JGMP22, JO22]. canonical [BvGKJ17, BCH^{+}22, FJM19]. capabilities [BSO20, BSLBG22, CK20, GGV^{+}21, GTB22, SDB19, SKB^{+}22, VPD19]. capability [GGV^{+}21, GTB22, SBO20, SGD17]. capability-passing [SBO20]. Capturing [KSSL18]. care [BPPS18]. carte [MMJ18, vdRPR^{+}22]. Case [WL23, BGWXP22, WH19]. cases [CLNL22]. cast [MHN19]. Casting [MHN19]. Casts [CCW18, LSM^{+}22]. Catala [MCP21]. Categorical [Mel19, ABG^{+}21]. categories [CK18, CS21, Ell17a]. category [KyKS22]. category-theoretic [KyKS22]. Causal [CC23, DTSLT23, GGN^{+}21]. Causality [CC23]. cause [WMW18]. causes [BFSK20]. Cedille [SJSM20]. Celsius [BL22]. centric [CCP^{+}18a, CPT19]. cérises [GTB22]. CertiCoq [PLA21]. Certificates [LCT^{+}23, CMS22]. certification [MMS^{+}22, USM22, UCWZ20]. Certified [SBM20, HH20, JBK22b, SScWS19]. Certifying [IEC22, WCMH19, WGP^{+}21]. CFA [JO22]. CFL [CMS22, KD23, Mil20, SWYZ22]. CFL-based [Mil20]. changes [BAY20, GRS^{+}21, MNT20]. changing [Str20]. channel [FDvdHP22, GCY^{+}20, SM20]. channel-based [FDvdHP22, SM20]. channels [BZKT21]. Chaperone [MP17]. Chase [CGL^{+}23]. CHC [GTU23]. check [FZL18]. checked [MKM^{+}22]. checker [GJS20]. checkers [PES20, RvAP^{+}20, ZvAV22]. Checking [KTST23, KLV23, KPH22, AAJN18, AAJ^{+}19, BE19, CVG^{+}17, KLSV18, KRV19, KKT19, MKV22, SBF^{+}20, TU22, UAM17, YNIC19, ZvAV22]. checkpoints [MCL17]. chemistry [OLC^{+}18]. Chez [FDD^{+}19]. child [BS17]. chip [OLC^{+}18]. Choice [AB23, CHH^{+}23]. choreographies [HG22, MYZ20]. Circuit [KO23]. circuits [HRH^{+}21, Mat22]. clairvoyant [HH19]. Clarke [LYSM22]. class [BXMS19, EDWL21, HLL20, Kov20, ML20a, XCIL22, XPL^{+}22]. classes [SKR20, ZM17]. Classical [VLRH23, KMP19]. clauses [BOR18, KSG22]. Client [QKB21]. Client-server [QKB21]. Clojure [Hic20]. closed [CS21]. Closure [PA19a, DEO^{+}20]. closures [WBM^{+}21]. CLOTHO [RNDJ19]. cloud [ZPG^{+}17]. clustering [BS21]. CN [PMS^{+}23]. co [CFD^{+}17, CA18a]. co-design [CFD^{+}17]. Coalgebraic [JW23]. coarrays [RLS20]. Coarsening [SZD^{+}21]. coaxioms [ADZ17]. codata [BJSO20]. Code [BJP23, BBP23, MGB^{+}23, PMS^{+}23, ABF20, BJC^{+}22, BCH^{+}22, BT18, BSZL^{+}18, BAY20, BPB^{+}21, CL21, DFPG21, GS22, LHJ^{+}18, LWNN19, LMM^{+}17b, LYB^{+}19, LC21, MCF^{+}22, MT21, PvAPV22, RH22, SVR17, SCZW20, VPD19, VDvG^{+}21, VN18, WNC^{+}19, XCB^{+}18, YAY20, YMJ17, ZBG^{+}22]. Code-Generating [BJP23]. Coeffects [BDG^{+}22, DG22b]. cognitive [DNFK22]. Coherence [BXMS19]. coherency [CFL17]. cohesion [Kav19]. coin [CY19]. coinduction [San22]. cold [BBTK^{+}17]. Collapsible [FGS^{+}18]. Collapsing [AR18]. Collection [MCP23, BDCN20, BP21, DEH^{+}18, MP22, SHSO22, UAM17]. collections [DTM^{+}18, EE18]. collective [EWR19]. combating [GRS^{+}21]. combination [MR22, SdSCQ19]. combinatorics [Mel19]. combinators [WWP20]. Combining [DTSLT23, Lee21, LWUD21]. Communication [DFLL23, CHJ^{+}19, DEO^{+}20]. Commutativity [FKP23, CFKP22]. comonadic [Cho22]. comonads [CK20, HT18]. compact [CS21]. Compacting [BS21]. Comparative [WLJ^{+}23]. CompCert [BBP23, SCK^{+}20]. CompCertELF [WXWS20]. CompCertM [SCK^{+}20]. Competitive [MAH18]. Compilation [CA22, HHY^{+}21, LB23, BBP20, Cra19, KKPS21, Kov22, LVH^{+}22, OUM18, PBC^{+}21, RKV21, Ser21, VPD19, VCD^{+}21, WXWS20, WZSK22, XL21, XK21]. compile [RAT17]. compile-time [RAT17]. Compiler [BBP23, MTDC19, AHM^{+}17, BH22, BBF^{+}21, BBG^{+}20, BPB^{+}21, CKP22, GCS^{+}18, KKC^{+}17, LWY^{+}22, Mai17, PA19b, SSH^{+}19]. compilers [CSD^{+}21, CKA18, DELT17, LBR20, MSRH20, MJK^{+}17, PH21]. Compiling [COER19, Ell17a, PG21, SBO20, WBTR20, NWP^{+}18]. Complete [GFD19, AAR20, BKKM21, CK18, DK19, EE18, Pav20, SGL^{+}21, UST18, VTC^{+}18]. completely [vdRS22]. Completeness [CDG22]. completion [PvAPV22, WDS17]. Complexity [FISS20, KD23, WYT^{+}22, AD17, BE19, DHP18, INN21, KP22, LZR21, MP21a, WWC17]. Complexity-guided [WYT^{+}22]. Component [KO23, DEH^{+}18, MJ22, RRG^{+}21].

Component-Based [KO23, MJ22, RRG^{+}21]. components [GLSY20, KE19]. composed [BC18]. Composing [WMM20]. composition [KKCV17, RK22, SLDN17]. Compositional [DPQS18, KPE18, LSC^{+}22, MHLG23, PLA21, SDdSO22, VSC23, BCB21, BGOS18, GJS20, HKSS21, JRB^{+}22, KNNJ18, KSW18, SKE^{+}20, SGD17, ZBY^{+}21, ZXSD21]. Compositionality [GY23]. compositionally [DFD22]. comprehensive [BFSK20]. Computable [HSS23, SMC21]. Computation [RZ20, CT19, DSLH20, Ham17, KVT20, MW20a, Mul22, PGFP17]. computational [CH19, CS21, CTW21, Kes22]. computations [ABG^{+}21, ADZ17, OCDR17]. computer [NWP^{+}18]. computer-aided [NWP^{+}18]. Computing [FSX^{+}23, KCB^{+}23, VLRH23, AFF^{+}18, GTF^{+}20, JPBG19, PP22, SLJ20]. concatenation [HJL^{+}18]. concentration [BZSL19]. conceptual [BL22]. Concurrency [FG18, SL23, BFSK20, FDvdHP22, JRB^{+}22, KESJ18, KLSV18, MDBL20, PFD^{+}18, SM20, YM21]. Concurrency-aware [FG18]. Concurrent [FKP23, RBDO22, SP22, AG20, BP19, BHP^{+}19, EE18, FNB^{+}21, FKE^{+}20, FF20, FRS^{+}21, GSS^{+}22, KSW18, LF18, MJP20, MJ21, MWW22, NBDF19, ÖN22, PKSW21, SFH22, SRF^{+}20, TB19, UAM17]. condition [Jac21]. Conditional [SCL^{+}23, WL23]. conditioning [DK20, NcS17]. conditions [GKJ^{+}18, ZZX^{+}20]. cones [EPT18]. Confidential [FSX^{+}23]. configuration [SZD^{+}17]. Conflict [ZH18]. conjunction [LLSS22]. Connectivity [JBK22a]. Conquer [ADH^{+}23]. consensus [DEO^{+}20]. Consequence [KKZ23]. Consequence-Finding [KKZ23]. conservativity [KP18b]. Consistency [KLV23, SFH22, AAJ^{+}19, BE19, GKMB17, HMdL22, KSTM20, KEW^{+}20, RMV22].

Consistency-preserving [SFH22].

consistent [HKSS21, RNDJ19]. constant [BBG^{+}20]. constant-time [BBG^{+}20]. Constrained [ME17, BOR18, KSG22]. Constraint [ESCL22]. Constraint-based [ESCL22]. Constraints [DGGM23, JJR23, CCH^{+}18, CFLH^{+}22, HJL^{+}18, ML20a, ZHL^{+}20]. Constructing [KKA19]. construction [DSPC^{+}19, LYU^{+}22, SPV18]. constructors [JME20]. container [WYT^{+}22]. Context [BMTZ21, BMTZ22, BGM^{+}23, LL23, FSD21, JJO18, JJCO17, LSDZ22, LTMS18, LWNN19, LH22, LX19, PP20, PXA^{+}21, SWYZ22, TLM^{+}21]. context-aware [PXA^{+}21]. context-based [LWNN19]. Context-Bounded [BGM^{+}23, BMTZ21, BMTZ22]. Context-Free [BGM^{+}23, FSD21, LSDZ22]. context-sensitive [LH22, PP20, SWYZ22]. context-sensitivity [JJCO17]. Contextual [FCY^{+}20, SCL^{+}23, WCGC18, ZN21, BAY20, Jab20, JGMP22, TSKJB18, ZA22]. continuation [ABD21]. continuation-passing [ABD21]. Continuations [SU23, CA18b, COER19, PG21, TB19]. continuity [BCDG22, SMC19]. continuous [BKMMP19, WCGC18]. contract [AGR^{+}20, BEM^{+}21, MNTHV21, NGTHV18, SNJ^{+}19]. contraction [KPP21]. contracts [ASD^{+}21, CGL22, CWG^{+}22, END^{+}18, FGS^{+}18, GKJ^{+}18, GLTS22, GAGG^{+}18, HFS22, MP17, SGL^{+}21, WZS19, WCD17, WMW18, ZBS^{+}22]. Control [KCB^{+}23, AFF^{+}18, BGMW20, DAY20, FBP^{+}21, FKLP17, GSW21, HLL20, KKSL22, PSY^{+}20, Ram22, RMH21, SDB19, ZSM20]. control-flow [KKSL22]. controlled [MDBL20]. conventions [DAJE20]. conversion [AÖV18, Ada19, DMS17, PA19a]. convex [MMS^{+}22]. coordination [BNS20, MYZ20]. Copilot [BJP23]. Copy [XK21, UAM17].

Copy-and-patch [XK21]. copying [UAM17]. coq [BSZL^{+}18, ADH^{+}23, AHM^{+}17, BCH^{+}22, CHH^{+}23, KZK^{+}18, SM19, SBF^{+}20, yXZH^{+}20]. CoqQ [ZBS^{+}23]. Core [GSF^{+}23]. coroutine [Spi17]. coroutines [IEC22]. Corpse [MNTHV21]. correct [ASD^{+}21, DSPC^{+}19, FKE^{+}20, FSD21, KJK^{+}22, QSL17, SBF^{+}20]. correct-by-construction [DSPC^{+}19]. correction [LSSO18]. Correctly [KZA^{+}23, LAGN21, LN22, PG21]. Correctness [FSY^{+}18, ZDS23, BKV^{+}21, LSA18, PA19b]. correlated [ZPG^{+}17]. correlation [BAS^{+}18, GRB20]. Cosmo [MJP20]. cost [AMS20, ABD21, CPY20, DFS18, GLPS^{+}20, KNJ^{+}22, MH21, NSGH22, QGG19, RBG^{+}18, RGGH21, WKH20, WAA22]. cost-aware [CPY20, NSGH22]. costs [BDA20, CCW18, OD18]. Countable [AB23]. counter [HC21]. Counterexample [GRB20]. Counterexample-guided [GRB20]. countering [GFFS17]. counting [LL22, THL^{+}20]. counting-set [THL^{+}20]. coupled [VVB22]. coupling [AH18]. couplings [VVB22]. course [Chl21]. Coverage [LHP19, LWY^{+}22, GJS20]. Coverage-guided [LWY^{+}22]. CPS [BCRA18, DMS17, ST21]. CPU [IDSW21]. CPU/FPGA [IDSW21]. crashing [SYW^{+}21]. CRDTs [LPM^{+}22, NGR^{+}22, WMM20]. credits [SGT^{+}22]. crisis [EDWL21]. Cross [DEH^{+}18]. Cross-component [DEH^{+}18]. crowded [Str20]. Cryptographic [GSF^{+}23, VDvG^{+}21]. Cubical [VMA19, CH19]. CUDA [MH21]. Customizable [WKN^{+}23]. Cutting [CGL^{+}23]. cycle [MW20b]. Cyclic [KPP21, TU22]. cyclic-proof [TU22].

D [BAP20]. Dargent [CLO^{+}23]. dark [MHN19]. Data [CCP^{+}18a, CLO^{+}23, JJCO17, PSW21, RHdSOZ23, SSSC^{+}22, ZMSD23, ZDDJ21, BJSO20, Cha20a, CCP18b, CWG^{+}22, Coh18, DLC22, DAB^{+}21, ECK^{+}22, FMTSL20, GTF^{+}20, IS18, JJO18, KSG22, KPSJ19, KRV^{+}21, KSW18, LRGC19, LPR^{+}20, MGT21, MW20b, MM19, PJP^{+}18, QSL17, SSC^{+}17, WDS17, WCMH19, WBW^{+}20, WCBG17, YvGK18, YDJD21, YD22, YC22]. Data-centric [CCP^{+}18a]. data-dependence [CCP18b]. Data-driven [JJCO17, SSSC^{+}22, ZDDJ21, JJO18, SSC^{+}17]. data-structure [QSL17]. database [RNDJ19, RS20, WDLC18]. database-driven [WDLC18]. Dataflow [MKLR20, BBP20, HKVR21, SAB17]. Dataflow-based [MKLR20]. Datalog [BGC20, BGC23, ML20a, RMZ^{+}20, SBEV18]. Datatype [KCL22, JR21]. Datatype-generic [KCL22]. datatypes [EC22, JME20, SMC21, XEdSO20]. day [BPPS20]. dead [BBTSTH17]. Deadlock [CSM^{+}17, JB23, JBK22a, JBK22b, KP18a]. debugging [CN22, CG21, LZY^{+}20, PN17]. Decentralized [SSS^{+}23, ZBS^{+}22]. Decidability [AÖV18, KP22]. Decidable [KM23, MPAG20, CCH^{+}18, Ham17, HL20, KLO^{+}22, MP20, PLSS17]. Deciding [AAB^{+}21, AKK^{+}21, BCK^{+}21, MMK^{+}20]. declarative [CFKP22, HLV21, NPZ^{+}20, RvAP^{+}20]. Declassification [MHLG23]. decoders [DSPC^{+}19]. decompilation [GLTS22]. decompiling [NWP^{+}18]. decomposing [SPV18, WMM20]. Decomposition [BJSO20, RFH^{+}22]. Deconstructing [BDS23]. Deductive [CMP20, FMTSL20, LCT^{+}23, LWUD21]. Deep [SADC23, SGL^{+}21, ZPZX21]. DeepBugs [PS18]. DeepSEA [SScWS19]. default [LMM17a]. defeating [DMB20]. defects [RAJ^{+}17]. define [CR20]. defined [FKLP17]. Definitional [GCST19, DLNV17, PRT^{+}18, WDR18, vdRPR^{+}22]. Definitions [MHLG23, KP18b, RSY21].

DéjàVu [LMM^{+}17b]. delayed [ABM^{+}21]. Delimited [SU23, CA18b, FKLP17]. demand [PPSW20, SNCM19]. demand-driven [PPSW20, SNCM19]. Demystifying [WZD^{+}19]. Denotational [CLD20, ŚKV^{+}18, LCTS^{+}20]. dense [KRV^{+}21]. deoptimization [BBF^{+}21, FSY^{+}18]. Dependence [GRXB19, LVSZ23, BGHT22, CCP18b, ZYT^{+}19]. Dependence-Aware [LVSZ23, GRXB19]. dependencies [JRB^{+}22, PNPW22, SMTH20]. dependency [Cho22, TCJ20]. Dependent [CBTB20, SU23, AVW17, CFLH^{+}22, CEEW21, CA18a, CA18b, DFS18, ETG19, FKT21, GSB19, LZ18, MPAG20, NVD17, PT20, RKHL17, RL19, SJL19, TV20, WFC^{+}20, WVdAE17, WCVE19, YAC21]. Dependently [CDD^{+}19, ECK^{+}22, EGT22, FRS^{+}21, PH21, SM19, SRF^{+}20, Tej20, VMA19]. Dependently-typed [ECK^{+}22, PH21, SM19]. derivation [DSPC^{+}19, IEC22, LZ18]. Derivative [HBP19, JKT21]. derivative-based [JKT21]. derivatives [HBP19]. derive [ZvAV22]. derived [RvAP^{+}20]. Deriving [LA21, KPW18, KKSL22, PES20]. descendants [FMG^{+}20]. deserves [Coh18]. Design [CNRG19, BCC^{+}18, CFD^{+}17, GV19, MRG18, NWP^{+}18, WSG^{+}20]. Designing [TGKV20]. Designs [WLJ^{+}23, ZR21]. despite [HFS22]. dessert [GLPS^{+}20]. desugarings [BCB21]. Detecting [MNT20, RAJ^{+}17, WZS19, ZZX^{+}20, BZKT21, SCMS20, SYW^{+}21]. Detection [AZMV23, AMT17, BGOS18, DMB20, DRS21, GRXB19, GAGG^{+}18, LWNN19, MKV18, PS18, SLO19, WAA22, WCBG17, ZPZX21]. Deterministic [KVT20, SLDN17, UST18]. development [WKS18]. devices [PGFP17]. diagnosis [LSSO18, LSZ^{+}20, SSC^{+}17]. Diagrammatic [BHP^{+}19]. Diagrams [BDS23]. Diamonds [BGM21]. dictionaries [SHSO22]. dictionary [EZYS22]. dictionary-passing [EZYS22]. did [WC17]. Difference [XHdSO23, YF18]. differences [BFSK20]. Different [AGK23]. differentiable [AP20, SFVJ19, SZD^{+}21, SMC21, WZD^{+}19]. differential [ADN22, AH18, BCK^{+}21, KNSA20, NDA^{+}19, WCHRP17, ZRH^{+}19, ZRH^{+}20]. Differentially [ZHL^{+}20, SA19]. Differentially-private [ZHL^{+}20]. Differentiation [LHSM23, Ell18, Ell21, KJK^{+}22, LYU^{+}22, MP21b]. diffing [MS19]. DiffStream [KNSA20]. Digging [JGW^{+}20]. Dijkstra [MAA^{+}19, SZ21]. dimensional [DR19]. DimSum [SSS^{+}23]. Direct [MSBO23, ZGHH23, MKC18]. directed [CPB^{+}22, FSSW22, GTB22, OUM18, RNDJ19]. discipline [VHEZ21]. Discover [DTSLT23]. discovery [JGW^{+}20]. discrete [HVM20, SFRM20]. Disentanglement [WYFA20]. dispatch [BCG^{+}20, FCY^{+}20, PHSR19, PP20]. distancing [DG22a]. Distributed [CHJ^{+}19, GGN^{+}21, WJS23, WKS18, ABIK22, BvGKJ17, BGMW20, BVG19, DPQS18, GHMM20, GKMB17, GLSY20, HKSS21, JWJ^{+}21, KEW^{+}20, LMÖ^{+}22, MDSM19, OMN^{+}18, OMO19, PLSS17, SKS^{+}19, SWT18, SKE^{+}20, VHEZ21]. Distributing [HdSO21]. distribution [BBB^{+}17]. distributions [SFRM20]. divergence [RMH21]. Divergent [FCS^{+}23, ADZ17]. diversity [BJSO20]. Divide [ADH^{+}23]. Divide-and-Conquer [ADH^{+}23]. Do [GLPS^{+}20, AMP^{+}20, UdM22]. Does [LKS^{+}20, MTDC19]. domain [BKF20, PLS^{+}19, SDdSO22]. domain-specific [BKF20, PLS^{+}19, SDdSO22]. domains [SPV18]. don’t [GFFS17]. DOT [BGWXP22, GST^{+}20, KLL20, RL19]. Doubly [WJS23]. Doubly-Unbounded [WJS23]. Down [BOW^{+}23, BCB21, Lee21]. DProf [BVG19]. driven [GBR^{+}20, JJO18, JJCO17, MJK^{+}17, PPSW20, SSC^{+}17, SSSC^{+}22, SNCM19, VHEZ21, WDLC18, ZBS^{+}22, ZDDJ21]. DSL [ZYB^{+}18]. Dual [SV23, UTGK23, LYSM22, PWK^{+}22, ZRH^{+}20]. Dual-Numbers [SV23]. duality [HdSO21, PWK^{+}22]. Duet [NDA^{+}19]. duplicates [LMM^{+}17b]. Duplo [LJ20]. Durable [BGJ^{+}21, DRV23, ZFS^{+}19]. DWARF [BKN19]. DWARF-based [BKN19]. Dyck [CCP18b, KP22, LZR21, LSZ22, LZR23, SWYZ22]. Dyck-CFL [SWYZ22]. Dyck-reachability [LZR21, LSZ22]. Dynamic [AZMV23, FKRS23, MSI19, PP20, AFF^{+}18, BDA20, CCP^{+}18a, CPT19, CA22, DTM^{+}18, DPQS18, FSY^{+}18, JSXH20, KMGV22, LAF^{+}20, LP20, LSZ22, LSC^{+}22, MKV18, MRG18, Pav20, SPKT18, WCBG17]. dynamic-priority [LSC^{+}22]. dynamically [CHJ^{+}19, INN21]. dynamically-instantiated [CHJ^{+}19]. dynamism [BCC^{+}18]. DynamiTe [LAF^{+}20]. Dynaplex [INN21].

E-Graphs [CKN^{+}23]. e-matching [ZWWT22]. early [Sym20]. easy [Par20]. ECROs [DFPG21]. edge [DWZ20]. edit [MGL^{+}19, ZBG^{+}22]. Editorial [Wad17]. edits [PBS^{+}17]. education [CDH17]. Effect [BSO18, MJK^{+}17, SU23, XBH^{+}20, BPPS20, BSO20, FKLP17, GLBP22, GCS^{+}18, KKPS21, LLB^{+}20, SBO20, XL21, XCIL22, YW21, ZM19, dVP21]. Effect-driven [MJK^{+}17]. Effectful [BBP23, DG22a, KyKS22, MJ22, SRF^{+}20]. Effective [KLSV18, KRV19, LVSZ23, ZGSN17, MN18, TTS18]. effectively [BKV^{+}21, GAGG^{+}18, Pav20]. effectiveness [WZS20]. Effects [BSO20, BSLBG22, HLL20, PvdR23, Ahm18, AP21, BPPS18, BPPS19, BAS^{+}18, DG22b, FKLP17, HK22, HT18, HC21, MvdP20, NPWW22, PT20, STI20, SM20, SB19, WBTR20, ZN21]. efficiency [ADV21, HLL20]. Efficient [AAM^{+}21, CFL17, GGV^{+}21, JKXH23, KKPS21, KRV^{+}21, LSZ22, SFVJ19, SV23, ZFS^{+}19, ACN18, AWA21, AYB^{+}22, BCD22, GTB22, GXD20, KJK^{+}22, LA21, MS19, MNTHV21, MT21, ONK^{+}17, SCJG21, SC20, SBM20, SAB17, WNC^{+}19, WHZ^{+}19, XL21, YSHZ21]. efficiently [HJL^{+}18]. egg [WNW^{+}21]. Elaborating [CA18a]. Elaboration [Kov20, PvdR23]. elaborator [KCL22]. elementary [LN22]. Elements [GD23, LP20]. ELF [WXWS20]. Eliminating [MV20, VDvG^{+}21]. elimination [CLNL22, DS17, OBS^{+}22, PT20]. Elipmoc [GLTS22]. elision [LP20]. Emacs [MS20]. embedded [PZR^{+}17]. embedding [NA18, SCZW20]. embedding-projection [NA18]. embeddings [Chl21, LW22a, SDdSO22, WWGW20]. embracing [UdM22]. empirical [CAMS20, GCS^{+}18, KSTM20, MHN19, PP22]. empirically [TGKV20]. Empowering [MT18]. Enabling [MSRH20, WJS23, ZBS^{+}22]. encapsulation [SDB19, TSKJB18]. encoders [DSPC^{+}19, GXD20]. End [CC22, RFH^{+}22]. End-to-end [CC22]. end-users [RFH^{+}22]. Endpoint [GY23, CHJ^{+}19]. Enforcing [MGB^{+}23, NDA^{+}19, SDB19]. engineering [DAY20]. enhanced [San22]. enhancement [PXA^{+}21]. Enhancing [MKV18]. entangled [KGdV^{+}22]. Entanglement [WAA22, YMC22]. enumeration [Lee21, vdRS22]. enumerative [BPP20]. environments [FZSN20]. epoch [RV18]. EPR [PLSS17]. Equality [PT23b, EGT22, KFM21, NWZ^{+}21, PT22, WNW^{+}21]. Equational [GSF^{+}23, LL23, Mor19]. Equations [SM19]. equivalence [AAJ^{+}19, BCG^{+}21, CRMA20, Jab20, KPH22, Mil20, WCGC18, WDLC18, ZA22].

Equivalences [TTS18, TSKJB18]. erasure [SBF^{+}20, Tej20]. error [CN22, SEG^{+}22, WCC17, WCH^{+}19]. errors [BBZ18, CN22, DdVMY19, LSSO18, SSC^{+}17, SLO19, WC17, ZZX^{+}20]. essence [DLC22, Ell18, Par20]. estimation [MSRH20]. Eta [RHdSOZ23]. Ethereum [BEM^{+}21, GKJ^{+}18, GLTS22, LGTS20, SGL^{+}21, WZS19]. eval [BCG^{+}20, GDBK^{+}21, PGIY20]. evaluate [LGFD21]. Evaluating [CBA17]. evaluation [AVW17, AK20, BT18, BP18, LBH^{+}18, LCOC20, MKC18, MH21, PPSW20, PNWT22]. even [JO22]. event [BAS^{+}18, VHEZ21, ZBS^{+}22]. event-driven [VHEZ21, ZBS^{+}22]. events [Jac21]. eventual [GKMB17]. Every [Coh18]. everywhere [KMV19]. evidence [XL21]. evidently [XBH^{+}20]. Evolution [MS20, CNH20, Ing20]. evolving [HH20]. Exact [CMcS23, HVM20, SMC19]. Example [CGL^{+}23, ASMS20, JSXH20, VLG21, WFB^{+}20]. example-based [ASMS20]. examples [BFV18, GS22, GDM20, IS18, YAY20]. ExceLint [BBZ18]. Exceptional [FLMD19, PTFT19]. Executable [GB20, ZBY^{+}21]. executables [ZYT^{+}19]. Executing [KZA^{+}23]. Execution [LS23, MGB^{+}23, AJRG21, BGC18, CKT18, CB17, CN22, GCY^{+}20, KHLD22, MCL17, MNB^{+}22, MKV22, NPZ^{+}20, SLHR22, WBTR20, WMLK18, ZBS^{+}22]. exhaustive [AJRG21, CPKG17]. existential [EDWL21]. existentials [DK19]. Expansion [RHdSOZ23]. expectation [ABH^{+}21, BKKM21]. expectation-based [BKKM21]. expectations [WKH20]. Expected [BKK^{+}23, LHSM23, ABD21, BEG^{+}18, WFC^{+}20, WKH20]. experience [ACM^{+}18, AHM^{+}17, BRS18, BS17, BSZL^{+}18, CDD^{+}19, FDD^{+}19, HP19, HTLS22, RK22, TRWS22, YW19]. experiment [RFH^{+}22]. experiments [TBB^{+}19]. explain [RSPC17]. explodes [BT18]. Exploiting [KNNR17, CFL17]. exploring [CKP22]. Exposing [GCY^{+}20]. expressing [HLK^{+}20]. Expression [LKM23, SJSM20]. Expressions [LC23, LP20, MKTD17, PHXD19]. Expressive [DGGM23, BKKM21, FKLP17, LWUD21, NDA^{+}19, XLD20]. expressiveness [PMD21]. ext4 [KKRV21]. extended [SWYZ22]. Extending [RMV22]. Extensible [RHdSOZ23, KKCV17, KJJ^{+}18, MM19, SRF^{+}20, WNW^{+}21]. extension [YvGK18]. extensionality [BGG^{+}20]. extensions [KKCV17]. extent [QRS22]. extracting [HWC^{+}21]. extraction [CLD20, KMLD20].

f [JGMP22, RHdSOZ23]. F# [Sym20]. F* [PZR^{+}17]. failure [LSZ^{+}20]. failures [HKSS21, ZPG^{+}17]. Fair [CP22, LNO^{+}21]. Fairness [MWA19, ADDN17, BZSL19, UCWZ20]. FairSquare [ADDN17]. Familia [ZM17]. Familial [Hir19]. families [ME17]. family [ZM17]. Farkas [LFY^{+}22]. Fast [CVG^{+}17, CN22, GTF^{+}20, JW23, Pav20, WNW^{+}21, BKN19, BPB^{+}21, LH22, LX19, PUW^{+}21, SV21, Tit22, USM22, WCR19, WSH^{+}19, XK21]. Faster [Spi17, WMLK18]. Fat [ZCH23]. Fault [Per18, WBdG^{+}23, BGC18, HKSS21, LZ17, MDSM19, VHEZ21]. fault-aware [HKSS21]. fault-tolerant [MDSM19, VHEZ21]. faults [SCMS20, VREV19]. Featherweight [GHK^{+}20]. features [COHY17, JLO20]. Feedback [GBR^{+}20]. Feedback-driven [GBR^{+}20]. Feller [BCDG22]. Few [SADC23]. Few-Pixel [SADC23]. FFT [Ell17b]. fibrations [MPV18]. fibred [Ahm18]. file [SZD^{+}17]. files [WXWS20]. Finding [AZMT18, BT18, KKZ23, LRV^{+}22, RS20, BBZ18, DNFK22]. Fine [KD23, FSK^{+}22, Kes22, MP21a]. Fine-Grained [KD23, FSK^{+}22, Kes22, MP21a]. Finitary [OUM18]. finite [KM22, WDS17]. fire [PT20]. First [XCIL22, EDWL21, HLL20, Kov20, ML20a, MKV18, NJA20, PHL^{+}18, WBE20]. First-class [XCIL22, EDWL21, HLL20, Kov20, ML20a]. first-order [PHL^{+}18]. fitch [BGM19, VRC22]. fitch-style [BGM19, VRC22]. fix [BSPC19]. fixed [WC17]. fixing [FGS^{+}18]. Fixpoint [BKMMP19, UTGK23, CTR^{+}20, KVT20]. Fixpoints [ML20a, GMOO21, MPB^{+}22]. flaky [LWW^{+}20]. FlashFill [CGL^{+}23]. FlashProfile [PJP^{+}18]. Flexible [yKMUW22]. floating [Ada19, LAGN21, ZZX^{+}20, ZGS^{+}22]. floating-point [ZZX^{+}20, ZGS^{+}22]. flow [AFF^{+}18, DAY20, Kav19, KKSL22, KSW18, PSW21, PSY^{+}20, Ram22, RMH21, SDB19, SGL^{+}21, SCZW20, ZSM20, ZK22]. Flow2Vec [SCZW20]. FlowCFL [Mil20]. Fluent [RG23, YNIC19]. fly [MGL^{+}19]. focusing [RZ20]. FOL [MPB^{+}22]. fold [JGW^{+}20]. forever [BGM21, SZ21]. forking [LyXW21]. Formal [BBG^{+}20, FS22, HPRW21, JPBG19, MJ21, YZZ22, DAB^{+}21, EUR^{+}17, GB20, LRS^{+}20, PNWT22, SLJ20, ZBY^{+}21, Rei17]. formalising [RWV19, RMV22]. formalization [BP19, ZdSOS19]. formalized [HVK19]. formalizing [RL19]. Formally [BBF^{+}21, BBP23]. Format [CKA18]. formats [DSPC^{+}19, GLR^{+}20]. formula [BBZ18, BJC^{+}22]. formulas [KM22]. Formulog [BGC20]. Fortran [RLS20]. forward [SCJG21]. Foundation [ZDS23, PNWT22, SLJ20]. Foundational [ZBS^{+}23, ZSL^{+}22]. Foundations [BBBK17, KSTM20, LMP18, AFH^{+}18, BGWXP22, DAB^{+}21, HVK19, JPBG19, JJKD18, OBL^{+}18]. FPGA [IDSW21]. FPL [PUW^{+}21]. FR [HVK19]. Fractional [DMS22, CS21]. fragments [GM21, HL20]. frame [LL22]. Framework [AGK23, BL22, CTR^{+}20, EUR^{+}17, GSS^{+}22, KKCV17, KJJ^{+}18, LBH^{+}18, LZ18, LJ20, MA18, NSGH22, PJP^{+}18, SSK17, SHW^{+}20, SSK22, WCHRP17]. Free [BGM^{+}23, JB23, AJSW17, BDYG^{+}21, CN22, Coh18, FSD21, GAGG^{+}18, LSDZ22, MW20b, RST20, TTS18, YvGK18, ZGS^{+}22, ZFS^{+}19, ZvAV22]. freedom [JBK22a, JBK22b]. FreezeML [ESCL22]. frequency [ZHL^{+}20]. friendly [WCC17]. Full [CdV20, BCDG22, JDT22]. Fully [Cra19, JTD21, KyKS22, SYW^{+}21, KMP19, RL19, VPD19]. fully-abstract [KMP19]. function [CCH^{+}18, FCY^{+}20, Kov20, LN22, OBL^{+}18, ONK^{+}17]. Functional [DTSLT23, LC23, LB23, NWP^{+}18, RG23, ŚKG18, SHSO22, ZMSD23, ACM^{+}18, ACF17, AK20, AWA21, AG20, BH22, BWB^{+}21, Ber17, BG18, BBRM17, CDH17, Cha20b, Chl21, CRMA20, DLNV17, DA20, DMS17, Ell17b, EHAO18, HLK^{+}20, HP19, HWC^{+}21, HG22, HP22, HdSO21, JBK22b, KMLD20, KSSL18, LSSO18, LC21, MCF^{+}22, MNB^{+}22, OVCH19, Par20, PA19b, PN17, Per18, PH21, QGG19, Ram22, RSPC17, SC20, SFVJ19, SLO19, SM19, SAFF^{+}17, SC18, SJSM20, SYW^{+}21, SB19, UdM22, VS20, WWC17, WDR18, YF18]. functional-imperative [QGG19]. functions [BGJ^{+}21, CFLH^{+}22, EPT18, KSG22, RK22, SPKT18, SHdM20, SMC21]. Functorial [DLNS21]. functors [BGPT19, MLMD19]. Fusing [TRWS22, SSK17]. fusion [YW21]. Futhark [EHAO18]. future [JLP^{+}20, KSSL18, MWW22]. futures [CSM^{+}17]. FuzzFactory [PLS^{+}19]. Fuzzi [ZRH^{+}19]. Fuzzing [LS23, SM20, LWY^{+}22, MTDC19, MR22, PLS^{+}19, SYW^{+}21].

gadget [BPB^{+}21]. GADT [BGWXP22]. GADT-style [BGWXP22]. Galois [PNPW22]. game [CY19, Mel19, VMS^{+}22, dVPJ20]. games [ACM^{+}18, BKMMP19, FK18]. gap [LMÖ^{+}22, WDR18]. Garbage [MCP23, BDCN20, DEH^{+}18, MP22, UAM17]. garden [LyXW21]. gas [GKJ^{+}18]. Gauss [BLSS21]. GC [CFD^{+}17]. General [HP23, CR20, KJJ^{+}18, LYU^{+}22, MMS^{+}22, SSK17]. Generalizable [JXXH21]. Generalized [XL21, XHdSO23, DS17, Mil20]. generate [LAGN21]. Generating [BJP23, BP21, LPP18, LCT^{+}23, Mat22, WCH^{+}19, YNIC19, COHY17]. Generation [BBP23, LVSZ23, CPKG17, CL21, LFY^{+}22, RNDJ19, RGSNT17, SSSW21, SPKT18, WH19]. Generative [PWZS21]. generator [ASD^{+}21, JKT21]. generators [BLF^{+}19, KKSL22, LPP18, Mat22, NPZ^{+}20, SV21]. Generic [DFS18, Ell17b, EZYS22, KPW18, EC22, KCL22, SKB^{+}22, Yal17, dVPJ20]. Geo [BBB^{+}17]. Geo-distribution [BBB^{+}17]. geometric [GHMM23]. Geometry [CC23, GYK^{+}20]. Getafix [BSPC19]. Getting [PJD^{+}21, MAH18]. Ghost [LHC^{+}23, CMP20]. GhostCell [YDJD21]. Girard [Kes22]. GitHub [LMM^{+}17b, TRWS22]. Giving [HC21]. global [DFPG21]. globally [MA18]. GLORE [DS17]. glossy [ACM^{+}18]. GNN [WBS^{+}22]. GNN-based [WBS^{+}22]. Go [KSW18, BCB21, CSD^{+}21, EZYS22, CHJ^{+}19, GHK^{+}20]. good [LPP18, PT22]. goSLP [MA18]. GPU [AAM^{+}21, CNRG19, EHAO18, KCB^{+}23, SSR^{+}21]. GPU-based [CNRG19]. GPUs [CPKG17]. graded [CEEW21, yKMUW22, OLE19]. Gradients [RPF^{+}23]. grading [CRMA20, HP19]. Gradual [CL17, CLPS19, ITVW17, NLA19, TLT19, VTV18, WBW^{+}20, YTO23, BBTSTH17, CCW18, CCEW18, CC20, ETG19, EGT22, FGS^{+}18, FKT21, GFD19, ISI17, JTD21, LTTD22, LGFD21, MLBTT22, MP20, MSI19, MNTHV21, MT17, MT21, PCAGG21, RAT17, SCJG21]. Graduality [NA18, NJA20]. Gradually [MGT21]. Grafs [HLV21]. Grained [KD23, FSK^{+}22, Kes22, MP21a]. grammar [YNIC19]. grammars [AM17, HBP19, JKT21]. graph [HLV21, JLO20, KKSL22, Mil20, WCMH19, WWGW20, ZYB^{+}18]. graph-based [JLO20]. graph-manipulating [WCMH19]. graphics [DELT17, GYK^{+}20]. GraphIt [ZYB^{+}18]. Graphs [ASW23, CKN^{+}23, BLSS21, DAY20, JBK22a, Mul22, PBS^{+}17, RMH21, ZvAV22]. greedy [Ber17]. Gries [RLV20]. Grisette [LB23]. Groovy [Kin20]. ground [RWV19]. Grounded [BJP23]. groupoids [CKS22]. guarantees [BVG19, OMN^{+}18]. Guarded [SFH^{+}20, BGM21, MV19]. guards [GJS20, Rei17]. Guide [LVSZ23]. Guided [BDA20, JKXH23, LS23, GJJ^{+}20, GRB20, JPR22, KHDR21, LHP19, LTMS18, LWY^{+}22, MCF^{+}22, MJ22, MPB^{+}22, RMZ^{+}20, RFH^{+}22, WH19, WYT^{+}22]. Guiding [JSXH20]. GV [JBK22b].

Halide [CC22, NAJ^{+}20]. handcrafting [JLO20]. Handle [BPPS18]. handlers [BPPS18, BPPS20, BSO18, BSO20, FKLP17, GLBP22, KKPS21, SBO20, XBH^{+}20, XL21, XCIL22, ZM19, ZN21, dVP21]. Handling [Ahm18, CA18b, ZSM20, KRV19]. happens [MKV18]. happens-before [MKV18]. harder [HKGK20]. hardware [BGC18, CVS^{+}17]. harmonizing [CCW18]. Harper [SJSM20]. Haskell [BBN^{+}18, BSZL^{+}18, CDD^{+}19, HVH20, JGW^{+}20, KFEJ19, KRB^{+}21, LPR^{+}20, WVdAE17, WCVE19, XPL^{+}22]. HDL [FMG^{+}20]. Heap [MCP23, GM21, GFFS17, KSP22, MP22, MMK^{+}20, TLKC18, TGSM19, WGP^{+}21, WBW^{+}20]. heap-manipulating [MMK^{+}20, WGP^{+}21]. Heaps [GFFS17, EWR19]. Hefty [PvdR23]. help [RFH^{+}22]. Herbarium [SAFF^{+}17]. heterogeneous [PGFP17, SMM17]. heuristics [COHY17, JLO20]. HFL [KTST23]. Hidden [WSG^{+}20]. High [GLBP22, MCP23, CVS^{+}17, HLK^{+}20, HPRW21, LHJ^{+}18, LBH^{+}18, LBCRK22, SGDL20, SNB^{+}19, SM19, WNC^{+}19, XK21, ZYB^{+}18]. High-Level [MCP23, GLBP22, CVS^{+}17, HPRW21, LHJ^{+}18, LBCRK22, SGDL20, SM19, WNC^{+}19, XK21]. high-performance [HLK^{+}20, LBH^{+}18, SNB^{+}19, ZYB^{+}18]. Higher [ABG^{+}21, BOR18, CH19, GMS^{+}23, JB23, JJR23, KFEJ19, MHLG23, PPSW20, PT23a, PvdR23, ABG^{+}17, AK20, BWB^{+}21, BBN^{+}18, CK18, CGM18, DK20, DK19, EPT18, EHAO18, GM21, GGN^{+}21, HG22, HTLS22, Jab20, KJK^{+}22, LYU^{+}22, MP17, MJ20, NDA^{+}19, NGTHV18, RGGH21, RK22, ŚKV^{+}18, SPKT18, SMC21, SC18, SG21, UST18, VMA19, WCD17, ZdSOS19]. Higher-Order [GMS^{+}23, JB23, JJR23, MHLG23, PT23a, PvdR23, ABG^{+}21, BOR18, KFEJ19, PPSW20, ABG^{+}17, AK20, BWB^{+}21, BBN^{+}18, CGM18, DK20, EPT18, EHAO18, GM21, GGN^{+}21, HG22, HTLS22, Jab20, KJK^{+}22, LYU^{+}22, MP17, MJ20, NDA^{+}19, NGTHV18, RGGH21, RK22, ŚKV^{+}18, SPKT18, SMC21, SC18, SG21, UST18, WCD17]. higher-rank [DK19]. higher-ranked [ZdSOS19]. Highly [HFS22, FKE^{+}20]. highly-concurrent [FKE^{+}20]. Hindley [MSI19]. History [RLS20, SHK^{+}20, CB17, Hic20, Kin20, MHR20, MWW22, ML20b, Sym20, VHSS20]. hoc [JME20]. HOL [KP18b]. Holes [YGG^{+}23, OVCH19]. hood [CDH17]. Horn [BOR18, END^{+}18, JJR23, KSG22]. Horn-ICE [END^{+}18]. hot [BBTK^{+}17]. hs [BSZL^{+}18]. hs-to-coq [BSZL^{+}18]. hull [MMS^{+}22]. Hybrid [GY23, EZYS22]. Hygienic [CW20]. hyperproperties [GMOO21]. hypersafety [DFD22].

I/O [SJL19]. ICE [END^{+}18]. IDEal [SAB17]. idempotence [SJL19]. Identifying [VN18, FZSN20]. identity [LP20]. idioms [LSM^{+}22]. Igloo [SKE^{+}20]. illogical [HFS22]. imaging [AAM^{+}21]. impacts [BPB^{+}21]. Imperative [RSPC17, BCH^{+}22, HWC^{+}21, MCF^{+}22, NPZ^{+}20, PRT^{+}18, QGG19]. imperativity [UdM22]. implementation [BCD22, CNRG19, GV19, XCB^{+}18]. Implementations [FSX^{+}23, DEO^{+}20, LSA18]. Implementing [DD22, GSB19]. implications [FDvdHP22]. implicit [FSK^{+}22, KNNR17, Kov20, OBL^{+}18, ST21]. implicits [KMV19]. imprecision [CDG22]. Impredicative [PT23b]. impredicativity [SHJV20]. improvement [HH18]. Improving [JKXH23, LWNN19, CBA17, GM21, NAJ^{+}20]. Impure [CHH^{+}23, yXZH^{+}20]. in-place [Tit22]. in-situ [LSZ^{+}20]. incomplete [BGG^{+}20, MRdAP18]. inconsistencies [RH22]. incorporated [EJMmH17]. Incorrectness [O’H20, ZDS23, LRV^{+}22, RBDO22, YJY22, ZdAG22]. Incremental [HPC18, YHW20, ZvAV22, HH20, KPSL18, LH22, PES20]. Incrementalizing [SBEV18, DRS21]. independence [ACMZ21]. index [PJD^{+}21]. Indexed [AB23, DK19, GST^{+}20]. Indexing [SWYZ22, SKD21]. Induction [PWK^{+}22, HKGK20, San22]. Inductive [JKXH23, LC23, CH19, FSSW21, KKA19, LPP18, Lee21, MPB^{+}22, VMA19, ZXSD21]. inductive-inductive [KKA19]. industry [CDD^{+}19, TRWS22]. Inference [LRY23, LVSZ23, MRdAP18, AKL19, AYB^{+}22, AD17, EDWL21, ESCL22, FISS20, HGJ20, HMP^{+}22, HVM20, KNNJ18, KFM21, KSS20, LYRY20, LCTS^{+}20, LZ18, MSI19, NWZ^{+}21, Par20, PC22, PSW21, RRG^{+}21, ŚKV^{+}18, ŚKG18, SKB^{+}22, Tej20, VTV18, WNC^{+}19, XLD20, XEdSO20, ZdSOS19, ZDDJ21]. inferred [INN21]. Inferring [PKW17]. infinitesimal [Jac21]. Information [HSS23, AFF^{+}18, DNFK22, FISS20, HK22, Kav19, PSY^{+}20, ZK22]. Inhabitation [AGK23]. inheritance [WSG^{+}20]. initialization [AMT17, BL22, KLL20, LLB^{+}20, WSH^{+}19]. Initialize [WSH^{+}19]. inline [CGL22, WSG^{+}20]. Inlining [DPM23]. input [CPKG17, RGSNT17, WH19]. insensitive [GBTB21]. instances [BPPS20]. instantiated [CHJ^{+}19]. instantiation [LMP18]. instruction [SBM20]. Instrumentation [WCBG17]. Integer [LZR23, LSM^{+}22]. integer-pointer [LSM^{+}22]. integrated [MT18, SKS^{+}19]. integrating [IS18, RV18]. Intel [IDSW21, RLV20, RWNV20, RMV22]. Intel-x86 [RLV20, RWNV20, RMV22]. IntelliMerge [SZZ^{+}19]. Intensional [JR21]. Inter [TCJ20]. Inter-theory [TCJ20]. Interact [BJP23]. Interaction [CC23, yXZH^{+}20, ADV21, SZ21, YW21, ZPZX21]. Interactive [GDM20, ABF20, FBP^{+}21, KJJ^{+}18, MDSM19, SSSC^{+}22, ZGSN17]. interfaces [ZM17]. Interleaved [LZR23, KP22, LZR21]. Interleaved-Dyck [LZR23]. interlocked [SBM20]. Intermediate [KCB^{+}23]. intermittent [MCL17, SJL19, SLJ20]. Internalizing [ACMZ21]. Interpretable [ZR21]. Interpretation [Lem23, BGMW20, BPPS18, CDG22, CS21, EHAO18, FSSW22, FDvdHP22, GLR^{+}20, KE19, Kes22, LYU^{+}22, LLSS22, MJ20, RMH21, ZYT^{+}19]. interpretations [BGG^{+}20]. interpreter [Tit22]. interpreters [AR18, CNRG19, DLNV17, KPE18, PRT^{+}18, WDR18, WCR19, YZZ22, ZRH^{+}20, vdRPR^{+}22]. Intersection [DdVMY19, DFD21, XHdSO23, CL17, HdSO21, MSdSOK20, MPV18, MT18, WMW18]. Interval [WAPJ22, WWGW20]. intervals [SG21]. Intrinsically [PRT^{+}18, RKV21, vdRPR^{+}22]. Intrinsically-typed [PRT^{+}18, vdRPR^{+}22].

Introduction [OBS^{+}22]. intuitionistic [Kes22]. invariant [FISS20, KCBR18, LFY^{+}22]. invariants [END^{+}18, FSSW21, PWK^{+}22, WWC17]. invertible [MW20a]. IR [LWY^{+}22, SSH^{+}19, ZBY^{+}21]. IR-pass [LWY^{+}22]. Iris [GST^{+}20, dVPJ20]. irrelevance [GCST19]. Isabelle [KP18b]. Isabelle/HOL [KP18b]. iso [ZdSOZ20]. iso-recursive [ZdSOZ20]. Isolation [KNJ^{+}22, BKV^{+}21, KNNJ18, LRS^{+}20, LSC^{+}22]. isomorphism [SdSCQ19]. iteration [SHW^{+}20]. itself [JGMP22]. IVT [WHZ^{+}19].

Jacobians [LYSM22]. Java

[BP19, CNRG19, DRS21, LH22, MHN19, MKTD17, MV20]. JavaDL [DRS21]. JavaScript [AMT17, AZMT18, CVG^{+}17, GTT21, MLT17, MNT20, SMN^{+}18, Ser21, WBE20]. JaVerT [SMN^{+}18]. JIT [BBF^{+}21, BBP23, PBC^{+}21, XCB^{+}18]. job [WBS^{+}22]. joint [LWY^{+}22]. Jones [BP18]. Jones-optimal [BP18]. Julia [BCG^{+}20, BCC^{+}18, NBP^{+}18, PBC^{+}21]. Just [BPP20]. Just-in-time [BPP20]. JVM [CSD^{+}21, LMM17a].

Kahn [BM20]. Kami [CVS^{+}17]. Katara [LPM^{+}22]. Kater [KLV23]. Keep [FZL18]. kernels [MH21]. Kind [XEdSO20]. Kindly [RST20]. Kinds [DAJE20, CT19]. Kirk [HFS22]. Kleene [SFH^{+}20, ZdAG22]. knew [RAT17]. know [BG18]. Knowing [RvAP^{+}20]. knowledge [RAT17]. Known [SV23]. Koord [GHMM20]. Kripke [KHLD22].

Label [FKT21, TV20, LZ18]. Label-dependent [TV20]. labels [BPPS20, HC21, RKV21]. laboratories [OLC^{+}18]. laboratories-on-a-chip [OLC^{+}18]. LabVIEW [Kod20]. lambda [BMP20, CdV20, FKT21, MKTD17, AGK23, Mor19, YW19, YTO23]. lambda-calculus [BMP20, CdV20]. Language [LCT^{+}23, PvAPV22, SKS^{+}19, VLRH23, XTZ^{+}23, AP20, AK20, BKKM21, BBN^{+}18, BRS18, BBP20, BAP20, CPB^{+}22, CC22, DSLH20, GDM20, GHMM20, HGJ20, HTLS22, JJKD18, KKCV17, Kin20, KPH22, KM20, KPSL18, LSDZ22, MW20a, MCP21, NDA^{+}19, Pue17, RRG^{+}21, RGSNT17, SSS^{+}23, SFVJ19, SNB^{+}19, SScWS19, SC18, UdM22, VHSS20, VMA19, WCGC18, WWC17, YWDD17, ZPG^{+}17]. Language-Agnostic [LCT^{+}23, XTZ^{+}23]. Language-integrated [SKS^{+}19]. Language-parametric [PvAPV22, KPSL18]. Languages [KM23, RG23, WKN^{+}23, ACF17, BKF20, BJC^{+}22, COHY17, CKS22, CFD^{+}17, DTM^{+}18, Ell21, KPSL18, LCTS^{+}20, MRG18, PRT^{+}18, SPKT18, SDdSO22, XK21]. large [EHAO18, GDBK^{+}21, HFS22, KMV19, LWW^{+}20, LLZ^{+}22]. large-scale [GDBK^{+}21, KMV19, LWW^{+}20, LLZ^{+}22]. larger [RFH^{+}22]. late [KMP19]. Later [SGT^{+}22]. Lattice [HSS23, SBEV18]. lattice-based [SBEV18]. lattices [BKMMP19]. law [MCP21]. Layered [VMS^{+}22, YZZ22]. layering [HT18]. Layout [CLO^{+}23, BFV18, PETK19]. laziness [FZL18, GV19]. Lazy [DKPS23, HT18, ZXSD21]. Leak [JB23]. leaks [BGM19, GCY^{+}20, VDvG^{+}21]. leaky [JRB^{+}22]. Learning [BOW^{+}23, CKN^{+}23, FSSW21, JLO20, KM22, KM23, MDBL20, SSC^{+}17, SADC23, WWGW20, WLJ^{+}23, WCC17, BSPC19, BPP20, CR19, COHY17, CWF^{+}19, END^{+}18, KLMM22, LWNN19, PS18, SZD^{+}17, SC18, WCH^{+}19, ZBG^{+}22]. Learning-based [MDBL20]. least [MPB^{+}22]. left [OBS^{+}22]. lemma [BG18, LFY^{+}22, SSSC^{+}22, TLKC18]. lemmas [MPB^{+}22]. lenient [CKT18]. lenses [HPC18, MMF^{+}18, MFP^{+}18, MMF^{+}19]. LER [DS17]. LER-notation [DS17]. Less [KKZ23, SY19, JME20]. Leto [BGC18]. Level [MCP23, BBKO22, CT19, CVS^{+}17, GLBP22, GCS^{+}18, HPRW21, KFEJ19, Kov22, LHJ^{+}18, LBCRK22, MA18, PZR^{+}17, SGDL20, SM19, WNC^{+}19, XK21, ZRH^{+}19]. levels [BKV^{+}21]. Leveraging [AMPS19, RAT17]. lexically [BPPS20]. Lexicographic [ACN18]. libraries [DA18, GRS^{+}21, LBH^{+}18, LAGN21, MJ22]. Library [BOW^{+}23, LB23, SL23, MRX^{+}19, MNT20, WLH^{+}17, ZDDJ21]. lie [GFFS17]. life [MW20b]. Lifting [FKRS23, LPM^{+}22, SC20]. lightweight [ADN22, BSO20, SCK^{+}20]. like [COHY17, SSK22]. Limit [XTZ^{+}23]. limited [LL22]. limiting [CDG22]. Limits [EBP^{+}23, KSP22]. Linear [LHC^{+}23, LLZ^{+}22, LZR23, VPD19, WL23, AKK^{+}21, AG20, BNS20, BHP^{+}19, BMP20, FK18, GTF^{+}20, KLO^{+}22, KCBR18, LMZ19, LFY^{+}22, NDA^{+}19, QKB21, SDB19, SFH^{+}20, ZHL^{+}20, BBN^{+}18]. Linearity [CGM18, BBN^{+}18]. Linearizability [DRV23, MK23, VSC23, EE18]. Linearize [RPF^{+}23]. Linearly [SKB^{+}22]. link [DA18, LJ20]. Linked [PNPW22]. linking [BDA20, SCK^{+}20, SKE^{+}20]. linters [HFS22]. Liquid [KWR^{+}20, PSY^{+}20, HVH20, VTV18, LPR^{+}20]. Liquidate [HVH20]. Lisp [MS20]. literal [OA18]. Live [OVCH19, YGG^{+}23, LCOC20]. LiveDroid [FZSN20]. liveness [BGM21, BMTZ21, PHL^{+}18]. living [Mai17]. LLVM [LHJ^{+}18, ZBY^{+}21]. Local [CJ17, GGV^{+}21, LLSS22, PBS^{+}17, UdM22, dVPJ20]. localization [LZ17]. localizations [WCMH19]. localizing [SSC^{+}17]. Locally [Pit23]. locations [MNT20]. Lock [BS17, Coh18, KRV19, MW20b, ZFS^{+}19]. lock-free [Coh18, MW20b, ZFS^{+}19]. Lock-step [BS17]. Locks [JB23].

Logarithm [FW22]. logging [CFL17]. Logic [DPM23, KDR23, MCP23, MK23, PT23a, PMS^{+}23, UTGK23, ZDS23, ABG^{+}17, AG20, BGHT22, BHL20, BDYG^{+}21, Cha20b, CR20, CTR^{+}20, DMS22, FRS^{+}21, GSS^{+}22, GGN^{+}21, GLSY20, HBK20, JBK22a, JLP^{+}20, KJJ^{+}18, KLMM22, LRV^{+}22, LLSS22, MP22, MJP20, MWW22, NBDF19, NGR^{+}22, O’H20, PHL^{+}18, QKB21, RLV20, RBDO22, SKE^{+}20, SC18, SYW^{+}21, SRF^{+}20, TLKC18, TGSM19, VPD19, YM21, YJY22, ZRH^{+}19, ZdAG22, dVP21]. Logical [AB23, GST^{+}20, GBTB21, KyKS22, LSSO18, NSGH22, SLO19, TSKJB18]. logics [ABG^{+}21, FNB^{+}21, KM22, MHRV20]. Logo [SHK^{+}20]. longitudinal [LWW^{+}20]. look [SHJV20]. loop [DS17, GCS^{+}18, PGIY20]. loops [KLO^{+}22, MSBK22, SSK22]. LooPy [FBP^{+}21]. loving [GM21]. low [BJC^{+}22, HKGK20, LHJ^{+}18, PZR^{+}17, SGDL20]. low-code [BJC^{+}22]. low-level [LHJ^{+}18, PZR^{+}17, SGDL20]. Lower [FCS^{+}23, GJS20, HKGK20]. LR [YNIC19]. LXM [SV21].

m [Rei17]. machine [BBTK^{+}17, BCD22, MRG18, Rot21, SVR17]. machine-code [SVR17]. machines [GTB22, IEC22, WDR18]. macro [CW20]. Macros [BKF20, CBTB20]. made [GJK^{+}21, KKT19, PLSS17, Par20]. MadMax [GKJ^{+}18]. Makam [SC18]. Making [LNO^{+}21, TLM^{+}21, XHdSO23, ACF17, JME20]. management [PVV^{+}17]. managers [CG21]. Manifest [BP17]. manipulating [MMK^{+}20, WCMH19, WGP^{+}21, ZSL^{+}22]. Manipulation [ZGHH23, MKC18]. manipulations [QSL17]. manual [PVV^{+}17]. many [HFS22, KPSL18]. map [LMM^{+}17b]. marriage [RRG^{+}21]. masses [BSO18, ML20a, Pue17]. match [BBKO22, GJS20, JGMP22, JR21]. matcher [SJSM20]. Matching [LKM23, YGG^{+}23, BGWXP22, CR20, CTR^{+}20, CA18a, Tej20, THL^{+}20, ZWWT22]. math [LAGN21]. math.h [LSA18]. Mathematical [GMS^{+}23, BGK^{+}22]. MATLAB [ML20b]. matter [MTDC19]. means [AG20]. Measurable [EPT18]. measure [Jac21]. measurement [XLD20, ZR21]. mechanical [ZdSOS19]. Mechanized [BBP20, GBTB21, TB19, BL22]. meets [DJKD20, KCL22]. Memory [KLV23, MGB^{+}23, SL23, ZCH23, ABM^{+}21, BMTZ21, BCG^{+}21, CFL17, Coh18, CAL18, CG21, DD22, DJKD20, DJR18, GGN^{+}21, HWC^{+}21, IDSW21, JJR20, JRB^{+}22, KSP22, KSTM20, KRB^{+}21, LGTS20, LNO^{+}21, ML21, MMK^{+}20, MJ21, MW20b, MKV22, PVV^{+}17, RV18, RMV22, UAM17, WZSK22]. Memory-Safe [MGB^{+}23, HWC^{+}21]. merge [CBA17, SDL18, ZH18]. Mergeable [KPSJ19]. merging [PNWT22, SZZ^{+}19]. Merlin [BRS18]. Message [DFLL23, Wad17]. Message-Passing [DFLL23]. messages [WCC17]. messaging [MAN17]. Meta [KM23, KHLD22, WCR19]. meta-programming [KHLD22, WCR19]. Meta-theorem [KM23]. metaprogramming [EUR^{+}17, JGMP22]. Metatheory [KLV23, FS22]. method [BCG^{+}20, JBK22a, KH21, WHZ^{+}19]. Methodology [FSX^{+}23]. methods [LF18]. microarchitectural [RH22]. Microservice [KZA^{+}23]. Migrating [CCEW18, GJK^{+}21]. migration [GTT21, PCAGG21]. Milner [MSI19]. Minimality [XTZ^{+}23]. Minimization [JW23]. mining [KNNR17]. misconfigurations [ZPZX21]. mismatches [HFS22]. missed [CKP22]. mitigating [BPB^{+}21]. mix [PT20]. Mixed [LMZ19, Chl21]. ML [MHR20, WR18]. MLstruct [PC22]. mobile [FZSN20, RFH^{+}22]. modal [BGM19, GSB19, KJJ^{+}18, OLE19, RRG^{+}21, VRC22, ZN21]. Modalities [Kav19, AB20]. modality [SGT^{+}22]. mode [KJK^{+}22, RSY21]. Model [KLV23, MKV22, MPB^{+}22, SVR17, UAM17, AAJN18, AAJ^{+}19, BL22, BAY20, CL21, DNFK22, EPT18, HKVR21, HKSS21, JJR20, JDKD20, KLSV18, KRV19, KKT19, MLT17, ML21, MJ21, MDSM19, RV18, SCMS20, TU22, WZSK22, YAC21]. Model-assisted [SVR17]. model-checking [TU22]. model-finding [DNFK22]. Model-guided [MPB^{+}22]. Modeling [MH21, BKPS18, JWJ^{+}21, LGTS20, SMC19, SGL^{+}21]. Models [BJP23, DFLL23, KMO23, AJRG21, BGK^{+}22, BGC18, BCG^{+}21, HHY^{+}21, KyKS22, LNO^{+}21, LMÖ^{+}22, NPWW22, PFD^{+}18, RWV19, RRG^{+}21, SSR^{+}21, UAM17, VLG21, YAY20]. modern [WCD17]. modes [BP19, LN22]. Modification [SU23]. Modular [HFCG19, KO23, NPWW22, NGR^{+}22, PETK19, PvdR23, TGSM19, UTGK23, WBM^{+}21, ZBY^{+}21, AMPS19, AMS20, CVS^{+}17, GGN^{+}21, LW22b, ŚKG18, SCK^{+}20, WCR19]. modularity [AGR^{+}20]. module [Cra19]. modules [EHAO18, HH20]. Modulo [LL23, ASD^{+}21, KSG22, MR22]. Modus [MSdSOK20]. Moebius [JGMP22]. moment [MSBK22]. monad [JDT22]. Monadic [BH22, Cho22, RBG^{+}18, SLDN17, FKLP17, TSKJB18, YZZ22]. Monads [DKPS23, AHLM20, HT18, Hir19, yKMUW22, KHLD22, MAA^{+}19, SZ21]. monitoring [EE18]. monitors [CMP20, FSK^{+}22, GFD19]. MonkeyDB [BKV^{+}21]. monomorphisation [EZYS22]. monotone [FSSW22]. monotonic [AFH^{+}18]. MOOC [CDH17]. morphisms [NBDF19]. MoSeL [KJJ^{+}18]. Mossad [DMB20]. mostly [BBTSTH17]. motion [MYZ20]. movies [ACF17]. moving [SCJG21]. MSL [JJR23]. MSWasm [MGB^{+}23]. Mtac2 [KZK^{+}18]. Much [HKSS21, MTDC19]. Multi [ADV22, CC23, RRG^{+}21, SSS^{+}23, MKV22, SEP^{+}20]. multi-actor [SEP^{+}20]. multi-execution [MKV22]. Multi-language [SSS^{+}23]. Multi-modal [RRG^{+}21]. Multi-token [CC23]. multicopy [PKSW21, PFD^{+}18]. multicopy-atomic [PFD^{+}18]. multicore [CFKP22, MJP20]. multiparadigm [VHSS20]. Multiparty [GY23, JBK22b, MYZ20, CPY20, GPP^{+}21, SY19, VHEZ21, ZFH^{+}20]. multiple [LN22, PHSR19, USM22]. multiplexing [DA18]. multiplicative [AG20]. multithreaded [BMTZ21]. must [GM21]. must-alias [GM21]. mutation [BDG^{+}22, LWY^{+}22, PWZS21]. mutations [WZS20].

name [MM19, PS18, RvAP^{+}20, SSSW21]. name-based [PS18]. Nameless [Pit23, RKV21]. names [XCIL22]. Narcissus [DSPC^{+}19]. Native [BBP23, EC22]. Natural [QSL17, BGPT19, CPB^{+}22, GDM20, LMP18, YWDD17]. Near [WLJ^{+}23, KNJ^{+}22, WAA22]. Near-Optimal [WLJ^{+}23]. near-zero [WAA22]. near-zero-cost [KNJ^{+}22]. nearly [SFH^{+}20]. Necessity [MEND22]. need [BBBK17, Bie22, BCD22, HH19]. needa [BG18]. negation [BMP20]. negative [BGHT22, BPB^{+}21, CS21, DNFK22]. nested [SSK22, WYFA20, ZA22]. nested-parallel [WYFA20]. NetKAT [VS20]. nets [Kes22]. Network [WLJ^{+}23, BGMW20, GSW21, MMS^{+}22, WWGW20]. networked [SEP^{+}20]. networks [LWNN19, USM22, UCWZ20, WAPJ22]. Neural [DAY20, JPR22, BLF^{+}19, LWNN19, MCF^{+}22, MMS^{+}22, USM22, UCWZ20, WWGW20, WAPJ22]. neural-backed [BLF^{+}19]. neural-guided [MCF^{+}22]. Neurosymbolic [BJC^{+}22, SEG^{+}22]. never [KMP19]. Newly [GM21]. Newly-single [GM21]. next [MHRV20, PA19b, CNH20]. night [BPPS20]. NK [VS20]. No [DMS17, KHLD22]. No-brainer [DMS17]. nominal [MT21, Rot21, WZSK22]. nominally [MT17]. Non [KCBR18, LC23, CFL17, CAL18, FF20, KPH22, LAF^{+}20, LMZ19, PVV^{+}17, RMV22, SYW^{+}21, UST18]. non-blocking [FF20, PVV^{+}17]. non-crashing [SYW^{+}21]. non-deterministic [UST18]. Non-linear [KCBR18, LMZ19]. Non-recursive [LC23]. non-strict [KPH22]. non-temporal [RMV22]. non-termination [LAF^{+}20]. non-volatile [CFL17, CAL18]. Nondeterminism [AB23]. Nondeterministic [CHH^{+}23, WZS19]. Noninterference [HP23, AB19, GBTB21, ZR21]. Normalization [AVW17, VRC22, BP18, ETG19]. notation [DS17, OA18]. novice [SSC^{+}17]. nullable [MvdP21]. number [LYSM22, SV21]. Numbers [SV23]. numerical [CN22, SPV18].

O [SJL19]. Object [CAL18, ZGHH23, BS21, BL22, FG18, JO22, KLL20, KEW^{+}20, LLB^{+}20, LX19, RKS18, RKHL17, SGD17, VMS^{+}22, WXWS20]. object-based [VMS^{+}22]. Object-Oriented [ZGHH23, CAL18, FG18, JO22, KEW^{+}20]. object-sensitive [LX19]. object-sensitivity [RKS18]. Objective [CNH20]. Objective-C [CNH20]. Objects [ZGHH23, BGWXP22, EJMmH17, GAGG^{+}18, LyXK^{+}22, LF18, WHZ^{+}19]. Oblivious [YD22, DSLH20]. observability [AC20]. Observational [PT22, PT23b, ML21]. Obsidian [CAMS20]. OCaml [BRS18, CDH17, LJ20, MJP20, SDW^{+}20]. occurrence [CLNL22]. Once [RPF^{+}23, WSH^{+}19]. One [BCB21, KPSL18, LN22]. Online [GAGG^{+}18, DLC22, TBB^{+}19]. Only [RPF^{+}23, BBTSTH17]. onto [SDW^{+}20]. op [NGR^{+}22, WMM20]. op-based [NGR^{+}22, WMM20]. Operational [GMS^{+}23, SL23, HH18, Hir19, KKSL22, Mai17, PFD^{+}18]. Operations [XHdSO23, EWR19, MHN19]. operator [WZS20]. opportunities [VN18]. optics [BG18]. Optimal [AAJN18, AAJ^{+}19, CCP18b, GTU23, KSP22, MPV21, SFRM20, WLJ^{+}23, BP18, KMGV22]. optimisation [LJ20]. Optimised [LRY23]. Optimization [BSRC19, CWG^{+}22, FMTSL20, HHT20, LBCRK22, MV20, MOP21, SZD^{+}21, VCD^{+}21]. optimizations [BPB^{+}21, CKP22, FSY^{+}18, GSS^{+}22, HLK^{+}20, LHJ^{+}18, LP20, PLA21, PP20, SHdM20, SSH^{+}19]. optimize [RAT17]. optimized [MA18]. optimizer [HRH^{+}21]. optimizing [BCG^{+}20, Mai17, OUM18]. Oracle [JKXH23, ZGS^{+}22]. Oracle-free [ZGS^{+}22]. Oracle-Guided [JKXH23]. oracles [LVH^{+}22]. Orca [CFD^{+}17]. Order [DFLL23, GMS^{+}23, FAM23, JB23, JJR23, MHLG23, PT23a, PvdR23, ABG^{+}17, ABG^{+}21, AK20, BWB^{+}21, BBN^{+}18, BOR18, CCP^{+}18a, CPT19, CGM18, DK20, EPT18, EHAO18, FS22, GM21, GGN^{+}21, GCY^{+}20, Ham17, HKVR21, HG22, HTLS22, Jab20, KFEJ19, KMGV22, KJK^{+}22, LYU^{+}22, MP17, MJ20, NDA^{+}19, NGTHV18, PHL^{+}18, PPSW20, RGGH21, RK22, ŚKV^{+}18, SPKT18, SMC21, SC18, SG21, UST18, WCD17]. order-aware [HKVR21]. Order-Theoretic [FAM23]. Oriented [ZGHH23, CAL18, FG18, FRS^{+}21, GRS^{+}21, JO22, KEW^{+}20]. Origins [BAP20, CNH20]. ornamentation [WR18]. other [MM19]. ours [JLP^{+}20]. out-of-gas [GKJ^{+}18]. out-of-order [GCY^{+}20]. out-of-thin-air [OD18]. Outcome [ZDS23]. output [DNFK22, GRS^{+}21, KNSA20]. output-oriented [GRS^{+}21]. overflow [TML^{+}22]. overhead [MV20]. Overloading [RHdSOZ23]. Overwatch [ZBG^{+}22]. Owicki [RLV20]. ownership [CAMS20]. Oz [VHSS20].

P [GS17, VS20]. P/Taint [GS17]. p4 [DAB^{+}21]. packet [LW22b]. page [PETK19]. pairs [NA18]. paradigm [BGK^{+}22]. Paradoxes [Jac21]. Parallel [ASW23, DHP18, AJRG21, AWA21, AG20, CSM^{+}17, Ell17b, FJM19, HKVR21, KVT20, MP21a, MRG18, Mul22, SLDN17, UCWZ20, WYFA20]. parallelism [KRV^{+}21, MA18, MAH18, MWA19, PJD^{+}21, RLS20, SDW^{+}20]. parallelism-preserving [PJD^{+}21]. Parallelization [DTM^{+}18, Mor19]. Parameterized [AAR20, JWJ^{+}21]. Parametric [HH18, NVD17, CHJ^{+}19, CVS^{+}17, KPSL18, PvAPV22]. Parametricity [DPP18, AJSW17, AB19, LTTD22, NJA20, TTS18, TLT19]. parse [SEG^{+}22]. parser [JKT21, WWP20]. Parsing [DA20, GP22, HBP19]. Partial [DFLL23, JME20, AC20, BP18, CCP^{+}18a, CPT19, DLNS21, KMGV22, LBH^{+}18, LF18, LX19, CDG22]. Partially [YvGK18, MW20a]. partially-invertible [MW20a]. Partially-static [YvGK18]. partition [MN18]. partitioning [RS20]. partitions [LSC^{+}22]. pass [LWY^{+}22, MMK^{+}20]. Passing [DFLL23, ABD21, EZYS22, SBO20, XL21]. past [KSSL18]. patch [XK21]. Path [DRV23, MPAG20, MV19, RL19, ZYT^{+}19]. path-dependent [RL19]. pathologies [PBC^{+}21]. pathology [FGS^{+}18]. paths [LyXW21]. Pattern [YGG^{+}23, BGWXP22, CA18a, DRS21, GJS20, JGMP22, JR21, Tej20]. pattern-match [GJS20, JR21]. patterns [SGD17, ZBG^{+}22]. Paxos [PLSS17]. payment [WZS19]. PCF [MP21b]. pearl [ACF17, BH22, Ber17, BG18, BBRM17, Cha20b, Chl21, DLNV17, DA20, DMS17, HLK^{+}20, HdSO21, KSSL18, Par20, PA19b, Per18, PH21, Ram22, SAFF^{+}17, SC18, SJSM20, SB19, UdM22, WDR18, YF18]. penultimate [WZD^{+}19]. per-path [ZYT^{+}19]. perfect [SMTH20]. Perfectly [UCWZ20]. performance [BCC^{+}18, CCW18, GF18, HLK^{+}20, LBH^{+}18, PBC^{+}21, Ser21, SNB^{+}19, WSG^{+}20, ZYB^{+}18]. Periodic [WL23]. Permchecker [CG21]. permissions [YDJD21]. PERs [PT23a]. PerSeVerE [KKRV21]. Persistence [Pue17, RV18]. Persistency [RWNV20, KL21, KKRV21, RV18, RWV19, RMV22]. Persistent [KRB^{+}21, RLV20, AAB^{+}21]. perspective [CLPS19]. perturbation [ASMS20]. Petr4 [DAB^{+}21]. phases [UAM17]. physicist [KH21]. pi [DdVMY19]. pi-calculus [DdVMY19]. pipeline [LW22b, RK22]. pipelines [AAM^{+}21, HKVR21, MV20, Spi17]. Pirouette [HG22]. Pixel [SADC23]. place [Tit22]. placement [PGFP17]. plagiarism [DMB20]. PlanAlyzer [TBB^{+}19]. plane [ECK^{+}22]. planes [BGMW20, DAB^{+}21, GSW21]. platform [CVS^{+}17]. platforms [PP22]. Plausible [LTTD22]. play [BS17]. Plotkin [BM20]. point [Ada19, LAGN21, PJD^{+}21, ZZX^{+}20, ZGS^{+}22]. Pointer [MW20b, BDCN20, JLO20, LSM^{+}22, LTMS18, LH22, LX19, MP21a, SHdM20, TLM^{+}21]. pointer-based [SHdM20]. Pointers [ZCH23]. pointful [PJD^{+}21]. points [BS21, GS17, JJO18, JJCO17, RKS18]. points-to [BS21, GS17, JJO18, JJCO17, RKS18]. Policy [HP23, PWSD20]. polling [BC18]. Polyadic [MPV18]. polyhedral [CL21, YAC21]. Polymorphic [MvdP20, PHSR19, BBN^{+}18, ISI17, STI20, WHZ^{+}19, ZdSOS19]. Polymorphism [FAM23, BSO20, DK19, HH18, OUM18, ST21, ZM17]. Polynomial [GHMM23, HP23, LN22]. Pomsets [JJR20]. Ponens [MSdSOK20]. pools [BMTZ22]. portable [SSH^{+}19]. positive [DFD21, DNFK22]. POSIX [GB20]. possible [BCRA18]. Possibly [FCS^{+}23]. post [LJ20, ZK22]. post-link [LJ20]. Power [DGGM23, FKLP17, KPP21, MKV18, TLM^{+}21]. powered [CWG^{+}22]. PPI [CNH20]. PPI/Stepstone [CNH20]. Practical [AMT17, EC22, BBN^{+}18, CT19, Ham17, RSY21, SPV18, WWC17, XCB^{+}18, ZYT^{+}19]. pre [ABH^{+}21, RRG^{+}21, VLG21]. pre-expectation [ABH^{+}21]. pre-trained [RRG^{+}21, VLG21]. Precise [EWR19, GPP^{+}21, JJO18, LGTS20, XLD20, CVG^{+}17, MMS^{+}22, SCJG21, SGL^{+}21, SAB17, SCZW20, TLM^{+}21, WCH^{+}19]. Precision [LTMS18, LX19, LWUD21]. Precision-guided [LTMS18]. Precision-preserving [LX19]. preconditions [JJR20]. predicate [SB19, YHW20]. prediction [KP18a, KFM21, MPV21, Mul22, Pav20]. predictive [GRXB19, MKV18]. preemptive [LRS^{+}20]. Prefix [LS23]. Prefix-Guided [LS23]. Presburger [PUW^{+}21]. presence [BCG^{+}20, TSKJB18]. presentations [yKMUW22]. Preserving [DPM23, LVSZ23, BBG^{+}20, BCRA18, FZSN20, LX19, MPV21, PJD^{+}21, SFH22]. pretty [Ber17]. preventing [ZPG^{+}17]. PRIMA [MMS^{+}22]. Primal [UTGK23, PWK^{+}22]. Primal-Dual [UTGK23, PWK^{+}22]. principal [Par20, PC22]. Principality [DR19]. principled [BL22, WR18]. principles [DNFK22]. printer [Ber17]. printf [Ada19]. priorities [CFLH^{+}22, MAH18]. priority [LSC^{+}22]. privacy [ADN22, AH18, BCK^{+}21, NDA^{+}19, SKS^{+}19, WCHRP17, ZRH^{+}19, ZRH^{+}20]. privacy-aware [SKS^{+}19]. private [SA19, ZHL^{+}20]. Probabilistic [AB23, BZSL19, CMcS23, DWH23, FCS^{+}23, LRY23, LHSM23, LVSZ23, SSSW21, YTO23, ACN18, ABG^{+}21, ABH^{+}21, ADDN17, ABM^{+}21, AYB^{+}22, AMS20, BEG^{+}18, BHL20, BKKM21, DK20, EPT18, GSW21, HKGK20, HVM20, HFCG19, Jac21, LYRY20, LCTS^{+}20, MSBK22, NcS17, NPWW22, OMN^{+}18, VS20, WCGC18, WFC^{+}20, ZA22]. probabilistically [DSLH20]. probabilities [Jac21]. probability [JSXH20, SFRM20]. problem [KNNR17]. Problems [BGC23, DZ23, BBRM17, LY18]. ProbNV [GSW21]. processes [KMP19, XCB^{+}18]. processing [DLC22, KNSA20, KM20, SLDN17, SFVJ19]. processor [ZR21]. processors [SBM20]. produce [LN22]. products [WMM20]. Profile [HMP^{+}22]. profiler [BVG19]. profiles [PJP^{+}18]. profiling [MOP21, ZHL^{+}20]. profunctor [BG18]. Program [CRMA20, GJJ^{+}20, KTST23, LW22a, LCT^{+}23, LCOC20, MRX^{+}19, SV23, WDS18, WL23, XTZ^{+}23, YSHZ21, ABG^{+}21, ADDN17, ACM^{+}18, BLF^{+}19, BLSS21, CDG22, COHY17, DG22a, FBP^{+}21, GSS^{+}22, GBR^{+}20, GRS^{+}21, HKGK20, HC21, FW22, INN21, KGdV^{+}22, Lee21, LA21, LBCRK22, MHRV20, MCF^{+}22, MWW22, MV20, NAJ^{+}20, OLE19, RLV20, RRG^{+}21, SWYZ22, SDL18, SFH22, SBEV18, WWD18, WWGW20, WCR19, ZYT^{+}19]. program-counter [HC21]. programmable [BGC18, LCTS^{+}20, OA18]. Programmatic [ZGHH23]. Programmers [BJP23, AMP^{+}20, LC21]. Programming [AJRG21, AC20, CGL^{+}23, CMcS23, DKPS23, DWZ20, LZR23, LB23, PGIY20, SWT18, ZGHH23, AP20, AWA21, AYB^{+}22, BGM19, BGM21, BGK^{+}22, BBKO22, BAP20, BBRM17, CT19, CDH17, CHJ^{+}19, CS21, CFKP22, CKS22, DPQS18, DMSM18, EPT18, ECK^{+}22, EHAO18, EGT22, EC22, FG18, FRS^{+}21, GYK^{+}20, GHMM20, GTF^{+}20, HP19, HHY^{+}21, Jac21, JSXH20, JKL^{+}22, JJKD18, KHLD22, Kin20, KFEJ19, KH18, KCL22, KEW^{+}20, LSSO18, LBH^{+}18, LRGC19, LW22b, ML20a, MRG18, MCP21, MDSM19, NWP^{+}18, OVCH19, OLC^{+}18, PJD^{+}21, PN17, Per18, PZR^{+}17, SSSW21, SMM17, SEP^{+}20, ŚKG18, SNJ^{+}19, SFVJ19, SHSO22, SZD^{+}21, SMC21, SLO19, SM19, SC18, VLG21, VMA19, VHEZ21, WZD^{+}19, WCR19, Yal17]. Programs [CHH^{+}23, DTSLT23, FKP23, FCS^{+}23, GHMM23, LHC^{+}23, LC23, LRY23, LHSM23, ZBS^{+}23, ACN18, ABG^{+}17, AZMT18, AG20, AMS20, BvGKJ17, BNS20, BKPS18, BWB^{+}21, BEG^{+}18, BHY^{+}20, BKKM21, BMTZ21, CR19, CSD^{+}21, Cha20b, CRMA20, CSM^{+}17, DK20, FJM19, GDBK^{+}21, GLPS^{+}20, HHT20, HH20, HWC^{+}21, HVM20, HFCG19, Jab20, JO22, KNSA20, KMLD20, LRV^{+}22, LLSS22, LYRY20, LZ17, LY18, LZY^{+}20, MYZ20, MHN19, MMK^{+}20, MRdAP18, MNB^{+}22, MNT20, MJ20, NBDF19, NcS17, NGTHV18, QGG19, RLV20, RMZ^{+}20, RSPC17, RFH^{+}22, SC20, SA19, SFH^{+}20, SLHR22, SRF^{+}20, TB19, UST18, WCMH19, WXWS20, WFC^{+}20, WZSK22, WGP^{+}21, WYFA20, yXZH^{+}20, YMDW21, YJY22, YMC22, ZA22, ZSL^{+}22, ZGS^{+}22]. Progress [LF18, SSR^{+}21]. Project [PVV^{+}17]. Projection [GY23, LZY^{+}20, NA18]. Projection-based [LZY^{+}20]. Prolog [SC18]. Promises [GJK^{+}21, AZMT18, MLT17]. Proof [KMO23, LCT^{+}23, MK23, USM22, BL22, CTR^{+}20, FBG20, FRS^{+}21, GCST19, Kes22, MMKK18, RKHL17, TU22]. proof-irrelevance [GCST19]. proof-nets [Kes22]. proof-oriented [FRS^{+}21]. Proof-Relevant [KMO23]. Proofs [GSF^{+}23, GTU23, KMO23, AH18, AAC^{+}18, AG20, BHY^{+}20, BP21, FV20, KPE18, KJJ^{+}18, KPP21, LAF^{+}20, LMP18, SSSC^{+}22]. propagation [Lee21, SFH22]. properties [AKK^{+}21, ASMS20, BZSL19, BMTZ21, BGG^{+}20, OCDR17, TGSM19]. Property [FSSW22, JPR22, LHP19]. Property-directed [FSSW22]. prophecy [JLP^{+}20]. Propositional [EGT22]. Propositions [RC21, FDvdHP22]. propositions-as-sessions [FDvdHP22]. Propositions-as-types [RC21]. protection [Bie22]. Proto [FKRS23]. Proto-Quipper [FKRS23]. Protocol [GY23]. Protocols [GSF^{+}23, BSRC19, CPY20, CFL17, FSD21, FSK^{+}22, Mai17, PLSS17, SWT18, ZFH^{+}20]. Prototyping [AHM^{+}17, SC18]. provable [GGV^{+}21]. Provably [AWA21, KJK^{+}22, PBS^{+}17, QSL17]. provably-correct [QSL17]. prove [Ham17]. Provenance [RMZ^{+}20, BDCN20]. Provenance-guided [RMZ^{+}20]. Proving [BEG^{+}18, DFD22, FKE^{+}20, WFC^{+}20, JBK22a, LSA18, SWT18, SM19, TSKJB18]. pruning [MKLR20]. pseudorandom [SV21]. PSO [BCG^{+}21]. PSPACE [AAR20]. PSPACE-complete [AAR20]. pure [SHdM20]. purely [UdM22]. Purity [JDT22, CK20, YMC22]. push [BC18, KMLD20]. pushdown [JKT21]. Pushing [XTZ^{+}23].

qualified [SKB^{+}22]. qualitative [LMÖ^{+}22]. quality [GS22]. quantifier [LMP18]. quantifiers [NVD17]. Quantitative [AGK23, GD23, OLE19, ZK22, LMÖ^{+}22]. Quantum [HHT20, HK22, JKL^{+}22, KO23, LLSS22, PP22, VLRH23, YJY22, YC22, YMC22, ZBS^{+}23, BHY^{+}20, CdV20, HRH^{+}21, KH21, LY18, LZY^{+}20, LVH^{+}22, MSRH20]. quasiquotes [PVSK18]. Qubit [SdSCQ19]. Queries [ZMSD23, CPB^{+}22, SKS^{+}19, WCB18, ZA22]. Query [WLJ^{+}23, AHM^{+}17, KM20, RS20, YWDD17]. Question [JKXH23]. queue [MJ21]. quick [SHJV20]. QuickChecking [MJK^{+}17]. QuickSilver [JWJ^{+}21]. Quipper [FKRS23]. Qunity [VLRH23]. quotient [KKA19, MMF^{+}18].

R [Cha20a, GV19, GJK^{+}21, GDBK^{+}21, TGKV20]. Race [AZMV23, AMT17, BGOS18, GRXB19, MKV18, Pav20, WCBG17]. RacerD [BGOS18]. races [MPV21]. Racket [FDD^{+}19]. Racketensis [SAFF^{+}17]. radio [Mai17]. Raising [WKH20]. RAM [KSS20]. Random [HTLS22, LBR20, MN18, OMO19, WCGC18]. Randomized [OMN^{+}18, WBdG^{+}23, SLHR22, WFC^{+}20]. randomness [GP22]. rank [DK19]. ranked [ZdSOS19]. ranking [ACN18]. rational [NBP^{+}18]. RaTT [BGM19]. Reachability [BWB^{+}21, KD23, LZR23, AAB^{+}21, CCP18b, CMS22, FSSW22, KP22, LSDZ22, LZR21, LSZ22, Mil20, SWYZ22, TGSM19]. Reactive [BC18, DTSLT23, BGM19, BGM21, DMSM18, PN17, Per18]. read [PGIY20]. reads [AAJ^{+}19, BCG^{+}21]. reads-from [AAJ^{+}19, BCG^{+}21]. Ready [BSZL^{+}18]. real [BSZL^{+}18, LRV^{+}22, LSM^{+}22, SMC19]. real-world [BSZL^{+}18, LSM^{+}22]. reasonable [ADV22, FKR20]. Reasonably [OA18, MLBTT22, PTFT19]. Reasoning [ABIK22, ADZ17, LyXW21, YW21, ZA22, ZDS23, AC20, BKKM21, BLSS21, BGWXP22, BSLBG22, CB18, CTR^{+}20, EWR19, HVH20, HBK20, KZK^{+}18, KNNJ18, KCBR18, LLSS22, MLT17, Mor19, ÖN22, OLE19, PLSS17, RLV20, SGT^{+}22, WCB18, YZZ22, YMC22, ZK22]. Rebuilding [FDD^{+}19]. Recalling [AFH^{+}18]. reclamation [Coh18, MW20b]. recommendation [LYB^{+}19]. reconciled [BCC^{+}18]. Reconciling [HSS23, LHJ^{+}18, VCD^{+}21, KRV^{+}21]. reconfigurable [ABIK22]. reconstruction [NBP^{+}18]. Record [XHdSO23]. Recovering [CK20]. recovery [CAL18]. Recurrence [KMLD20, AKK^{+}21, CLD20, INN21]. Recurrences [WL23]. Recursion [ADH^{+}23, LL23, BGM21, CGM18, WCGC18, ZA22]. recursions [SSK22]. Recursive [CHH^{+}23, CMcS23, LC23, ZZdSO23, KSG22, LMZ19, MNB^{+}22, MV19, PMD21, Ram22, RSY21, SSK17, WBW^{+}20, yXZH^{+}20, ZdSOZ20]. reducible [RMH21]. Reducing [PHL^{+}18, KSP22]. Reduction [AHLM20, XTZ^{+}23, CCP^{+}18a, CPT19, KMGV22, Mil20, Mor19]. Reductions [FV20, YAC21]. redundancy [DS17, LSDZ22]. refactoring [CWG^{+}22, SZZ^{+}19, VN18]. refactoring-aware [SZZ^{+}19]. Reference [LL22]. references [Jab20]. Refinement [CT19, CLO^{+}23, PMS^{+}23, RKS18, SCL^{+}23, VTC^{+}18, AKL19, CJ17, GJJ^{+}20, JR21, PSW21, SKE^{+}20, SNCM19, TML^{+}22, UST18, VVB22, WDS18]. refinements [LPR^{+}20, RBG^{+}18, ZFH^{+}20]. Reflection [SHTZ^{+}19, FKLP17, KCL22, VTC^{+}18]. Reflection-aware [SHTZ^{+}19]. Refunctionalization [WDR18]. Regex [CFLH^{+}22, THL^{+}20]. Regex-dependent [CFLH^{+}22]. Regions [WJS23]. regression [SHTZ^{+}19, YHW20]. Regular [HGJ20, LKM23, AKK^{+}21, PHXD19, SJSM20]. regular-expression [SJSM20]. reinforcement [CWF^{+}19]. related [CSD^{+}21]. relation [TSKJB18]. Relational [AKL^{+}23, BHY^{+}20, CWF^{+}19, GHHW18, MvdP21, QGG19, WWD18, ZWWT22, ABG^{+}17, ASMS20, BFV18, BPPS18, DG22b, FMTSL20, HPC18, MHRV20, RBG^{+}18, TB19, WCB18]. Relations [AB23, BDS23, AKL19, GST^{+}20, GBTB21, INN21, KyKS22, LPP18, MJ20]. Relatively [BKKM21, UST18]. Relativization [PT23a]. Relaxed [SL23, DJKD20, DJR18, JJR20, JRB^{+}22]. relaxed-memory [JRB^{+}22]. release [AAJN18, DD22, YM21]. release-acquire [AAJN18, YM21]. Relevant [KMO23].

Reliable [BKN19, KKCV17]. reloaded [SM19]. Relooper [Ram22]. repair [BJC^{+}22, PHXD19, SEG^{+}22, ZGS^{+}22]. ReplaceAll [CCH^{+}18]. replacement [SKR20, WYT^{+}22]. replacing [VN18]. replaying [KSSL18]. replicated [KPSJ19, LRGC19, LPR^{+}20]. replication [KESJ18]. report [ACM^{+}18, AHM^{+}17, BRS18, BS17, BSZL^{+}18, CDD^{+}19, FDD^{+}19, HP19, HTLS22, RK22, TRWS22, YW19]. Representation [KCB^{+}23, ACMZ21, LWNN19]. representations [KRV^{+}21, LN22]. Representing [CHH^{+}23, yXZH^{+}20]. reset [BBP20, WZD^{+}19]. Resolution [MSdSOK20, BXMS19, RvAP^{+}20, ZGSN17, ZH18]. Resolvable [PCHB23]. resolve [KNNR17]. resolved [EDWL21]. Resource [DWH23, HVH20, KH21, KWR^{+}20, MSRH20]. Resource-Aware [DWH23]. resourceful [SGT^{+}22]. resources [DMS22]. responsive [MWA19]. Restricting [AM17]. restriction [STI20]. results [LN22, OD18]. Rethinking [KEW^{+}20]. Retrofitting [SDW^{+}20]. Return [JO22]. reusable [KE19]. reuse [BPB^{+}21, DFS18, LL22]. reusing [CB17, HH20]. Reverse [SV23, DAY20, KJK^{+}22]. reverse-mode [KJK^{+}22]. reversible [CS21, CKS22]. revisited [Ada19, HMP^{+}22, SY19, TLT19]. Revisiting [ZdSOZ20]. reviver [MNTHV21]. revocation [GGV^{+}21]. rew [CTW21]. Rewrite [NWZ^{+}21, HLK^{+}20, LA21]. rewrites [LBCRK22]. Rewriting [GD23, HGJ20, NAJ^{+}20]. Rich [BEM^{+}21]. rig [CKS22]. right [MAH18, OBS^{+}22]. ring [SHSO22]. RNN [KSS20]. road [CKP22]. robot [RFH^{+}22]. robotics [GHMM20, MYZ20]. Robust [ASW23, BFV18, SGD17, SMC19]. Robustness [SADC23, MEND22, ML21]. role [CHJ^{+}19, WCVE19]. role-parametric [CHJ^{+}19]. roles [FG18]. root [BFSK20, WMW18]. rounded [LAGN21, LN22]. rounding [LN22]. rows [MM19]. RRB [Pue17]. RRB-vectors [Pue17]. Ruby [KFM21]. Rule [BGC23, MMKK18, NWZ^{+}21, SZD^{+}17]. rules [GS22, LA21]. runST [TSKJB18]. runtime [DdVMY19, FZSN20, LZY^{+}20]. Runtimes [BKK^{+}23, SM20]. Rust [HP22, AMPS19, AMP^{+}20, ESDH21, EBP^{+}23, JJKD18, JDKD20, LHC^{+}23, PXA^{+}21, WBM^{+}21, YDJD21]. RustBelt [DJKD20, JJKD18]. Ryū [Ada19].

S [Cha20a]. Safe

[EBP^{+}23, FSX^{+}23, KESJ18, LW22b, MGB^{+}23, VVB22, AAC^{+}18, BL22, BP18, CT19, DMSM18, HWC^{+}21, KEW^{+}20, MS19, OLC^{+}18, PBS^{+}17, PA19a, PVV^{+}17, ZM19]. Safer [PXA^{+}21, SNJ^{+}19, ESDH21]. Safety [KP18b, ZCH23, CCW18, FV20, FJM19, GTB22, JR21, MMK^{+}20, PHL^{+}18, PXA^{+}21]. same [CY19]. Samples [AZMV23]. sampling [ABM^{+}21, SFRM20, ZYT^{+}19]. sandboxing [PWSD20, SGDL20]. SAT [GXD20]. Satisfiability [MR22]. Satune [GXD20]. saturation [NWZ^{+}21, WNW^{+}21]. SAVI [EJMmH17]. Scala [GST^{+}20, KMV19]. Scalability [LWUD21]. Scalable [LFY^{+}22, SEP^{+}20, WBS^{+}22, JJO18, JR21, MMS^{+}22, SLO19]. ScalaLoci [WKS18]. scale [DFPG21, GDBK^{+}21, KMV19, LWW^{+}20, LLZ^{+}22]. Scaling [CGL^{+}23, CDH17, HVM20]. scan [Ell17b]. schedulers [LRS^{+}20, WBS^{+}22]. scheduling [AAM^{+}21, LBCRK22, Mel19, RvAP^{+}20, SBM20]. Scheme [FDD^{+}19]. schemes [BCK^{+}21, CGM18]. science [Cha20a]. Scilla [SNJ^{+}19]. scope [AAC^{+}18, BSLBG22, PKW17, ZvAV22]. scope-based [BSLBG22]. scoped [BPPS20]. Scopes [vAPRV18]. Scott [BM20, HSS23]. Scripting [WKN^{+}23]. scripts [CB17, KM17, SMTH20, WDS17]. Sealing [SHdM20, LTTD22]. Seam [PBS^{+}17]. search [JPR22, KW21, LYB^{+}19, PWK^{+}22, PKSW21, TU22]. Searching [KGdV^{+}22]. second [FS22, Ham17]. second-order [FS22, Ham17]. SecRSL [YM21]. Secure [AFF^{+}18, FSX^{+}23, HC21, VCD^{+}21]. securing [JJKD18]. Security [MHLG23, LZ18, YM21]. Segal [CK18]. Selection [BGC23, JKXH23, RAJ^{+}17, SHTZ^{+}19]. Selective [MLMD19, TLM^{+}21, WWP20]. Semantic [VLG21, JRB^{+}22, LMÖ^{+}22, PMD21, PvAPV22, WWGW20]. semantically [JDT22]. Semantics [DK20, GMS^{+}23, JKL^{+}22, KHDR21, SSS^{+}23, AAJN18, ABG^{+}21, AAC^{+}18, BBP20, BGJ^{+}21, CY19, Chl21, CEEW21, DLNS21, FBG20, GJK^{+}21, GLPS^{+}20, GB20, HT18, HC21, Hir19, IDSW21, KMP19, KKRV21, KKSL22, LCTS^{+}20, Mai17, Mel19, MRdAP18, RV18, RWV19, RWNV20, RMV22, SSSW21, SMC21, SB19, VMS^{+}22, ZBY^{+}21, ZA22]. semantics-aware [FBG20]. Semantics-guided [KHDR21]. Semi [AYB^{+}22, GBR^{+}20, SHSO22, CK18]. semi-ring [SHSO22]. Semi-Segal [CK18]. semi-supervised [GBR^{+}20]. Semi-symbolic [AYB^{+}22]. semicolon [JRB^{+}22]. semidirect [WMM20]. Seminaïve [AK20]. Semiring [LP20]. semistructured [CBA17]. sensitive [LH22, LX19, PP20, SWYZ22, SZ21]. sensitivity [ABH^{+}21, BEG^{+}18, JO22, JJCO17, LTMS18, LX19, RKS18, TLM^{+}21, WFC^{+}20]. separate [WXWS20]. separating [LLSS22, YDJD21]. Separation [Cha20b, DPM23, MCP23, MK23, PMS^{+}23, BWB^{+}21, BGHT22, BHL20, BDYG^{+}21, DMS22, FNB^{+}21, FRS^{+}21, GSS^{+}22, GGN^{+}21, HBK20, JBK22a, JLP^{+}20, KJJ^{+}18, LLSS22, MP22, MJP20, NBDF19, NGR^{+}22, RBDO22, SKE^{+}20, SRF^{+}20, TLKC18, TGSM19, VPD19, YM21, dVP21].

Separation-Logic [PMS^{+}23].

separation-logic-verified [VPD19]. Seq [SNB^{+}19]. Seq2Parse [SEG^{+}22]. sequences [AKK^{+}21, ZBG^{+}22]. Sequential [LRGC19, AAJ^{+}19, Cha20b, DFPG21]. sequentialization [BvGKJ17, FJM19]. serializable [SEP^{+}20]. Series [ASW23, KM20]. server [BRS18, LMM17a, QKB21]. Serverless [KZA^{+}23, AFF^{+}18, BGJ^{+}21, JPBG19]. services [BBB^{+}17, BC18, WCD17]. Session [DWH23, GY23, BP17, CY19, CPY20, CHJ^{+}19, DHP18, FLMD19, HBK20, ITVW17, JBK22b, SY19, TV20, VHEZ21]. session-type [HBK20]. sessions [CP22, FDvdHP22, GPP^{+}21, MP17, QKB21]. set [BSZL^{+}18, THL^{+}20]. Sets [Pit23, BS21, BPB^{+}21, PJD^{+}21, ZFS^{+}19]. seven [BBRM17]. SFI [KNJ^{+}22]. shader [DELT17]. shadow [CN22]. shadows [GDBK^{+}21]. Shannon [HSS23]. share [DA18]. shared [BMTZ21, IDSW21, RC21]. shared-memory [BMTZ21]. ShareJIT [XCB^{+}18]. sharing [BP17, BDG^{+}22, EJMmH17, WHZ^{+}19, XCB^{+}18]. SHARP [LH22]. shell [GB20]. shielding [VREV19]. shift [WZD^{+}19]. shift/reset [WZD^{+}19]. shifting [LKS^{+}20]. Shiftry [KSS20]. shot [WCH^{+}19]. should [CSM^{+}17]. side [BZKT21, CY19, GCY^{+}20]. side-channel [GCY^{+}20]. sides [CY19]. Signature [STI20]. signatures [AHLM20]. SigVM [ZBS^{+}22]. silent [ZPZX21]. Silver [CLO^{+}23]. Simple [AB19, AP20, BCD22, Ell18, JJR20, Par20, RKHL17]. Simplicitly [OBL^{+}18]. simplification [Mor19]. Simplifying [PFD^{+}18, YAC21]. Simply [BGM19, BMP20]. SimTyper [KFM21]. simulation [BS17]. simulations [NBDF19]. Simuliris [GSS^{+}22]. since [HK20]. Single [LZR23, GM21, MMK^{+}20]. single-pass [MMK^{+}20]. Single-Source-Single-Target [LZR23]. site [JO22]. situ [LSZ^{+}20]. size [SP22]. sized [AVW17, AD17, Dan18]. sized-type [AD17]. sketches [YMDW21]. sketching [LCOC20]. Skip [CB17]. Skipping [Chl21]. slicing [RKS18]. small [GTF^{+}20]. Smalltalk [Ing20]. Smalltalk-72 [Ing20]. smart [AGR^{+}20, BEM^{+}21, CGL22, CWG^{+}22, GKJ^{+}18, GLTS22, GAGG^{+}18, SNJ^{+}19, SGL^{+}21, WZS19, ZBS^{+}22]. Smoothness [LRY23]. SMT [BGC20, BGC23, MR22, PWZS21, SFH22, TCJ20, VTC^{+}18, WZS20]. SMT-based [BGC20]. snapshot [ÖN22]. snapshots [GFFS17]. snowflake [PVV^{+}17]. Soft [NGTHV18]. Software [DA18, TU22, DMB20, FF20, KRB^{+}21, SZZ^{+}19, SScWS19, ZHL^{+}20]. solid [SMC19]. Solidity [CGL22, TML^{+}22]. Solo [ADN22]. SolType [TML^{+}22]. solvability [AG22]. solved [HJL^{+}18]. Solver [BGC23, PCAGG21, RGSNT17, WLH^{+}17, dVPJ20]. solver-aided [RGSNT17, WLH^{+}17]. Solver-Based [BGC23, PCAGG21]. solvers [PWZS21, TCJ20, WZS20]. Solving [BGC23, CFLH^{+}22, GTU23, KSG22, UTGK23, WL23, BBRM17, MR22, SFH22]. Sound [BDCN20, BBTSTH17, DK19, EE18, KP18a, KE19, LHSM23, LVSZ23, MT17, SMC19, GRXB19, KFM21, LCTS^{+}20, MNTHV21, Pav20, RvAP^{+}20, YMC22]. Soundly [MGB^{+}23, SKE^{+}20]. soundness [BL22, GST^{+}20, GF18, KPE18, RKHL17]. Source [LZR23, GCS^{+}18, YMJ17]. source-level [GCS^{+}18]. Space [MCP23, ADV22, AWA21, BGM19, FKR20, GLPS^{+}20, MP22, PA19a, SCJG21, SHW^{+}20, ZH18]. space-efficient [AWA21, SCJG21]. spaces [KGdV^{+}22]. SpaceSearch [WLH^{+}17]. Sparcl [MW20a]. sparse [CKA18, CA22, GTF^{+}20, HHY^{+}21, KNNR17, SHW^{+}20]. spatial [OCDR17]. specialization [BP18, FCY^{+}20]. specialization-safe [BP18]. specific [BKF20, BGC18, JLO20, PLS^{+}19, SDdSO22]. Specification [AKL19, FSX^{+}23, GY23, MJ22, AMPS19, CVS^{+}17, GGN^{+}21, KHLD22, KNNR17, Rei17, WVdAE17, WBM^{+}21, XPL^{+}22]. Specification-guided [MJ22]. Specifications [BGM^{+}23, BNS20, BDYG^{+}21, BEM^{+}21, GDM20, HWC^{+}21, MEND22, MRX^{+}19, NPZ^{+}20, RvAP^{+}20, SZD^{+}17, SZ21, SCMS20, WCH^{+}19, ZDDJ21]. Specifying [NBDF19, SSR^{+}21, BGK^{+}22]. SpecSafe [BZKT21]. spectrum [GF18]. speculation [BBF^{+}21]. speculative [BZKT21, FSY^{+}18, VDvG^{+}21]. speed [BC18, PXA^{+}21]. Speeding [WCB18, MKLR20]. speedup [HLL20]. split [AGLK18]. splits [HdSO21]. splittable [SV21]. SPMD [RLS20, RMH21]. spotting [HFS22]. spreadsheet [BBZ18]. spreadsheets [IS18]. Spy [dVPJ20]. SQL [BCH^{+}22, ZXSD21, ZMSD23]. SQLizer [YWDD17]. Squeak [Ing20]. SSA [Lem23]. ST [JDT22]. stability [GCS^{+}18, PBC^{+}21]. Stable [MJ20, EPT18]. stack [BKN19, GTB22, SDB19]. Stacked [JDKD20]. stage [JGMP22]. Staged [Kov22, WCR19, WWP20, Yal17]. stages [SMM17]. Staging [XPL^{+}22, WBTR20]. Stainless [HVK19]. Standard [MHR20]. start [WSH^{+}19]. state [AFH^{+}18, FZSN20, IEC22, RC21, TSKJB18, VN18]. stateful [BGJ^{+}21, NGTHV18]. stateless [AAJN18, AAJ^{+}19, KLSV18, KRV19, KMGV22]. Static [EHAO18, Mul22, PGFP17, SMM17, SNCM19, ZPZX21, ADN22, BGC20, BGOS18, JTD21, LGTS20, LYSM22, MRdAP18, MOP21, PvAPV22, SHTZ^{+}19, SGL^{+}21, YvGK18, ZGSN17]. Statically [ABM^{+}21, PCHB23, ZFH^{+}20, CHJ^{+}19, DA18, LC21, NDA^{+}19, PVSK18]. statically-typed [CHJ^{+}19, LC21, PVSK18]. Steel [FRS^{+}21]. SteelCore [SRF^{+}20]. Step [AB23, BS17, GST^{+}20, SKD21]. step-by-step [GST^{+}20]. Step-Indexed [AB23, GST^{+}20]. step-indexing [SKD21]. Stepstone [CNH20]. StkTokens [SDB19]. stochastic [LYRY20]. storage [FMTSL20]. stores [LRGC19, RMV22]. strangers [CSM^{+}17]. strategies [HLK^{+}20]. Strategy [FK18]. Stratified [FKP23]. stream [KNSA20, MV20]. streaming [AYB^{+}22, KM20]. StreamQL [KM20]. streams [ABM^{+}21]. Strict [HT18, GJK^{+}21, KPH22]. String [DGGM23, HJL^{+}18, CCH^{+}18, CFLH^{+}22, TCJ20]. stripped [DAY20]. stroll [SAFF^{+}17]. Strong [SJSM20, BBBK17, BVG19, BCD22, GKMB17]. strongest [ZK22]. strongly [HKSS21]. Structural [RK22, BAY20, Hir19, JSXH20, LYB^{+}19, MS19, MT21, PC22]. Structurally [LC23]. Structure [GLR^{+}20, Coh18, QSL17]. structured [EWR19, MGT21, Ram22, ZH18]. Structures [DKPS23, CHJ^{+}19, FBP^{+}21, KSW18, MW20b, PKSW21, WCMH19, WBW^{+}20, YC22]. students [WC17]. Study [Rot21, BFSK20, CSD^{+}21, CGL22, CAMS20, GDBK^{+}21, GCS^{+}18, KMV19, LWW^{+}20, MHN19, PP22]. Style [MSBO23, BGM19, BGWXP22, ML21, SBO20, VRC22]. subclass [VN18]. Subcubic [CMS22]. subgraph [SdSCQ19]. sublanguage [YW19]. substitution [PT20]. subsumes [BM20]. subtracting [YF18]. Subtraction [XHdSO23]. subtype [WHZ^{+}19]. Subtyping [ZZdSO23, GPP^{+}21, MPAG20, MSdSOK20, MT18, NBP^{+}18, Par20, Rot21, SG21, YdSO17, ZdSOZ20]. sugar [PKW17]. suggestions [MGL^{+}19]. Super [ACF17]. superior [JO22]. supermartingales [ACN18]. superoptimization [MKLR20]. superposition [YC22]. supervised [CR19, GBR^{+}20]. superword [MA18]. sure [DFD21, HFCG19, MMKK18]. surviving [GKJ^{+}18]. swapping [SdSCQ19]. swift [BSRC19]. Symbolic [CKT18, Ell21, LB23, NcS17, SGL^{+}21, SLHR22, AYB^{+}22, BT18, GCY^{+}20, HBP19, KHLD22, KKT19, MSRH20, PPSW20, PNWT22, TLKC18, WCB18, WBTR20, YSHZ21]. symbolic-heap [TLKC18]. symmetric [BJSO20, CKS22, MMF^{+}19, PHSR19]. Symmetries [CKS22]. symptoms [BFSK20]. Synbit [YMDW21]. synchronization [FSK^{+}22, MPV21, Mel19]. synchronization-preserving [MPV21]. synchronizing [DTM^{+}18]. synchronous [BBP20, CY19, GTT21]. synchrony [DWZ20]. synergistic [MR22]. syntactic [PKW17]. syntax [ABF20, FS22, KPSL18, YNIC19]. syntaxes [AAC^{+}18]. synth [PGIY20]. Synthesis [BNS20, BGC23, BOW^{+}23, CWG^{+}22, DTSLT23, GHMM23, JKXH23, KO23, LC23, WDS17, WLJ^{+}23, ASMS20, BPP20, BLF^{+}19, BLSS21, BFV18, CB18, CPB^{+}22, FK18, FBP^{+}21, FBG20, GBR^{+}20, GRS^{+}21, GDM20, GJJ^{+}20, HPRW21, JGW^{+}20, JXXH21, JPR22, KHDR21, KCBR18, Lee21, MRX^{+}19, MCF^{+}22, MGL^{+}19, MNB^{+}22, MJ22, MPB^{+}22, NAJ^{+}20, PWSD20, QSL17, RMZ^{+}20, RRG^{+}21, SSSC^{+}22, SVR17, TLKC18, WDS18, WWD18, WYT^{+}22, WGP^{+}21, YWDD17, ZXSD21, ZGS^{+}22]. synthesis-aided [JGW^{+}20]. Synthesis-as-Rule-Selection [BGC23]. synthesis-based [CB18]. Synthesis-powered [CWG^{+}22]. synthesising [BCB21]. Synthesizing [AH18, ASD^{+}21, FSK^{+}22, GS22, KMD^{+}22, KLMM22, MMF^{+}18, MFP^{+}18, MMF^{+}19, SKR20, SZD^{+}17, SA19, END^{+}18, GXD20, LPM^{+}22, PJP^{+}18, YMDW21]. System [SU23, CEEW21, CFD^{+}17, HWC^{+}21, JGMP22, KPP21, LLB^{+}20, MAN17, NDA^{+}19, NAJ^{+}20, PWSD20, RSY21, SScWS19, SKE^{+}20, UST18, WKS18, HVK19]. systematic [PES20]. Systems [PMS^{+}23, WJS23, AB20, ABIK22, BHP^{+}19, CB18, CBTB20, CAMS20, DFPG21, DPQS18, GKMB17, HKSS21, IDSW21, JWJ^{+}21, LWUD21, LLZ^{+}22, LMÖ^{+}22, MMJ18, OMN^{+}18, OMO19, Pue17, RNDJ19, RS20, SJL19]. SyTeCi [Jab20].

t [KPP21]. tactics [KZK^{+}18]. TacTok [FBG20]. Tail [LL23]. taint [GS17, GS17]. taken [CKP22]. Taking [KCB^{+}23]. Taming [AGR^{+}20, CC20, KL21, LSDZ22, CTW21]. tandem [LZ17]. Tangents [RPF^{+}23]. Tape [BDS23]. Target [LZR23]. tasks [CSM^{+}17]. taxation [KNJ^{+}22]. Taylor [BM20]. Teaching [ACM^{+}18, HP19]. technique [SZZ^{+}19]. techniques [Dan18]. technology [CW20]. Template [GHMM23, XPL^{+}22]. Template-Based [GHMM23]. Temporal [SU23, UTGK23, ZCH23, BNS20, DHP18, GDM20, GTB22, GLSY20, LRS^{+}20, LSC^{+}22, RMV22]. temps [GTB22]. tensor [CKA18, CA22, KKC^{+}17, LWY^{+}22, LBCRK22, SHW^{+}20]. tensor-program [LBCRK22]. term [HGJ20, NAJ^{+}20]. Termination [GTU23, HH20, ACN18, CP22, DFD21, GBTB21, HFCG19, LAF^{+}20, LY18, MMKK18, SZ21, SKD21, WFC^{+}20]. termination-insensitive [GBTB21]. termination-sensitive [SZ21]. Test [SPKT18, ASD^{+}21, CPKG17, KM17, RNDJ19, RGSNT17, SHTZ^{+}19]. test-input [CPKG17]. Testing [DEO^{+}20, PN17, WBdG^{+}23, ZRH^{+}20, AJRG21, BKV^{+}21, DPQS18, DELT17, HTLS22, FW22, KNSA20, KSTM20, KM17, LHP19, LZY^{+}20, LBR20, MN18, MDBL20, OMN^{+}18, OMO19, PWZS21, SSR^{+}21, WZS20]. tests [LWW^{+}20, LZ17, SFH^{+}20, ZdAG22]. text [GLR^{+}20]. textual [ABF20]. their [AHLM20, AAC^{+}18, RSPC17]. them [DA18]. theorem [KM23]. Theorems [AJSW17, BDYG^{+}21, PA19b]. Theoretic [FAM23, KyKS22]. theoretical [BGWXP22]. theories [DLNS21, Ham17]. Theory [ASW23, KKZ23, VSC23, AÖV18, AG22, CH19, CTW21, DG22b, FSSW22, GSB19, HMdL22, Kov22, MLBTT22, NLA19, NVD17, PTFT19, RGGH21, SG21, TCJ20]. thin [OD18]. Thread [DMSM18, BMTZ22]. Thread-safe [DMSM18]. threats [TBB^{+}19]. three [SDL18, ZRH^{+}19]. three-level [ZRH^{+}19]. three-way [SDL18]. Thriving [Str20]. tiers [FLMD19]. Tight [AGLK18, MSBO23]. Time [HP23, BPP20, BBG^{+}20, EWR19, FKR20, KM20, MV20, NJA20, RAT17, SFH^{+}20, WFC^{+}20, WSH^{+}19]. timeline [LRS^{+}20]. timelines [LSC^{+}22]. timing [GCY^{+}20]. TiML [WWC17]. TLA [KKT19]. TLC [GLSY20]. Tlön [LW22a]. together [KNNJ18, NJA20]. token [CC23, SdSCQ19]. tolerance [BGC18, MN18]. Tolerant [WBdG^{+}23, MDSM19, Per18, VHEZ21]. too [DA18, HFS22]. tool [KPSL18]. toolchain [CG21, SMN^{+}18]. tools [WLH^{+}17]. Top [BOW^{+}23, Lee21, ZdAG22]. Top-Down [BOW^{+}23, Lee21]. Tower [YC22]. towers [AR18]. Trace [LCTS^{+}20, OMO19, AKL19]. tracking [BWB^{+}21]. tractable [EE18]. trained [RRG^{+}21, VLG21]. transactional [BE19, DD22, KRB^{+}21, LyXK^{+}22, RWV19]. Transactions [DJR18]. transducers [CFLH^{+}22, HJL^{+}18]. transfer [USM22]. Transfinite [SKD21]. transformation [KPSL18, ST21, SHW^{+}20, WMLK18]. Transformations [SV23, ABD21, BP21, CN22, GBR^{+}20, GCS^{+}18, LA21]. transformer [SB19]. transformers [KMD^{+}22]. Transforming [LZ17]. Transitioning [MT21]. transitions [KNJ^{+}22]. transitive [LSDZ22]. Translating [BCH^{+}22, ESDH21, EBP^{+}23]. Translation [Lem23, ZMSD23, BCRA18, CC22, GRB20, HP22, JDT22, Ram22, ZXSD21]. transparent [WMLK18]. Transpilation [WKN^{+}23, MCF^{+}22]. transport [TTS18]. Transpose [RPF^{+}23]. transprecision [GTF^{+}20, PUW^{+}21]. traversals [FKE^{+}20, KPW18, KRV^{+}21, SSK17]. tree [AM17, KRV^{+}21, SSK17, WDS17]. tree-traversals [KRV^{+}21]. TreeFuser [SSK17]. Trees [CHH^{+}23, SZ21, yXZH^{+}20]. triangle [PT20]. Truly [KMGV22, ZBS^{+}22]. trustworthy [VREV19]. TSO [AAR20, AAB^{+}21, BCG^{+}21, KL21, RV18]. tunneling [JJO18, ZM19]. Turning [BBP23]. Twist [YMC22]. Two [CY19, Kov22]. two-level [Kov22]. Type [ADH^{+}23, BBKO22, BCRA18, CPB^{+}22, KM17, LVSZ23, PBC^{+}21, SU23, WH19, XHdSO23, AÖV18, AB20, AAC^{+}18, AD17, BXMS19, BSLBG22, CT19, CC20, CLNL22, CH19, CB18, CBTB20, CVG^{+}17, CWG^{+}22, CEEW21, CFD^{+}17, CAMS20, CTW21, DPP18, EDWL21, ESCL22, GSB19, GF18, GJJ^{+}20, HGJ20, HWC^{+}21, HBK20, HFS22, JME20, KFM21, KFEJ19, Kov22, LWUD21, LLB^{+}20, MLBTT22, Mil20, MS19, MSI19, MV19, ME17, NDA^{+}19, NLA19, NVD17, OUM18, PES20, PWZS21, Par20, PC22, PSW21, PTFT19, PCAGG21, RGGH21, RvAP^{+}20, SSC^{+}17, SBF^{+}20, SG21, UST18, VTV18, VN18, WZS20, WC17, WCC17, ZM17, ZdSOS19, ZvAV22]. Type-and-Effect [SU23, LLB^{+}20]. type-aware [PWZS21, WZS20]. Type-Based [ADH^{+}23, BSLBG22, Mil20]. type-cases [CLNL22]. type-checkers [ZvAV22]. type-checking [ZvAV22]. Type-directed [CPB^{+}22, OUM18]. type-error [WCC17]. Type-guided [WH19, GJJ^{+}20]. Type-level [BBKO22, CT19, KFEJ19]. Type-Preserving [LVSZ23, BCRA18]. type-safe [CT19, MS19]. type-theory [RGGH21]. typechecking [DK19]. typeclass [LPR^{+}20]. Typed [MSBO23, YGG^{+}23, BMP20, CHJ^{+}19, CSD^{+}21, CDD^{+}19, ECK^{+}22, EGT22, FRS^{+}21, HG22, JDT22, KZK^{+}18, LC21, OVCH19, PVSK18, PH21, PRT^{+}18, RKV21, SM19, SRF^{+}20, Tej20, VMA19, XPL^{+}22, vdRPR^{+}22]. Types [DWH23, GY23, LHC^{+}23, PT23a, PMS^{+}23, RHdSOZ23, XHdSO23, AVW17, ADV22, AJSW17, AMPS19, BP17, BWB^{+}21, BBKO22, BCRA18, CCEW18, CK18, CL17, CY19, CHJ^{+}19, CH19, CKT18, CS21, CA18b, DdVMY19, DFD21, Dan18, DHP18, DFS18, DK19, EDWL21, ETG19, FLMD19, GYK^{+}20, GFD19, HFS22, HdSO21, ITVW17, JBK22b, JGMP22, KSG22, KPSJ19, KKA19, KWR^{+}20, Kov20, LGFD21, LCTS^{+}20, LLZ^{+}22, LMZ19, LPR^{+}20, MPAG20, MvdP20, MvdP21, MPV18, MW20b, MP20, MV19, MM19, MT18, OBL^{+}18, OLE19, PC22, PMD21, RMV22, RKHL17, RL19, RC21, SY19, ST21, SKB^{+}22, SM20, TML^{+}22, TV20, TGKV20, VVB22, VMA19, WKH20, WVdAE17, WCVE19, WMW18, XLD20, YD22, ZN21, vAPRV18]. Types-to-PERs [PT23a]. TypeScript [KM17, WSG^{+}20]. typestate [CAMS20, CG21]. typing [BBTSTH17, CCW18, CC20, CL17, CLPS19, CLNL22, CSD^{+}21, CJ17, FGS^{+}18, FKT21, ISI17, MSI19, MNTHV21, MT17, MT21, RAT17, SCJG21, VHEZ21, YdSO17]. typing-related [CSD^{+}21]. typings [AGLK18].

UDF [ZXSD21]. ultimate [YW19]. unbiased [ZYT^{+}19]. Unbounded [WJS23, DMS22, GRXB19]. unchained [Pot17, UdM22]. Undecidability [HL20]. Undecidable [DZ23]. Understanding [MKTD17, BPB^{+}21, OD18]. Unfolding [CC23]. unidirectional [YMDW21]. unification [CKN^{+}23, JXXH21, MvdP20, MvdP21]. Unified [VLRH23, AB20, BBRM17, CTR^{+}20, GS17]. Unifying [AGK23, NPZ^{+}20, PVSK18, YdSO17, ZDS23, RGGH21, ZM17]. unimodular [SSK22]. unimodular-like [SSK22]. uninitialized [GGV^{+}21]. uninterpreted [SFH^{+}20]. union [CL17, CLNL22, HdSO21, MT18, WMW18]. unique [vdRS22]. uniqueness [SKB^{+}22]. UniRec [SSK22]. units [XLD20]. univalence [ACMZ21, VMA19]. Univalent [CK18, TTS18]. universal [DPP18, WAPJ22]. Universe [FAM23, AAC^{+}18, EC22]. Unix [HKVR21]. unleashing [TLM^{+}21]. Unrealizability [KDR23]. Unsafe [MGB^{+}23, AMP^{+}20]. unsound [HFS22]. unsoundness [GFFS17]. unstructured [Ram22]. unusual [WZS20]. unwinding [BKN19]. Up-to [Dan18]. upon [DS17]. usable [CAMS20]. usage [CEEW21, FSD21, HVH20, KNNR17]. usage-aware [CEEW21]. usage-based [KNNR17]. usages [BFSK20]. use [AMP^{+}20, GV19, KSP22, KMV19, MKTD17]. User [WKN^{+}23, FKLP17, WCC17]. User-Customizable [WKN^{+}23]. user-defined [FKLP17]. users [RFH^{+}22]. using [ACM^{+}18, AFF^{+}18, ASMS20, AHM^{+}17, BDCN20, CHJ^{+}19, CWF^{+}19, CR20, CTR^{+}20, CN22, Dan18, DAY20, DEO^{+}20, GGV^{+}21, GTB22, HP19, IS18, INN21, JGMP22, JPR22, KFM21, KLMM22, LHC^{+}23, LKM23, MCF^{+}22, MSRH20, MNB^{+}22, MV20, NWZ^{+}21, SDB19, SC18, SM20, WDS17, WDS18, YMDW21, ZvAV22].

v8 [Rei17]. v8-m [Rei17]. validation [CC22, GRB20, Rei17, ŚKV^{+}18]. Validity [KTST23, TBB^{+}19]. Value [CPT19, AG22, DNFK22, FKR20, HH19, KMLD20, ST21, SGL^{+}21, SNCM19, SCZW20]. Value-centric [CPT19]. value-flow [SGL^{+}21]. value-flow-based [SCZW20]. Values [LHSM23]. variability [SC20]. variability-aware [SC20]. variable [KM22, WFC^{+}20]. variable-dependent [WFC^{+}20]. variables [CFLH^{+}22, JLP^{+}20, WCGC18]. variance [PHSR19, Rot21]. variants [LAGN21]. Variational [LRY23, JKL^{+}22, LYRY20, WMLK18]. various [UAM17]. Vector [LKM23]. vectors [Pue17]. Veracity [CFKP22]. Verification [AKL^{+}23, BGM^{+}23, DPM23, FKP23, FSX^{+}23, KTST23, SSS^{+}23, SU23, SADC23, UTGK23, WJS23, WL23, ZBS^{+}23, AAR20, ADDN17, AMPS19, BBG^{+}20, BZSL19, BKKM21, BMTZ21, BMTZ22, BEM^{+}21, BOR18, CWF^{+}19, CVS^{+}17, CMP20, EUR^{+}17, GSW21, GGN^{+}21, HKGK20, HPRW21, HP22, HKSS21, HFCG19, JWJ^{+}21, JR21, KESJ18, KKRV21, LWUD21, LLZ^{+}22, MJ21, MNTHV21, NGTHV18, NGR^{+}22, PETK19, SMN^{+}18, SFH^{+}20, SCK^{+}20, SKE^{+}20, SFH22, SGD17, TGSM19, TB19, UST18, VTC^{+}18, WBW^{+}20, WBM^{+}21, WBS^{+}22, YHW20, ZSL^{+}22, SBF^{+}20]. Verification-Preserving [DPM23]. Verified [BBP23, CLO^{+}23, CL21, KHLD22, LVH^{+}22, LBCRK22, PZR^{+}17, SDL18, WZSK22, BBF^{+}21, BBP20, GLPS^{+}20, HRH^{+}21, LPM^{+}22, LYRY20, LyXK^{+}22, MAN17, VPD19, WXWS20, ZFH^{+}20]. Verifier [LCT^{+}23, FF20, HVK19]. Verifiers [DPM23]. verify [BSZL^{+}18]. Verifying [BvGKJ17, FSD21, FJM19, GKMB17, LHC^{+}23, LPR^{+}20, ML21, NAJ^{+}20, OCDR17, ONK^{+}17, PKSW21, PMS^{+}23, WDLC18, BGC18, DD22, GSS^{+}22, GHMM20, LSM^{+}22, LRS^{+}20, LSC^{+}22, WLH^{+}17, dVPJ20]. Verilog [FMG^{+}20]. Versatile [BAS^{+}18]. version [ZH18]. versus [DPP18, RK22]. Verus [LHC^{+}23]. VESPA [MOP21]. via [BvGKJ17, BZSL19, BPPS20, CK18, CWG^{+}22, FJM19, GTU23, HC21, JJO18, JSXH20, KyKS22, LWNN19, LZR23, LBCRK22, LYB^{+}19, MSdSOK20, MNTHV21, MMS^{+}22, NPWW22, PNPW22, RKS18, RS20, SMC19, SV23, WJS23, WCMH19, WCR19, YSHZ21, ZM19, ZH18, ZZX^{+}20]. View [DFLL23, AB20]. VIP [LSM^{+}22]. Virtual [BBTK^{+}17, LRS^{+}20, MRG18, LSC^{+}22]. virtuality [EJMmH17]. Visibility [ÖN22]. visibly [JKT21]. vision [AAM^{+}21]. Visitors [Pot17]. visual [ABF20]. visualisations [PNPW22]. Visualization [WFB^{+}20]. visualizations [CPB^{+}22]. VLIW [SBM20]. VM [RAT17]. volatile [CFL17, CAL18, FZSN20, LMM17a]. volatile-by-default [LMM17a].

wait [CSM^{+}17]. warmup [BBTK^{+}17]. WATCHER [LSZ^{+}20]. way [GHHW18, HLK^{+}20, SDL18]. waypoints [PLS^{+}19]. ways [SC18]. Weak [KKZ23, KLV23, RWV19, BKV^{+}21, FKR20, KNNJ18, LNO^{+}21, MJ21, RV18]. Weakening [WRPP19]. weakly [RNDJ19]. web [AMT17, CB17, IS18, PETK19]. WebAssembly [KNJ^{+}22, Tit22, WRPP19]. WebRelate [IS18]. Weighted [BGK^{+}22, KW21]. Well [CSD^{+}21, SV23, MT17, SDB19]. well-bracketed [SDB19]. Well-Known [SV23]. Well-typed [CSD^{+}21]. were [WC17]. Whatever [COER19]. Where [MHLG23, JGMP22]. Whip [WCD17]. Who [Rei17]. whole [WCR19, ZYT^{+}19]. whole-program [WCR19, ZYT^{+}19]. wild [KMV19]. Wildcards [Bie22]. within [WCMH19]. without [AJSW17, BGM19, BDA20, COER19, FLMD19, GCST19, JLO20, KNJ^{+}22, MCL17]. witness [AFH^{+}18, Bie22]. Witnessability [DZ23]. woods [SAFF^{+}17]. work [LKS^{+}20, RSPC17]. workgroup [SSR^{+}21]. World [BCG^{+}20, BSZL^{+}18, BZKT21, LSM^{+}22, Str20]. worst [WH19]. worst-case [WH19]. write [LC21, RFH^{+}22]. wrong [CSD^{+}21].

x86 [AAB^{+}21, KL21, RLV20, RWNV20, RMV22]. x86-TSO [AAB^{+}21, KL21].

YARPGen [LBR20]. years [WBE20].

Yoneda [BG18].

Z [KTST23]. zero [DFS18, Jac21, KNJ^{+}22, WAA22, WCH^{+}19]. zero-cost [DFS18]. zippers [DA20].

References

Abdulla:2021:DRU

[AAB^{+}21] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Ahmed Bouajjani, K. Narayan Kumar, and Prakash Saivasan. Deciding reachability under persistent x86-TSO. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 56:1–56:32, January 2021. URL

                                                    https://dl.acm.org/doi/10.
                                                    1145/3434337.

Allais:2018:TSS

[AAC^{+}18] Guillaume Allais, Robert Atkey, James Chapman, Conor McBride, and James McKinna. A type and scope safe universe of syntaxes with binding: their semantics and proofs. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):90:1–90:30, July 2018.

                                                    URL https://dl.acm.org/
                                                    doi/abs/10.1145/3236785.

Abdulla:2019:OSM

[AAJ^{+}19] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Bengt Jonsson, Magnus Lång, Tuan Phong Ngo, and Konstantinos Sagonas. Optimal stateless model checking for readsfrom equivalence under sequential consistency. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):150:1–150:29, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360576.

Abdulla:2018:OSM

[AAJN18] Parosh Aziz Abdulla, Mohamed Faouzi Atig, Bengt Jonsson, and Tuan Phong Ngo. Optimal stateless model checking under the releaseacquire semantics. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):135:1–135:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276505.

Anderson:2021:EAS

[AAM^{+}21] Luke Anderson, Andrew Adams, Karima Ma, Tzu-Mao Li, Tian Jin, and Jonathan Ragan-Kelley. Efficient automatic scheduling of imaging and vision pipelines for the GPU. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):109:1–109:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485486.

Abdulla:2020:PVU

[AAR20] Parosh Aziz Abdulla, Mohamed Faouzi Atig, and Rojin Rezvan. Parameterized verification under TSO is PSPACEcomplete. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 26:1–26:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371094.

Algehed:2019:SNP

[AB19] Maximilian Algehed and Jean-Philippe Bernardy. Simple noninterference from parametricity. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP): 89:1–89:22, July 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3341693.

Abel:2020:UVM

[AB20] Andreas Abel and Jean-Philippe Bernardy. A unified view of modalities in type systems. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 90:1–90:28, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408972.

Aguirre:2023:SIL

[AB23] Alejandro Aguirre and Lars Birkedal. Step-indexed logical relations for countable nondeterminism and probabilistic choice. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 2:1–2:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571195.

Avanzini:2021:CPT

[ABD21] Martin Avanzini, Gilles Barthe, and Ugo Dal Lago. On continuation-passing transformations and expected cost analysis. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 87:1–87:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473592.

Andersen:2020:AIV

[ABF20] Leif Andersen, Michael Ballantyne, and Matthias Felleisen. Adding interactive visual syntax to textual code. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):222:1–222:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428290.

Aguirre:2017:RLH

[ABG^{+}17] Alejandro Aguirre, Gilles Barthe, Marco Gaboardi, Deepak Garg, and Pierre-Yves Strub. A relational logic for higher-order programs. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 21:1–21:??, September 2017. CODEN ???? ISSN 2475-1421.

Aguirre:2021:HOP

[ABG^{+}21] Alejandro Aguirre, Gilles Barthe, Marco Gaboardi, Deepak Garg, Shin ya Katsumata, and Tetsuya Sato. Higher-order probabilistic adversarial computations: categorical semantics and program logics. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 93:1–93:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473598.

Aguirre:2021:PEC

[ABH^{+}21] Alejandro Aguirre, Gilles Barthe, Justin Hsu, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Christoph Matheja. A preexpectation calculus for probabilistic sensitivity. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):52:1–52:28, January

                                        2021. URL https://dl.acm.
                                        org/doi/10.1145/3434333.

Ahrens:2022:RAD

[ABIK22] Emma Ahrens, Marius Bozga, Radu Iosif, and Joost-Pieter Katoen. Reasoning about distributed reconfigurable systems. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):130:1– 130:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563293.

Atkinson:2021:SBM

[ABM^{+}21] Eric Atkinson, Guillaume Baudart, Louis Mandel, Charles Yuan, and Michael Carbin. Statically bounded-memory delayed sampling for probabilistic streams. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):115:1–115:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485492.

Atkinson:2020:PRP

[AC20] Eric Atkinson and Michael Carbin. Programming and reasoning with partial observability. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):200:1– 200:28, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428268.

Andersen:2017:SLM

[ACF17] Leif Andersen, Stephen Chang, and Matthias Felleisen. Super 8 languages for making movies (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):30:1–30:??, September 2017. CODEN ???? ISSN 2475-1421.

Almeida:2018:THP

[ACM^{+}18] José Bacelar Almeida, Alcino Cunha, Nuno Macedo, Hugo Pacheco, and José Proença. Teaching how to program using automated assessment and functional glossy games (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):82:1–82:17, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236777.

Angiuli:2021:IRI

[ACMZ21] Carlo Angiuli, Evan Cavallo, Anders Mörtberg, and Max Zeuner. Internalizing representation independence with univalence. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 12:1–12:30, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434293.

Agrawal:2018:LRS

[ACN18] Sheshansh Agrawal, Krishnendu Chatterjee, and Petr Novotný. Lexicographic ranking supermartingales: an efficient approach to termination of probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 34:1–34:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Avanzini:2017:AST

[AD17] Martin Avanzini and Ugo Dal Lago. Automating sizedtype inference for complexity analysis. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 43:1–43:??, September 2017. CODEN ???? ISSN 2475-1421.

Adams:2019:RRP

[Ada19] Ulf Adams. Ryū revisited: printf floating point conversion. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):169:1– 169:23, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360595.

Albarghouthi:2017:FPV

[ADDN17] Aws Albarghouthi, Loris D’Antoni, Samuel Drews, and Aditya V. Nori. FairSquare: probabilistic verification of program fairness. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):80:1– 80:??, October 2017. CODEN ???? ISSN 2475-1421.

Abreu:2023:TBA

[ADH^{+}23] Pedro Abreu, Benjamin Delaware, Alex Hubers, Christa Jenkins, J. Garrett Morris, and Aaron Stump. A type-based approach to divide-and-conquer recursion in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 3:1–3:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571196.

Abuah:2022:SLS

[ADN22] Chiké Abuah, David Darais, and Joseph P. Near. Solo: a lightweight static analysis for differential privacy. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):150:1–150:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563313.

Accattoli:2021:EI

[ADV21] Beniamino Accattoli, Ugo Dal Lago, and Gabriele Vanoni. The (in)efficiency of interaction. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 51:1–51:33, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434332.

Accattoli:2022:MTR

[ADV22] Beniamino Accattoli, Ugo Dal Lago, and Gabriele Vanoni. Multi types and reasonable space. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 119:1–119:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547650.

Ancona:2017:RDC

[ADZ17] Davide Ancona, Francesco Dagnino, and Elena Zucca. Reasoning on divergent computations with coaxioms. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):81:1–81:??, October 2017. CODEN ???? ISSN 2475-1421.

Alpernas:2018:SSC

[AFF^{+}18] Kalev Alpernas, Cormac Flanagan, Sadjad Fouladi, Leonid Ryzhyk, Mooly Sagiv, Thomas Schmitz, and Keith Winstein. Secure serverless computing using dynamic information flow control. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):118:1–118:26, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276488.

Ahman:2018:RWF

[AFH^{+}18] Danel Ahman, Cédric Fournet, Cătălin Hriţcu, Kenji Maillard, Aseem Rastogi, and Nikhil Swamy. Recalling a witness: foundations and applications of monotonic state. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 65:1–65:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Aschieri:2020:PMP

[AG20] Federico Aschieri and Francesco A.

Genco. Par means parallel: multiplicative linear logic proofs as concurrent functional programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 18:1–18:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371086.

Accattoli:2022:TCV

[AG22] Beniamino Accattoli and Giulio Guerrieri. The theory of callby-value solvability. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):121:1–121:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547652.

Arrial:2023:QID

[AGK23] Victor Arrial, Giulio Guerrieri, and Delia Kesner. Quantitative inhabitation for different Lambda calculi in a unifying framework. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):51:1–51:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571244.

Accattoli:2018:TTS

[AGLK18] Beniamino Accattoli, Stéphane Graham-Lengrand, and Delia Kesner. Tight typings and split bounds. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):94:1–94:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236789.

Albert:2020:TCS

[AGR^{+}20] Elvira Albert, Shelly Grossman, Noam Rinetzky, Clara Rodríguez-Núñez, Albert Rubio, and Mooly Sagiv. Taming callbacks for smart contract modularity. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):209:1–209:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428277.

Albarghouthi:2018:SCP

[AH18] Aws Albarghouthi and Justin Hsu. Synthesizing coupling proofs of differential privacy. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 58:1–58:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Ahrens:2020:RMT

[AHLM20] Benedikt Ahrens, André Hirschowitz, Ambroise Lafont, and Marco Maggesi. Reduction monads and their signatures. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):31:1–31:29, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371099.

Auerbach:2017:PQC

[AHM^{+}17] Joshua S. Auerbach, Martin Hirzel, Louis Mandel, Avraham Shinnar, and Jérôme Siméon. Prototyping a query compiler using Coq (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 9:1–9:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Ahman:2018:HFA

[Ahm18] Danel Ahman. Handling fibred algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 7:1–7:??, January 2018. CO-DEN ???? ISSN 2475-1421.

AlAwar:2021:PEM

[AJRG21] Nader Al Awar, Kush Jain, Christopher J. Rossbach, and Milos Gligoric. Programming and execution models for parallel bounded exhaustive testing. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):166:1–166:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485543.

Ahmed:2017:TFF

[AJSW17] Amal Ahmed, Dustin Jamner, Jeremy G. Siek, and Philip Wadler. Theorems for free for free: parametricity, with and without types. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):39:1–39:??, September 2017. CODEN ???? ISSN 2475-1421.

Arntzenius:2020:SVE

[AK20] Michael Arntzenius and Neel Krishnaswami. Seminaïve evaluation for a higher-order functional language. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 22:1–22:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371090.

Almagor:2021:DRP

[AKK^{+}21] Shaull Almagor, Toghrul Karimov, Edon Kelmendi, Joël Ouaknine, and James Worrell. Deciding ω-regular properties on linear recurrence sequences. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 48:1–48:24, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434329.

Antonopoulos:2019:SIT

[AKL19] Timos Antonopoulos, Eric Koskinen, and Ton Chanh Le. Specification and inference of trace refinement relations. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):178:1– 178:30, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360604.

Antonopoulos:2023:AAR

[AKL^{+}23] Timos Antonopoulos, Eric Koskinen, Ton Chanh Le, Ramana Nagasamudram, David A. Naumann, and Minh Ngo. An algebra of alignment for relational verification. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):20:1–20:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571213.

Adams:2017:RGT

[AM17] Michael D. Adams and Matthew Might. Restricting grammars with tree automata. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):82:1–82:??, October 2017. CODEN ???? ISSN 2475-1421.

Astrauskas:2020:HDP

[AMP^{+}20] Vytautas Astrauskas, Christoph Matheja, Federico Poli, Peter Müller, and Alexander J. Summers. How do programmers use unsafe Rust? Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):136:1–136:27, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428204.

Astrauskas:2019:LRT

[AMPS19] Vytautas Astrauskas, Peter Müller, Federico Poli, and Alexander J. Summers. Leveraging Rust types for modular specification and verification. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):147:1– 147:30, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360573.

Avanzini:2020:MCA

[AMS20] Martin Avanzini, Georg Moser, and Michael Schaper. A modular cost analysis for probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):172:1–172:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428240.

Adamsen:2017:PIR

[AMT17] Christoffer Quist Adamsen, Anders Møller, and Frank Tip. Practical initialization race detection for JavaScript web applications. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):66:1–66:??, October 2017. CODEN ???? ISSN 2475-1421.

Abel:2018:DCT

[AÖV18] Andreas Abel, Joakim Öhman, and Andrea Vezzosi. Decidability of conversion for type theory in type theory. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 23:1–23:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Abadi:2020:SDP

[AP20] Martín Abadi and Gordon D.

Plotkin. A simple differentiable programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 38:1–38:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371106.

Ahman:2021:AE

[AP21] Danel Ahman and Matija Pretnar. Asynchronous effects. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 24:1–24:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434305.

Amin:2018:CTI

[AR18] Nada Amin and Tiark Rompf.

Collapsing towers of interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 52:1–52:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Astorga:2021:SCC

[ASD^{+}21] Angello Astorga, Shambwaditya Saha, Ahmad Dinkins, Felicia Wang, P. Madhusudan, and Tao Xie. Synthesizing contracts correct modulo a test generator. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):104:1–104:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485481.

An:2020:AEB

[ASMS20] Shengwei An, Rishabh Singh, Sasa Misailovic, and Roopsha Samanta. Augmented example-based synthesis using relational perturbation properties. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 56:1–56:24, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371124.

Alur:2023:RTS

[ASW23] Rajeev Alur, Caleb Stanford, and Christopher Watson. A robust theory of series parallel graphs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 37:1–37:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571230.

Abel:2017:NES

[AVW17] Andreas Abel, Andrea Vezzosi, and Theo Winterhalter. Normalization by evaluation for sized dependent types. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):33:1–33:??, September 2017. CODEN ???? ISSN 2475-1421.

Arora:2021:PSE

[AWA21] Jatin Arora, Sam Westrick, and Umut A. Acar. Provably spaceefficient parallel functional programming. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 18:1–18:33, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434299.

Atkinson:2022:SSI

[AYB^{+}22] Eric Atkinson, Charles Yuan, Guillaume Baudart, Louis Mandel, and Michael Carbin. Semisymbolic inference for efficient streaming probabilistic programming. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):184:1–184:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563347.

Alimadadi:2018:FBP

[AZMT18] Saba Alimadadi, Di Zhong, Magnus Madsen, and Frank Tip. Finding broken promises in asynchronous JavaScript programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):162:1– 162:26, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276532.

AlThokair:2023:DRD

[AZMV23] Mosaad Al Thokair, Minjian Zhang, Umang Mathur, and Mahesh Viswanathan. Dynamic race detection with O(1) samples. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 45:1–45:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571238.

Bright:2020:ODP

[BAP20] Walter Bright, Andrei Alexandrescu, and Michael Parker. Origins of the D programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):73:1–73:38, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386323.

Bracevac:2018:VEC

[BAS^{+}18] Oliver Bracevac, Nada Amin, Guido Salvaneschi, Sebastian Erdweg, Patrick Eugster, and Mira Mezini. Versatile event correlation with algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):67:1–67:31, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236762.

Brody:2020:SMC

[BAY20] Shaked Brody, Uri Alon, and Eran Yahav. A structural model for contextual code changes. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):215:1–215:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428283.

Bernstein:2017:GDA

[BBB^{+}17] Philip A. Bernstein, Sebastian Burckhardt, Sergey Bykov, Natacha Crooks, Jose M.

Faleiro, Gabriel Kliot, Alok Kumbhare, Muntasir Raihan Rahman, Vivek Shah, Adriana Szekeres, and Jorgen Thelin. Geo-distribution of actorbased services. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):107:1–107:??, October 2017. CODEN ???? ISSN 2475-1421.

Balabonski:2017:FSC

[BBBK17] Thibaut Balabonski, Pablo Barenbaum, Eduardo Bonelli, and Delia Kesner. Foundations of strong call by need. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):20:1–20:??, September 2017. CODEN ???? ISSN 2475-1421.

Barriere:2021:FVS

[BBF^{+}21] Aurèle Barrière, Sandrine Blazy, Olivier Flückiger, David Pichardie, and Jan Vitek. Formally verified speculation and deoptimization in a JIT compiler. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 46:1–46:26, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434327.

Barthe:2020:FVC

[BBG^{+}20] Gilles Barthe, Sandrine Blazy, Benjamin Grégoire, Rémi Hutin, Vincent Laporte, David Pichardie, and Alix Trieu. Formal verification of a constanttime preserving C compiler.

Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):7:1–7:30, January

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3371075.

Blanvillain:2022:TLP

[BBKO22] Olivier Blanvillain, Jonathan Immanuel Brachthäuser, Maxime Kjaer, and Martin Odersky. Type-level programming with match types. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 37:1–37:24, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498698.

Bernardy:2018:LHP

[BBN^{+}18] Jean-Philippe Bernardy, Mathieu Boespflug, Ryan R. Newton, Simon Peyton Jones, and Arnaud Spiwack. Linear Haskell: practical linearity in a higher-order polymorphic language. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 5:1–5:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Bourke:2020:MSV

[BBP20] Timothy Bourke, Lélio Brun, and Marc Pouzet. Mechanized semantics and verified compilation for a dataflow synchronous language with reset. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 44:1–44:29, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371112.

Barriere:2023:FVN

[BBP23] Aurèle Barrière, Sandrine Blazy, and David Pichardie. Formally verified native code generation in an effectful JIT: Turning the CompCert backend into a formally verified JIT compiler. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 9:1–9:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571202.

Byrd:2017:UAS

[BBRM17] William E. Byrd, Michael Ballantyne, Gregory Rosenblatt, and Matthew Might. A unified approach to solving seven programming problems (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 8:1–8:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Barrett:2017:VMW

[BBTK^{+}17] Edd Barrett, Carl Friedrich Bolz-Tereick, Rebecca Killick, Sarah Mount, and Laurence Tratt. Virtual machine warmup blows hot and cold. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):52:1–52:??, October 2017. CODEN ???? ISSN 2475-

Bauman:2017:SGT

[BBTSTH17] Spenser Bauman, Carl Friedrich Bolz-Tereick, Jeremy Siek, and Sam Tobin-Hochstadt. Sound gradual typing: only mostly dead. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):54:1– 54:??, October 2017. CODEN ???? ISSN 2475-1421.

Barowy:2018:EAF

[BBZ18] Daniel W. Barowy, Emery D. Berger, and Benjamin Zorn. ExceLint: automatically finding spreadsheet formula errors. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):148:1– 148:26, October 2018. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3276518.

Burckhardt:2018:RCC

[BC18] Sebastian Burckhardt and Tim Coppieters. Reactive caching for composed services: polling at the speed of push. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):152:1–152:28, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276522.

Bartha:2021:OGS

[BCB21] Sándor Bartha, James Cheney, and Vaishak Belle. One down, 699 to go: or, synthesising compositional desugarings. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):122:1–122:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485499.

Bezanson:2018:JDP

[BCC^{+}18] Jeff Bezanson, Jiahao Chen, Benjamin Chung, Stefan Karpinski, Viral B. Shah, Jan Vitek, and Lionel Zoubritzky. Julia: dynamism and performance reconciled by design. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):120:1–120:23, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276490.

Biernacka:2022:SEI

[BCD22] Małgorzata Biernacka, Witold Charatonik, and Tomasz Drab. A simple and efficient implementation of strong call by need by an abstract machine. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 94:1–94:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549822.

Barthe:2022:FCF

[BCDG22] Gilles Barthe, Raphaëlle Crubillé, Ugo Dal Lago, and Francesco Gavazzo. On Feller continuity and full abstraction. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):120:1– 120:??, August 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3547651.

Belyakova:2020:WAJ

[BCG^{+}20] Julia Belyakova, Benjamin Chung, Jack Gelinas, Jameson Nash, Ross Tate, and Jan Vitek. World age in Julia: optimizing method dispatch in the presence of eval. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):207:1–207:26, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428275.

Bui:2021:RET

[BCG^{+}21] Truc Lam Bui, Krishnendu Chatterjee, Tushar Gautam, Andreas Pavlogiannis, and Viktor Toman. The reads-from equivalence for the TSO and PSO memory models. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):164:1–164:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485541.

Benzaken:2022:TCS

[BCH^{+}22] Véronique Benzaken, Évelyne Contejean, Mohammed Houssem Hachmaoui, Chantal Keller, Louis Mandel, Avraham Shinnar, and Jérôme Siméon. Translating canonical SQL to imperative code in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):83:1–83:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527327.

Barthe:2021:DAD

[BCK^{+}21] Gilles Barthe, Rohit Chadha, Paul Krogmeier, A. Prasad Sistla, and Mahesh Viswanathan. Deciding accuracy of differential privacy schemes. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 8:1–8:30, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434289.

Bowman:2018:TPC

[BCRA18] William J. Bowman, Youyou Cong, Nick Rioux, and Amal Ahmed. Type-preserving CPS translation of Σ and Π types is not not possible. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 22:1–22:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Bartell:2020:GLD

[BDA20] Sean Bartell, Will Dietz, and Vikram S. Adve. Guided linking: dynamic linking without the costs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):145:1–145:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428213.

Banerjee:2020:SGC

[BDCN20] Subarno Banerjee, David Devecsery, Peter M. Chen, and Satish Narayanasamy. Sound garbage collection for C using pointer provenance. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):176:1–176:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428244.

Bianchini:2022:CSM

[BDG^{+}22] Riccardo Bianchini, Francesco Dagnino, Paola Giannini, Elena Zucca, and Marco Servetto. Coeffects for sharing and mutation. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):156:1– 156:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563319.

Bonchi:2023:DCR

[BDS23] Filippo Bonchi, Alessandro Di Giorgio, and Alessio Santamaria. Deconstructing the calculus of relations with tape diagrams. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 64:1–64:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571257.

Birkedal:2021:TFS

[BDYG^{+}21] Lars Birkedal, Thomas Dinsdale-Young, Armaël Guéneau, Guilhem Jaber, Kasper Svendsen, and Nikos Tzevelekos. Theorems for free from separation logic specifications. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):81:1–81:29, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3473586.

Biswas:2019:CCT

[BE19] Ranadeep Biswas and Constantin Enea. On the complexity of checking transactional consistency. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):165:1–165:28, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360591.

Barthe:2018:PES

[BEG^{+}18] Gilles Barthe, Thomas Espitau, Benjamin Grégoire, Justin Hsu, and Pierre-Yves Strub. Proving expected sensitivity of probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):57:1–57:??, January 2018. CODEN ???? ISSN 2475-1421.

Bram:2021:RSE

[BEM^{+}21] Christian Bräm, Marco Eilers, Peter Müller, Robin Sierra, and Alexander J. Summers. Rich specifications for Ethereum smart contract verification. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):146:1–146:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485523.

Bernardy:2017:PGP

[Ber17] Jean-Philippe Bernardy. A pretty but not greedy printer (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP):6:1–6:??, September 2017. CODEN ???? ISSN 2475-1421.

Bagherzadeh:2020:ACB

[BFSK20] Mehdi Bagherzadeh, Nicholas Fireman, Anas Shawesh, and Raffi Khatchadourian. Actor concurrency bugs: a comprehensive study on symptoms, root causes, API usages, and differences. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):214:1–214:32, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428282.

Bielik:2018:RRL

[BFV18] Pavol Bielik, Marc Fischer, and Martin Vechev. Robust relational layout synthesis from examples for Android. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):156:1–156:29, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276526.

Boisseau:2018:WYN

[BG18] Guillaume Boisseau and Jeremy Gibbons. What you needa know about Yoneda: profunctor optics and the Yoneda lemma (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):84:1–84:27, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236779.

Boston:2018:LVA

[BGC18] Brett Boston, Zoe Gong, and Michael Carbin. Leto: verifying application-specific hardware fault tolerance with programmable execution models. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):163:1– 163:30, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276533.

Bembenek:2020:FDS

[BGC20] Aaron Bembenek, Michael Greenberg, and Stephen Chong. Formulog: Datalog for SMTbased static analysis. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):141:1–141:31, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428209.

Bembenek:2023:SAS

[BGC23] Aaron Bembenek, Michael Greenberg, and Stephen Chong. From SMT to ASP: Solverbased approaches to solving Datalog synthesis-as-rule-selection problems. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):7:1–7:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571200.

Bruni:2020:AEP

[BGG^{+}20] Roberto Bruni, Roberto Giacobazzi, Roberta Gori, Isabel Garcia-Contreras, and Dusko Pavlovic. Abstract extensionality: on the properties of incomplete abstract interpretations. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 28:1–28:28, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371096.

Bao:2022:SLN

[BGHT22] Jialu Bao, Marco Gaboardi, Justin Hsu, and Joseph Tassarotti. A separation logic for negative dependence. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):57:1–57:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498719.

Burckhardt:2021:DFS

[BGJ^{+}21] Sebastian Burckhardt, Chris Gillum, David Justo, Konstantinos Kallas, Connor McMahon, and Christopher S. Meiklejohn. Durable functions: semantics for stateful serverless. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):133:1–133:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485510.

Batz:2022:WPP

[BGK^{+}22] Kevin Batz, Adrian Gallus, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Tobias Winkler. Weighted programming: a programming paradigm for specifying mathematical models. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):66:1–66:30, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527310.

Bahr:2019:SRF

[BGM19] Patrick Bahr, Christian Uldal Graulund, and Rasmus Ejlers Møgelberg. Simply RaTT: a fitch-style modal calculus for reactive programming without space leaks. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):109:1–109:27, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341713.

Bahr:2021:DFL

[BGM21] Patrick Bahr, Christian Uldal Graulund, and Rasmus Ejlers Møgelberg. Diamonds are not forever: liveness in reactive programming with guarded recursion. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 2:1–2:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434283.

Baumann:2023:CBV

[BGM^{+}23] Pascal Baumann, Moses Ganardi, Rupak Majumdar, Ramanathan S. Thinniyam, and Georg Zetzsche. Contextbounded verification of contextfree specifications. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):73:1–73:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571266.

Beckett:2020:AID

[BGMW20] Ryan Beckett, Aarti Gupta, Ratul Mahajan, and David Walker. Abstract interpretation of distributed network control planes. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 42:1–42:27, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371110.

Blackshear:2018:RCS

[BGOS18] Sam Blackshear, Nikos Gorogiannis, Peter W. O’Hearn, and Ilya Sergey. RacerD: compositional static race detection. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):144:1– 144:28, October 2018. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3276514.

Blanchette:2019:BBN

[BGPT19] Jasmin Christian Blanchette, Lorenzo Gheri, Andrei Popescu, and Dmitriy Traytel. Bindings as bounded natural functors. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 22:1–22:34, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290335.

Boruch-Gruszecki:2022:CDT

[BGWXP22] Aleksander Boruch-Gruszecki, Radosław Waśko, Yichen Xu, and Lionel Parreaux. A case for DOT: theoretical foundations for objects with pattern matching and GADTstyle reasoning. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):179:1–179:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563342.

Bahr:2022:MCC

[BH22] Patrick Bahr and Graham Hutton. Monadic compiler calculation (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):93:1–93:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547624.

Barthe:2020:PSL

[BHL20] Gilles Barthe, Justin Hsu, and Kevin Liao. A probabilistic separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 55:1–55:30, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371123.

Bonchi:2019:DAL

[BHP^{+}19] Filippo Bonchi, Joshua Holland, Robin Piedeleu, Paweł Sobociński, and Fabio Zanasi. Diagrammatic algebra: from linear to concurrent systems. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):25:1–25:28, January

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3290338.

Barthe:2020:RPQ

[BHY^{+}20] Gilles Barthe, Justin Hsu, Mingsheng Ying, Nengkun Yu, and Li Zhou. Relational proofs for quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 21:1–21:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371089.

Bierhoff:2022:WNW

[Bie22] Kevin Bierhoff. Wildcards need witness protection. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):138:1–138:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563301.

Bavishi:2022:NRL

[BJC^{+}22] Rohan Bavishi, Harshit Joshi, José Cambronero, Anna Fariha, Sumit Gulwani, Vu Le, Ivan Radiček, and Ashish Tiwari. Neurosymbolic repair for low-code formula languages. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):164:1– 164:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563327.

Barke:2023:GCH

[BJP23] Shraddha Barke, Michael B.

James, and Nadia Polikarpova. Grounded copilot: How programmers interact with codegenerating models. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):78:1–78:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586030.

Binder:2020:DDS

[BJSO20] David Binder, Julian Jabs, Ingo Skupin, and Klaus Ostermann. Decomposition diversity with symmetric data and codata. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 30:1–30:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371098.

Ballantyne:2020:MDS

[BKF20] Michael Ballantyne, Alexis King, and Matthias Felleisen. Macros for domain-specific languages. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):229:1– 229:29, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428297.

Batz:2023:CAE

[BKK^{+}23] Kevin Batz, Benjamin Lucien Kaminski, Joost-Pieter Katoen, Christoph Matheja, and Lena Verscht. A calculus for amortized expected runtimes. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 67:1–67:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571260.

Batz:2021:RCV

[BKKM21] Kevin Batz, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Christoph Matheja. Relatively complete verification of probabilistic programs: an expressive language for expectation-based reasoning. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 39:1–39:30, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434320.

Baldan:2019:FGC

[BKMMP19] Paolo Baldan, Barbara König, Christina Mika-Michalski, and Tommaso Padoan. Fixpoint games on continuous lattices. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 26:1–26:29, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290339.

Bastian:2019:RFD

[BKN19] Théophile Bastian, Stephen Kell, and Francesco Zappa Nardelli. Reliable and fast DWARF-based stack unwinding. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):146:1– 146:24, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360572.

Bao:2018:AMC

[BKPS18] Wenlei Bao, Sriram Krishnamoorthy, Louis-Noel Pouchet, and P. Sadayappan. Analytical modeling of cache behavior for affine programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 32:1–32:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Biswas:2021:PME

[BKV^{+}21] Ranadeep Biswas, Diptanshu Kakwani, Jyothi Vedurada, Constantin Enea, and Akash Lal. MonkeyDB: effectively testing correctness under weak isolation levels. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):132:1–132:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485546.

Blaudeau:2022:CFS

[BL22] Clément Blaudeau and Fengyun Liu. A conceptual framework for safe object initialization: a principled and mechanized soundness proof of the Celsius model. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):151:1–151:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563314.

Bavishi:2019:ANB

[BLF^{+}19] Rohan Bavishi, Caroline Lemieux, Roy Fox, Koushik Sen, and Ion Stoica. AutoPandas: neural-backed generators for program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):168:1–168:27, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360594.

Bavishi:2021:PGP

[BLSS21] Rohan Bavishi, Caroline Lemieux, Koushik Sen, and Ion Stoica. Gauss: program synthesis by reasoning over graphs. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):134:1–134:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485511.

Barbarossa:2020:TSS

[BM20] Davide Barbarossa and Giulio Manzonetto. Taylor subsumes Scott, Berry, Kahn and Plotkin. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):1:1–1:23, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371069.

Brunel:2020:BST

[BMP20] Aloïs Brunel, Damiano Mazza, and Michele Pagani. Backpropagation in the simply typed lambda-calculus with linear negation. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 64:1–64:27, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371132.

Baumann:2021:CBV

[BMTZ21] Pascal Baumann, Rupak Majumdar, Ramanathan S. Thinniyam, and Georg Zetzsche. Context-bounded verification of liveness properties for multithreaded shared-memory programs. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 44:1–44:31, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434325.

Baumann:2022:CBV

[BMTZ22] Pascal Baumann, Rupak Majumdar, Ramanathan S. Thinniyam, and Georg Zetzsche. Context-bounded verification of thread pools. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):17:1–17:28, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498678.

Bansal:2020:SCP

[BNS20] Suguman Bansal, Kedar S. Namjoshi, and Yaniv Sa’ar. Synthesis of coordination programs from linear temporal specifications. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 54:1–54:27, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371122.

Burn:2018:HOC

[BOR18] Toby Cathcart Burn, C.- H. Luke Ong, and Steven J. Ramsay. Higher-order constrained horn clauses for verification. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 11:1–11:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Bowers:2023:TSL

[BOW^{+}23] Matthew Bowers, Theo X. Olausson, Lionel Wong, Gabriel Grand, Joshua B. Tenenbaum, Kevin Ellis, and Armando Solar-Lezama. Topdown synthesis for library learning. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 41:1–41:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571234.

Balzer:2017:MSS

[BP17] Stephanie Balzer and Frank Pfenning. Manifest sharing with session types. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):37:1–37:??, September 2017. CODEN ???? ISSN 2475-1421.

Brown:2018:JOP

[BP18] Matt Brown and Jens Palsberg.

Jones-optimal partial evaluation by specialization-safe normalization. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 14:1–14:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Bender:2019:FJC

[BP19] John Bender and Jens Palsberg.

A formalization of Java’s concurrent access modes. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):142:1–142:28, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360568.

Benedikt:2021:GCT

[BP21] Michael Benedikt and Pierre Pradic. Generating collection transformations from proofs. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 14:1–14:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434295.

Brown:2021:FUM

[BPB^{+}21] Michael D. Brown, Matthew Pruett, Robert Bigelow, Girish Mururu, and Santosh Pande. Not so fast: understanding and mitigating negative impacts of compiler optimizations on code reuse gadget sets. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):154:1–154:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485531.

Barke:2020:JTL

[BPP20] Shraddha Barke, Hila Peleg, and Nadia Polikarpova. Justin-time learning for bottom-up enumerative synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):227:1–227:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428295.

Biernacki:2018:HCR

[BPPS18] Dariusz Biernacki, Maciej Piróg, Piotr Polesiuk, and Filip Sieczkowski. Handle with care: relational interpretation of algebraic effects and handlers. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 8:1–8:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Biernacki:2019:AAE

[BPPS19] Dariusz Biernacki, Maciej Piróg, Piotr Polesiuk, and Filip Sieczkowski. Abstracting algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):6:1–6:28, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290319.

Biernacki:2020:BDL

[BPPS20] Dariusz Biernacki, Maciej Piróg, Piotr Polesiuk, and Filip Sieczkowski. Binders by day, labels by night: effect instances via lexically scoped handlers. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 48:1–48:29, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371116.

Bour:2018:MLS

[BRS18] Frédéric Bour, Thomas Refis, and Gabriel Scherer. Merlin: a language server for OCaml (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):103:1–103:15, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236798.

Breitner:2017:LSS

[BS17] Joachim Breitner and Chris Smith. Lock-step simulation is child’s play (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 3:1–3:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Barbar:2021:CPS

[BS21] Mohamad Barbar and Yulei Sui. Compacting points-to sets through object clustering. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):159:1– 159:27, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485547.

Brachthauser:2022:ECB

[BSLBG22] Jonathan Immanuel Brachthäuser, Philipp Schuster, Edward Lee, and Aleksander Boruch-Gruszecki. Effects, capabilities, and boxes: from scopebased reasoning to type-based reasoning and back. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):76:1–76:30, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527320.

Brachthauser:2018:EHM

[BSO18] Jonathan Immanuel Brachthäuser, Philipp Schuster, and Klaus Ostermann. Effect handlers for the masses. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):111:1–111:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276481.

Brachthauser:2020:ECE

[BSO20] Jonathan Immanuel Brachthäuser, Philipp Schuster, and Klaus Ostermann. Effects as capabilities: effect handlers and lightweight effect polymorphism. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):126:1–126:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428194.

Bader:2019:GLF

[BSPC19] Johannes Bader, Andrew Scott, Michael Pradel, and Satish Chandra. Getafix: learning to fix bugs automatically. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):159:1–159:27, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360585.

Barik:2019:OSP

[BSRC19] Rajkishore Barik, Manu Sridharan, Murali Krishna Ramanathan, and Milind Chabbi. Optimization of swift protocols. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):164:1– 164:27, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360590.

Breitner:2018:RSV

[BSZL^{+}18] Joachim Breitner, Antal Spector-Zabusky, Yao Li, Christine Rizkallah, John Wiegley, and Stephanie Weirich. Ready, set, verify! applying hs-to-coq to real-world Haskell code (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):89:1–89:16, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236784.

Bornholt:2018:FCE

[BT18] James Bornholt and Emina Torlak. Finding code that explodes under symbolic evaluation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):149:1–149:26, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276519.

Benavides:2019:DDP

[BVG19] Zachary Benavides, Keval Vora, and Rajiv Gupta. DProf: distributed profiler with strong guarantees. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 156:1–156:24, October 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3360582.

Bakst:2017:VDP

[BvGKJ17] Alexander Bakst, Klaus v. Gleissenthall, Rami Gökhan Kici, and Ranjit Jhala. Verifying distributed programs via canonical sequentialization. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):110:1–110:??, October 2017. CODEN ???? ISSN 2475-1421.

Bao:2021:RTT

[BWB^{+}21] Yuyan Bao, Guannan Wei, Oliver Bracevac, Yuxuan Jiang, Qiyang He, and Tiark Rompf. Reachability types: tracking aliasing and separation in higher-order functional programs. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):139:1– 139:32, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485516.

Bottu:2019:CTC

[BXMS19] Gert-Jan Bottu, Ningning Xie, Koar Marntirosian, and Tom Schrijvers. Coherence of type class resolution. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):91:1–91:28, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341695.

Brotzman:2021:PSD

[BZKT21] Robert Brotzman, Danfeng Zhang, Mahmut Taylan Kandemir, and Gang Tan. SpecSafe: detecting cache side channels in a speculative world. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):129:1–129:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485506.

Bastani:2019:PVF

[BZSL19] Osbert Bastani, Xin Zhang, and Armando Solar-Lezama. Probabilistic verification of fairness properties via concentration. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):118:1– 118:27, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360544.

Cockx:2018:EDC

[CA18a] Jesper Cockx and Andreas Abel. Elaborating dependent (co)pattern matching. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):75:1–75:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236770.

Cong:2018:HDC

[CA18b] Youyou Cong and Kenichi Asai. Handling delimited continuations with dependent types. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):69:1–69:31, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236764.

Chou:2022:CDS

[CA22] Stephen Chou and Saman Amarasinghe. Compilation of dynamic sparse tensor algebra. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):175:1– 175:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563338.

Cohen:2018:OOR

[CAL18] Nachshon Cohen, David T. Aksun, and James R. Larus. Object-oriented recovery for non-volatile memory. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):153:1–153:22, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276523.

Coblenz:2020:CAT

[CAMS20] Michael Coblenz, Jonathan Aldrich, Brad A. Myers, and Joshua Sunshine. Can advanced type systems be usable? An empirical study of ownership, assets, and typestate in Obsidian. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):132:1–132:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428200.

Chasins:2017:SBR

[CB17] Sarah Chasins and Rastislav Bodik. Skip blocks: reusing execution history to accelerate web scripts. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):51:1–51:??, October 2017. CODEN ???? ISSN 2475-1421.

Chandra:2018:BSB

[CB18] Kartik Chandra and Rastislav Bodik. Bonsai: synthesisbased reasoning for type systems. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 62:1–62:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Cavalcanti:2017:EIS

[CBA17] Guilherme Cavalcanti, Paulo Borba, and Paola Accioly. Evaluating and improving semistructured merge. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):59:1–59:??, October 2017. CODEN ???? ISSN 2475-1421.

Chang:2020:DTS

[CBTB20] Stephen Chang, Michael Ballantyne, Milo Turner, and William J. Bowman. Dependent type systems as macros. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 3:1–3:29, January 2020. URL

https://dl.acm.org/doi/abs/
10.1145/3371071.

Campora:2020:TTA

[CC20] John Peter Campora and Sheng Chen. Taming type annotations in gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):191:1–191:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428259.

Clement:2022:EET

[CC22] Basile Clément and Albert Cohen. End-to-end translation validation for the Halide language. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):84:1– 84:30, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527328.

Castellan:2023:GCM

[CC23] Simon Castellan and Pierre Clairambault. The geometry of causality: Multi-token geometry of interaction and its causal unfolding. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):24:1–24:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571217.

Campora:2018:MGT

[CCEW18] John Peter Campora, Sheng Chen, Martin Erwig, and Eric Walkingshaw. Migrating gradual types. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 15:1–15:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Chen:2018:WDA

[CCH^{+}18] Taolue Chen, Yan Chen, Matthew Hague, Anthony W. Lin, and Zhilin Wu. What is decidable about string constraints with the ReplaceAll function. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 3:1–3:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Chalupa:2018:DCD

[CCP^{+}18a] Marek Chalupa, Krishnendu Chatterjee, Andreas Pavlogiannis, Nishant Sinha, and Kapil Vaidya. Data-centric dynamic partial order reduction. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 31:1–31:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Chatterjee:2018:ODR

[CCP18b] Krishnendu Chatterjee, Bhavya Choudhary, and Andreas Pavlogiannis. Optimal Dyck reachability for data-dependence and alias analysis. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 30:1–30:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Campora:2018:CCH

[CCW18] John Peter Campora, Sheng Chen, and Eric Walkingshaw. Casts and costs: harmonizing safety and performance in gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):98:1–98:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236793.

Christiansen:2019:DTH

[CDD^{+}19] David Thrane Christiansen, Iavor S. Diatchki, Robert Dockins, Joe Hendrix, and Tristan Ravitch. Dependently typed Haskell in industry (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):100:1–100:16, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341704.

Campion:2022:PCA

[CDG22] Marco Campion, Mila Dalla Preda, and Roberto Giacobazzi. Partial (In)Completeness in abstract interpretation: limiting the imprecision in program analysis. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 59:1–59:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498721.

Canou:2017:SFP

[CDH17] Benjamin Canou, Roberto Di Cosmo, and Grégoire Henry. Scaling up functional programming education: under the hood of the OCaml MOOC. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 4:1–4:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Clairambault:2020:FAQ

[CdV20] Pierre Clairambault and Marc de Visme. Full abstraction for the quantum lambdacalculus. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 63:1–63:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371131.

Choudhury:2021:GDT

[CEEW21] Pritam Choudhury, Harley Eades III, Richard A. Eisenberg, and Stephanie Weirich. A graded dependent type system with a usage-aware semantics. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 50:1–50:32, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434331.

Clebsch:2017:OGT

[CFD^{+}17] Sylvan Clebsch, Juliana Franco, Sophia Drossopoulou, Albert Mingkun Yang, Tobias Wrigstad, and Jan Vitek. Orca: GC and type system co-design for actor languages. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):72:1–72:??, October 2017. CODEN ???? ISSN 2475-1421.

Chen:2022:VDM

[CFKP22] Adam Chen, Parisa Fathololumi, Eric Koskinen, and Jared Pincus. Veracity: declarative multicore programming with commutativity. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):186:1–186:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563349.

Cohen:2017:ELN

[CFL17] Nachshon Cohen, Michal Friedman, and James R. Larus. Efficient logging in non-volatile memory by exploiting coherency protocols. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):67:1–67:??, October 2017. CODEN ???? ISSN 2475-1421.

Chen:2022:SSC

[CFLH^{+}22] Taolue Chen, Alejandro Flores-Lamas, Matthew Hague, Zhilei Han, Denghang Hu, Shuanglong Kan, Anthony W. Lin, Philipp Rümmer, and Zhilin Wu. Solving string constraints with regex-dependent functions through transducers with priorities and variables. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):45:1–45:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498707.

Cronburg:2021:PPT

[CG21] Karl Cronburg and Samuel Z.

Guyer. Permchecker: a toolchain for debugging memory managers with typestate. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):149:1– 149:28, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485526.

Chaliasos:2022:SIA

[CGL22] Stefanos Chaliasos, Arthur Gervais, and Benjamin Livshits. A study of inline assembly in Solidity smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):165:1–165:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563328.

Cambronero:2023:FSP

[CGL^{+}23] José Cambronero, Sumit Gulwani, Vu Le, Daniel Perelman, Arjun Radhakrishna, Clint Simon, and Ashish Tiwari. Flash-Fill++: Scaling programming by example by cutting to the chase. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):

33:1–33:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571226.

Clairambault:2018:LHO

[CGM18] Pierre Clairambault, Charles Grellois, and Andrzej S. Murawski. Linearity in higherorder recursion schemes. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):39:1–39:??, January 2018. CODEN ???? ISSN 2475-1421.

Cavallo:2019:HIT

[CH19] Evan Cavallo and Robert Harper. Higher inductive types in cubical computational type theory. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):1:1–1:27, January

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3290314.

Chambers:2020:RDS

[Cha20a] John M. Chambers. S, R, and data science. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):84:1–84:17, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386334.

Chargueraud:2020:SLS

[Cha20b] Arthur Charguéraud. Separation logic for sequential programs (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):116:1–116:34, August

2020. URL https://dl.acm.
org/doi/10.1145/3408998.

Chappe:2023:CTR

[CHH^{+}23] Nicolas Chappe, Paul He, Ludovic Henrio, Yannick Zakowski, and Steve Zdancewic. Choice trees: Representing nondeterministic, recursive, and impure programs in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 61:1–61:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571254.

Castro:2019:DPU

[CHJ^{+}19] David Castro, Raymond Hu, Sung-Shik Jongmans, Nicholas Ng, and Nobuko Yoshida. Distributed programming using role-parametric session types in Go: statically-typed endpoint APIs for dynamicallyinstantiated communication structures. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 29:1–29:30, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290342.

Chlipala:2021:SBB

[Chl21]  Adam Chlipala.  Skipping

the binder bureaucracy with mixed embeddings in a semantics course (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):

94:1–94:28, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473599.

Choudhury:2022:MCA

[Cho22] Pritam Choudhury. Monadic and comonadic aspects of dependency analysis. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):172:1–172:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563335.

Cosman:2017:LRT

[CJ17] Benjamin Cosman and Ranjit Jhala. Local refinement typing. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 26:1–26:??, September 2017. CODEN ???? ISSN 2475-1421.

Capriotti:2018:UHC

[CK18] Paolo Capriotti and Nicolai Kraus. Univalent higher categories via complete Semi-Segal types. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 44:1–44:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Choudhury:2020:RPC

[CK20] Vikraman Choudhury and Neel Krishnaswami. Recovering purity with comonads and capabilities. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):111:1– 111:28, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408993.

Chou:2018:FAS

[CKA18] Stephen Chou, Fredrik Kjolstad, and Saman Amarasinghe. Format abstraction for sparse tensor algebra compilers. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):123:1–123:30, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276493.

Cao:2023:BLB

[CKN^{+}23] David Cao, Rose Kunkel, Chandrakana Nandi, Max Willsey, Zachary Tatlock, and Nadia Polikarpova. babble: Learning better abstractions with E-graphs and antiunification. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 14:1–14:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571207.

Chitre:2022:RTE

[CKP22] Khushboo Chitre, Piyus Kedia, and Rahul Purandare. The road not taken: exploring alias analysis based optimizations missed by the compiler. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):153:1–153:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563316.

Choudhury:2022:SRP

[CKS22] Vikraman Choudhury, Jacek Karwowski, and Amr Sabry. Symmetries in reversible programming: from symmetric rig groupoids to reversible programming languages. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):6:1–6:32, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498667.

Chang:2018:STL

[CKT18] Stephen Chang, Alex Knauth, and Emina Torlak. Symbolic types for lenient symbolic execution. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 40:1–40:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Castagna:2017:GTU

[CL17] Giuseppe Castagna and Victor Lanvin. Gradual typing with union and intersection types. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 41:1–41:??, September 2017. CODEN ???? ISSN 2475-1421.

Courant:2021:VCG

[CL21] Nathanaël Courant and Xavier Leroy. Verified code generation for the polyhedral model. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 40:1–40:24, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434321.

Cutler:2020:DRE

[CLD20] Joseph W. Cutler, Daniel R.

Licata, and Norman Danner. Denotational recurrence extraction for amortized analysis. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 97:1–97:29, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408979.

Castagna:2022:TCU

[CLNL22] Giuseppe Castagna, Mickaël Laurent, Kim Nguy˜ ên, and Matthew Lutze. On typecases, union elimination, and occurrence typing. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):13:1–13:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498674.

Chen:2023:DSB

[CLO^{+}23] Zilin Chen, Ambroise Lafont, Liam OConnor, Gabriele Keller, Craig McLaughlin, Vincent Jackson, and Christine Rizkallah. Dargent: a silver bullet for verified data layout refinement. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):47:1–47:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571240.

Castagna:2019:GTN

[CLPS19] Giuseppe Castagna, Victor Lanvin, Tommaso Petrucciani, and Jeremy G. Siek. Gradual typing: a new perspective. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 16:1–16:32, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290329.

Chiang:2023:ERP

[CMcS23] David Chiang, Colin McDonald, and Chung chieh Shan. Exact recursive probabilistic programming. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):98:1–98:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3586050.

Clochard:2020:DVG

[CMP20] Martin Clochard, Claude Marché, and Andrei Paskevich. Deductive verification with ghost monitors. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 2:1–2:26, January 2020. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3371070.

Chistikov:2022:SCC

[CMS22] Dmitry Chistikov, Rupak Majumdar, and Philipp Schepper. Subcubic certificates for CFL reachability. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):41:1–41:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498702.

Chowdhary:2022:FSE

[CN22] Sangeeta Chowdhary and Santosh Nagarakatte. Fast shadow execution for debugging numerical errors using error free transformations. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):190:1–190:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563353.

Cox:2020:OOC

[CNH20] Brad J. Cox, Steve Naroff, and Hansen Hsu. The origins of Objective-C at PPI/Stepstone and its evolution at NeXT. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):82:1–82:74, June 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3386332.

Celik:2019:DIA

[CNRG19] Ahmet Celik, Pengyu Nie, Christopher J. Rossbach, and Milos Gligoric. Design, implementation, and application of GPU-based Java bytecode interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):177:1–177:28, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360603.

Cong:2019:CCW

[COER19] Youyou Cong, Leo Osvald, Grégory M. Essertel, and Tiark Rompf. Compiling with continuations, or without? Whatever. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):79:1–79:28, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341643.

Cohen:2018:EDS

[Coh18] Nachshon Cohen. Every data structure deserves lock-free memory reclamation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):143:1–143:24, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276513.

Chae:2017:AGF

[COHY17] Kwonsoo Chae, Hakjoo Oh, Kihong Heo, and Hongseok Yang. Automatically generating features for learning program analysis heuristics for C-like languages. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):101:1–101:??, October 2017. CODEN ???? ISSN 2475-1421.

Ciccone:2022:FTB

[CP22] Luca Ciccone and Luca Padovani. Fair termination of binary sessions. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 5:1–5:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498666.

Chen:2022:TDS

[CPB^{+}22] Qiaochu Chen, Shankara Pailoor, Celeste Barnaby, Abby Criswell, Chenglong Wang, Greg Durrett, and Işil Dillig. Typedirected synthesis of visualizations from natural language queries. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):144:1–144:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563307.

Celik:2017:BET

[CPKG17] Ahmet Celik, Sreepathi Pai, Sarfraz Khurshid, and Milos Gligoric. Bounded exhaustive test-input generation on GPUs. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):94:1– 94:??, October 2017. CODEN ???? ISSN 2475-1421.

Chatterjee:2019:VCD

[CPT19] Krishnendu Chatterjee, Andreas Pavlogiannis, and Viktor Toman. Value-centric dynamic partial order reduction. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):124:1– 124:29, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360550.

Castro-Perez:2020:CCA

[CPY20] David Castro-Perez and Nobuko Yoshida. CAMP: cost-aware multiparty session protocols. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):155:1–155:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428223.

Cambronero:2019:AAS

[CR19] José P. Cambronero and Martin C. Rinard. AL: autogenerating supervised learning programs. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):175:1–175:28, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360601.

Chen:2020:GAD

[CR20] Xiaohong Chen and Grigore Rosu. A general approach to define binders using matching logic. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 88:1–88:32, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408970.

Crary:2019:FAM

[Cra19] Karl Crary. Fully abstract module compilation. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 10:1–10:29, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290323.

Clune:2020:PEA

[CRMA20] Joshua Clune, Vijay Ramamurthy, Ruben Martins, and Umut A. Acar. Program equivalence for assisted grading of functional programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):171:1–171:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428239.

Chen:2021:CIC

[CS21] Chao-Hong Chen and Amr Sabry. A computational interpretation of compact closed categories: reversible programming with negative and fractional types. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 9:1–9:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434290.

Chaliasos:2021:WTP

[CSD^{+}21] Stefanos Chaliasos, Thodoris Sotiropoulos, Georgios-Petros Drosos, Charalambos Mitropoulos, Dimitris Mitropoulos, and Diomidis Spinellis. Well-typed programs can go wrong: a study of typing-related bugs in JVM compilers. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):123:1–123:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485500.

Cogumbreiro:2017:DAP

[CSM^{+}17] Tiago Cogumbreiro, Rishi Surendran, Francisco Martins, Vivek Sarkar, Vasco T. Vasconcelos, and Max Grossman. Deadlock avoidance in parallel programs with futures: why parallel tasks should not wait for strangers. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):103:1–103:??, October 2017. CODEN ???? ISSN 2475-1421.

Caires:2019:RKT

[CT19] Luís Caires and Bernardo Toninho. Refinement kinds: type-safe programming with practical type-level computation. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):131:1– 131:30, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360557.

Chen:2020:TUP

[CTR^{+}20] Xiaohong Chen, Minh-Thai Trinh, Nishant Rodrigues, Lucas Peña, and Grigore Rosu. Towards a unified proof framework for automated fixpoint reasoning using matching logic. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):161:1–161:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428229.

Cockx:2021:TRT

[CTW21] Jesper Cockx, Nicolas Tabareau, and Théo Winterhalter. The taming of the rew: a type theory with computational assumptions. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 60:1–60:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434341.

Chaudhuri:2017:FPT

[CVG^{+}17] Avik Chaudhuri, Panagiotis Vekris, Sam Goldman, Marshall Roch, and Gabriel Levi. Fast and precise type checking for JavaScript. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):48:1–48:??, October 2017. CODEN ???? ISSN 2475-1421.

Choi:2017:KPH

[CVS^{+}17] Joonwon Choi, Muralidaran Vijayaraghavan, Benjamin Sherman, Adam Chlipala, and Arvind. Kami: a platform for high-level parametric hardware specification and its modular verification. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 24:1–24:??, September 2017. CODEN ???? ISSN 2475-1421.

Clinger:2020:HMT

[CW20] William D. Clinger and Mitchell Wand. Hygienic macro technology. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):80:1– 80:110, June 2020. URL https:

//dl.acm.org/doi/abs/10.1145/
3386330.

Chen:2019:RVU

[CWF^{+}19] Jia Chen, Jiayi Wei, Yu Feng, Osbert Bastani, and Isil Dillig. Relational verification using reinforcement learning. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):141:1–141:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360567.

Chen:2022:SPO

[CWG^{+}22] Yanju Chen, Yuepeng Wang, Maruth Goyal, James Dong, Yu Feng, and Işil Dillig. Synthesis-powered optimization of smart contracts via data type refactoring. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):145:1–145:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563308.

Castellan:2019:TSS

[CY19] Simon Castellan and Nobuko Yoshida. Two sides of the same coin: session types and game semantics: a synchronous side and an asynchronous side. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 27:1–27:29, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290340.

Dietz:2018:SMS

[DA18] Will Dietz and Vikram Adve.

Software multiplexing: share your libraries and statically link them too. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):154:1–154:26, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276524.

Darragh:2020:PZF

[DA20] Pierce Darragh and Michael D.

Adams. Parsing with zippers (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):108:1–108:28, August

2020. URL https://dl.acm.
org/doi/10.1145/3408990.

Doenges:2021:PFF

[DAB^{+}21] Ryan Doenges, Mina Tahmasbi Arashloo, Santiago Bautista, Alexander Chang, Newton Ni, Samwise Parkinson, Rudy Peterson, Alaia Solko-Breslin, Amanda Xu, and Nate Foster. Petr4: formal foundations for p4 data planes. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 41:1–41:32, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434322.

Downen:2020:KCC

[DAJE20] Paul Downen, Zena M. Ariola, Simon Peyton Jones, and Richard A. Eisenberg. Kinds are calling conventions. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):104:1– 104:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408986.

Danielsson:2018:TUS

[Dan18] Nils Anders Danielsson. Upto techniques using sized types. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 43:1–43:??, January 2018. CO-DEN ???? ISSN 2475-1421.

David:2020:NRE

[DAY20] Yaniv David, Uri Alon, and Eran Yahav. Neural reverse engineering of stripped binaries using augmented control flow graphs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):225:1–225:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428293.

Dalvandi:2022:IVR

[DD22] Sadegh Dalvandi and Brijesh Dongol. Implementing and verifying release–acquire transactional memory in C11. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):189:1–189:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563352.

DalLago:2019:ITR

[DdVMY19] Ugo Dal Lago, Marc de Visme, Damiano Mazza, and Akira Yoshimizu. Intersection types and runtime errors in the picalculus. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 7:1–7:29, January 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3290320.

Degenbaev:2018:CCG

[DEH^{+}18] Ulan Degenbaev, Jochen Eisinger, Kentaro Hara, Marcel Hlopko, Michael Lippautz, and Hannes Payer. Cross-component garbage collection. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):151:1–151:24, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276521.

Donaldson:2017:ATG

[DELT17] Alastair F. Donaldson, Hugues Evrard, Andrei Lascu, and Paul Thomson. Automated testing of graphics shader compilers. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):93:1– 93:??, October 2017. CODEN ???? ISSN 2475-1421.

Dragoi:2020:TCI

[DEO^{+}20] Cezara Dragoi, Constantin Enea, Burcu Kulahcioglu Ozkan, Rupak Majumdar, and Filip Niksic. Testing consensus implementations using communication closure. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):210:1–210:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428278.

DalLago:2021:ITP

[DFD21] Ugo Dal Lago, Claudia Faggian, and Simona Ronchi Della Rocca. Intersection types and (positive) almost-sure termination. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 32:1–32:32, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434313.

DOsualdo:2022:PHC

[DFD22] Emanuele D’Osualdo, Azadeh Farzan, and Derek Dreyer. Proving hypersafety compositionally. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2): 135:1–135:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3563298.

DiGiusto:2023:POV

[DFLL23] Cinzia Di Giusto, Davide Ferré, Laetitia Laversa, and Etienne Lozes. A partial order view of message-passing communication models. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 55:1–55:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571248.

DePorre:2021:PEB

[DFPG21] Kevin De Porre, Carla Ferreira, Nuno Preguiça, and Elisa Gonzalez Boix. ECROs: building global scale systems from sequential code. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):107:1–107:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485484.

Diehl:2018:GZC

[DFS18] Larry Diehl, Denis Firsov, and Aaron Stump. Generic zero-cost reuse for dependent types. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):104:1–104:30, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236799.

DalLago:2022:EPD

[DG22a] Ugo Dal Lago and Francesco Gavazzo. Effectful program distancing. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 19:1–19:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498680.

DalLago:2022:RTE

[DG22b] Ugo Dal Lago and Francesco Gavazzo. A relational theory of effects and coeffects. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):31:1–31:28, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498692.

Day:2023:EPS

[DGGM23] Joel D. Day, Vijay Ganesh, Nathan Grewal, and Florin Manea. On the expressive power of string constraints. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 10:1–10:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571203.

Das:2018:PCA

[DHP18] Ankush Das, Jan Hoffmann, and Frank Pfenning. Parallel complexity analysis with temporal session types. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):91:1–91:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236786.

Dang:2020:RMR

[DJKD20] Hoang-Hai Dang, Jacques-Henri Jourdan, Jan-Oliver Kaiser, and Derek Dreyer. RustBelt meets relaxed memory. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 34:1–34:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371102.

Dongol:2018:TRM

[DJR18] Brijesh Dongol, Radha Jagadeesan, and James Riely. Transactions in relaxed memory architectures. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 18:1–18:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Dunfield:2019:SCB

[DK19] Joshua Dunfield and Neelakantan R. Krishnaswami. Sound and complete bidirectional typechecking for higher-rank polymorphism with existentials and indexed types. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):9:1–9:28, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290322.

Dahlqvist:2020:SHO

[DK20] Fredrik Dahlqvist and Dexter Kozen. Semantics of higherorder probabilistic programs with conditioning. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):57:1–57:29, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371125.

Dash:2023:AML

[DKPS23] Swaraj Dash, Younesse Kaddar, Hugo Paquet, and Sam Staton. Affine monads and lazy structures for Bayesian programming. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 46:1–46:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571239.

Dexter:2022:EOD

[DLC22] Philip Dexter, Yu David Liu, and Kenneth Chiu. The essence of online data processing. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):157:1– 157:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563320.

DiLiberti:2021:FSP

[DLNS21] Ivan Di Liberti, Fosco Loregian, Chad Nester, and Paweł Sobociński. Functorial semantics for partial theories. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 57:1–57:28, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434338.

Darais:2017:ADI

[DLNV17] David Darais, Nicholas Labich, Phúc C. Nguyen, and David Van Horn. Abstracting definitional interpreters (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 12:1–12:??, September 2017. CODEN ???? ISSN 2475-1421.

Devore-McDonald:2020:MDS

[DMB20] Breanna Devore-McDonald and Emery D. Berger. Mossad: defeating software plagiarism detection. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):138:1– 138:28, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428206.

Davis:2017:NBC

[DMS17] Milo Davis, William Meehan, and Olin Shivers. Nobrainer CPS conversion (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 23:1–23:??, September 2017. CODEN ???? ISSN 2475-1421.

Dardinier:2022:FRU

[DMS22] Thibault Dardinier, Peter Müller, and Alexander J. Summers. Fractional resources in unbounded separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):163:1–163:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563326.

Drechsler:2018:TSR

[DMSM18] Joscha Drechsler, Ragnar Mogk, Guido Salvaneschi, and Mira Mezini. Thread-safe reactive programming. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):107:1–107:30, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276477.

Dyer:2022:ACP

[DNFK22] Tristan Dyer, Tim Nelson, Kathi Fisler, and Shriram Krishnamurthi. Applying cognitive principles to model-finding output: the positive value of negative information. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):79:1–79:29, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527323.

Dardinier:2023:VPI

[DPM23] Thibault Dardinier, Gaurav Parthasarathy, and Peter Müller. Verification-preserving inlining in automatic separation logic verifiers. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):102:1–102:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586054.

Devriese:2018:PVU

[DPP18] Dominique Devriese, Marco Patrignani, and Frank Piessens. Parametricity versus the universal type. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 38:1–38:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Desai:2018:CPT

[DPQS18] Ankush Desai, Amar Phanishayee, Shaz Qadeer, and Sanjit A. Seshia. Compositional programming and testing of dynamic distributed systems. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):159:1– 159:30, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276529.

Dudenhefner:2019:PAU

[DR19] Andrej Dudenhefner and Jakob Rehof. Principality and approximation under dimensional bound. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):8:1–8:29, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290321.

Dura:2021:PJA

[DRS21] Alexandru Dura, Christoph Reichenbach, and Emma Söderberg. JavaDL: automatically incrementalizing Java bug pattern detection. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):165:1–165:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485542.

DOsualdo:2023:PDL

[DRV23] Emanuele DOsualdo, Azalea Raad, and Viktor Vafeiadis. The path to durable linearizability. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 26:1–26:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571219.

Ding:2017:GGL

[DS17] Yufei Ding and Xipeng Shen.

GLORE: generalized loop redundancy elimination upon LER-notation. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):74:1–74:??, October 2017. CODEN ???? ISSN 2475-1421.

Darais:2020:LPO

[DSLH20] David Darais, Ian Sweet, Chang Liu, and Michael Hicks. A language for probabilistically oblivious computation. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 50:1–50:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371118.

Delaware:2019:NCC

[DSPC^{+}19] Benjamin Delaware, Sorawit Suriyakarn, Clément Pit-Claudel, Qianchuan Ye, and Adam Chlipala. Narcissus: correct-by-construction derivation of decoders and encoders from binary formats. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):82:1–82:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341686.

Daloze:2018:PDL

[DTM^{+}18] Benoit Daloze, Arie Tal, Stefan Marr, Hanspeter Mössenböck, and Erez Petrank. Parallelization of dynamic languages: synchronizing builtin collections. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):108:1–108:30, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276478.

Das:2023:CFA

[DTSLT23] Ria Das, Joshua B. Tenenbaum, Armando Solar-Lezama, and Zenna Tavares. Combining functional and automata synthesis to discover causal reactive programs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):56:1–56:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571249.

deVilhena:2021:SLE

[dVP21] Paulo Emílio de Vilhena and François Pottier. A separation logic for effect handlers. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 33:1–33:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434314.

deVilhena:2020:SGV

[dVPJ20] Paulo Emílio de Vilhena, François Pottier, and Jacques-Henri Jourdan. Spy game: verifying a local generic solver in Iris. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 33:1–33:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371101.

Das:2023:PRA

[DWH23] Ankush Das, Di Wang, and Jan Hoffmann. Probabilistic resource-aware session types. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):66:1–66:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571259.

Dragoi:2020:PES

[DWZ20] Cezara Dragoi, Josef Widder, and Damien Zufferey. Programming at the edge of synchrony. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):213:1–213:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428281.

Ding:2023:WUP

[DZ23] Shuo Ding and Qirun Zhang. Witnessability of undecidable problems. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 34:1–34:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571227.

Emre:2023:ALT

[EBP^{+}23] Mehmet Emre, Peter Boyland, Aesha Parekh, Ryan Schroeder, Kyle Dewey, and Ben Hardekopf. Aliasing limits on translating C to safe Rust. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):94:1–94:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586046.

Escot:2022:PGP

[EC22] Lucas Escot and Jesper Cockx.

Practical generic programming over a universe of native datatypes. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 113:1–113:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547644.

Eichholz:2022:DTD

[ECK^{+}22] Matthias Eichholz, Eric Hayden Campbell, Matthias Krebs, Nate Foster, and Mira Mezini. Dependently-typed data plane programming. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 40:1–40:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498701.

Eisenberg:2021:ECR

[EDWL21] Richard A. Eisenberg, Guillaume Duboc, Stephanie Weirich, and Daniel Lee. An existential crisis resolved: type inference for first-class existential types. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):

64:1–64:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473569.

Emmi:2018:SCT

[EE18] Michael Emmi and Constantin Enea. Sound, complete, and tractable linearizability monitoring for concurrent collections. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 25:1–25:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Eremondi:2022:PEG

[EGT22] Joseph Eremondi, Ronald Garcia, and Éric Tanter. Propositional equality for gradual dependently typed programming. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 96:1–96:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547627.

Elsman:2018:SIH

[EHAO18] Martin Elsman, Troels Henriksen, Danil Annenkov, and Cosmin E. Oancea. Static interpretation of higher-order modules in Futhark: functional GPU programming in the large. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 97:1–97:30, July 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3236792.

ElHajj:2017:SOS

[EJMmH17] Izzat El Hajj, Thomas B.

Jablin, Dejan Milojicic, and Wen mei Hwu. SAVI objects: sharing and virtuality incorporated. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):45:1–45:??, October 2017. CODEN ???? ISSN 2475-1421.

Elliott:2017:CC

[Ell17a] Conal Elliott. Compiling to categories. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 27:1–27:??, September 2017. CODEN ???? ISSN 2475-1421.

Elliott:2017:GFP

[Ell17b] Conal Elliott. Generic functional parallel algorithms: scan and FFT. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 7:1–7:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Elliott:2018:SEA

[Ell18]  Conal Elliott.  The sim-

ple essence of automatic differentiation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):70:1–70:29, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236765.

Elliott:2021:SAD

[Ell21] Conal Elliott. Symbolic and automatic differentiation of languages. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 78:1–78:18, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473583.

Ezudheen:2018:HIL

[END^{+}18] P. Ezudheen, Daniel Neider, Deepak D’Souza, Pranav Garg, and P. Madhusudan. Horn-ICE learning for synthesizing invariants and contracts. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):131:1–131:25, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276501.

Ehrhard:2018:MCS

[EPT18] Thomas Ehrhard, Michele Pagani, and Christine Tasson. Measurable cones and stable, measurable functions: a model for probabilistic higher-order programming. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 59:1–59:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Emrich:2022:CBT

[ESCL22] Frank Emrich, Jan Stolarek, James Cheney, and Sam Lindley. Constraint-based type inference for FreezeML. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):111:1–111:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547642.

Emre:2021:TCS

[ESDH21] Mehmet Emre, Ryan Schroeder, Kyle Dewey, and Ben Hardekopf. Translating C to safer Rust. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):121:1–121:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485498.

Eremondi:2019:ANG

[ETG19] Joseph Eremondi, Éric Tanter, and Ronald Garcia. Approximate normalization for gradual dependent types. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):88:1–88:30, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341692.

Ebner:2017:MFF

[EUR^{+}17] Gabriel Ebner, Sebastian Ullrich, Jared Roesch, Jeremy Avigad, and Leonardo de Moura. A metaprogramming framework for formal verification. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):34:1–34:??, September 2017. CODEN ???? ISSN 2475-1421.

Essertel:2019:PRS

[EWR19] Grégory M. Essertel, Guannan Wei, and Tiark Rompf. Precise reasoning with structured time, structured heaps, and collective operations. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):157:1–157:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360583.

Ellis:2022:GGG

[EZYS22] Stephen Ellis, Shuofei Zhu, Nobuko Yoshida, and Linhai Song. Generic go to go: dictionary-passing, monomorphisation, and hybrid. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):168:1–168:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563331.

Hou:2023:OTA

[FAM23] Kuen-Bang Hou (Favonia), Carlo Angiuli, and Reed Mullanix. An order-theoretic analysis of universe polymorphism. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 57:1–57:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571250.

First:2020:TSA

[FBG20] Emily First, Yuriy Brun, and Arjun Guha. TacTok: semantics-aware proof synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):231:1– 231:31, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428299.

Ferdowsifard:2021:PLI

[FBP^{+}21] Kasra Ferdowsifard, Shraddha Barke, Hila Peleg, Sorin Lerner, and Nadia Polikarpova. LooPy: interactive program synthesis with control structures. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):153:1–153:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485530.

Feng:2023:LBP

[FCS^{+}23] Shenghua Feng, Mingshuai Chen, Han Su, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Naijun Zhan. Lower bounds for possibly divergent probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):99:1–99:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586051.

Fluckiger:2020:CDF

[FCY^{+}20] Olivier Flückiger, Guido Chari, Ming-Ho Yee, Jan Jecmen, Jakob Hain, and Jan Vitek. Contextual dispatch for function specialization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):220:1–220:24, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428288.

Flatt:2019:RRC

[FDD^{+}19] Matthew Flatt, Caner Derici, R. Kent Dybvig, Andrew W. Keep, Gustavo E. Massaccesi, Sarah Spall, Sam Tobin-Hochstadt, and Jon Zeppieri. Rebuilding Racket on Chez Scheme (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):78:1–78:15, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341642.

Frumin:2022:BSP

[FDvdHP22] Dan Frumin, Emanuele D’Osualdo, Bas van den Heuvel, and Jorge A. Pérez. A bunch of sessions: a propositions-as-sessions interpretation of bunched implications in channel-based concurrency. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):155:1–155:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563318.

Flanagan:2020:AVB

[FF20] Cormac Flanagan and Stephen N. Freund. The anchor verifier for blocking and non-blocking concurrent software. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):156:1–156:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428224.

Faes:2018:CAO

[FG18] Michael Faes and Thomas R.

Gross. Concurrency-aware object-oriented programming with roles. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 130:1–130:30, October 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3276500.

Feltey:2018:CCF

[FGS^{+}18] Daniel Feltey, Ben Greenman, Christophe Scholliers, Robert Bruce Findler, and Vincent St-Amour. Collapsible contracts: fixing a pathology of gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):133:1–133:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276503.

Feldman:2020:CII

[FISS20] Yotam M. Y. Feldman, Neil Immerman, Mooly Sagiv, and Sharon Shoham. Complexity and information in invariant inference. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):5:1–5:29, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371073.

Fernando:2019:VSA

[FJM19] Vimuth Fernando, Keyur Joshi, and Sasa Misailovic. Verifying safety and accuracy of approximate parallel programs via canonical sequentialization. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):119:1– 119:29, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360545.

Farzan:2018:SSL

[FK18] Azadeh Farzan and Zachary Kincaid. Strategy synthesis for linear arithmetic games. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):61:1–61:??, January 2018. CODEN ???? ISSN 2475-1421.

Feldman:2020:PHC

[FKE^{+}20] Yotam M. Y. Feldman, Artem Khyzha, Constantin Enea, Adam Morrison, Aleksandar Nanevski, Noam Rinetzky, and Sharon Shoham. Proving highly-concurrent traversals correct. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):128:1–128:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428196.

Forster:2017:EPU

[FKLP17] Yannick Forster, Ohad Kammar, Sam Lindley, and Matija Pretnar. On the expressive power of user-defined effects: effect handlers, monadic reflection, delimited control. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):13:1–13:??, September 2017. CODEN ???? ISSN 2475-1421.

Farzan:2023:SCV

[FKP23] Azadeh Farzan, Dominik Klumpp, and Andreas Podelski. Stratified commutativity in verification algorithms for concurrent programs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):49:1–49:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571242.

Forster:2020:WCV

[FKR20] Yannick Forster, Fabian Kunze, and Marc Roth. The weak call-by-value λ-calculus is reasonable for both time and space. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 27:1–27:23, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371095.

Fu:2023:PQD

[FKRS23] Peng Fu, Kohei Kishida, Neil J.

Ross, and Peter Selinger. Protoquipper with dynamic lifting. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):11:1– 11:??, January 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571204.

Fu:2021:LDL

[FKT21] Weili Fu, Fabian Krause, and Peter Thiemann. Label dependent lambda calculus and gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):108:1–108:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485485.

Fowler:2019:EAS

[FLMD19] Simon Fowler, Sam Lindley, J. Garrett Morris, and Sára Decova. Exceptional asynchronous session types: session types without tiers. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 28:1–28:29, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290341.

Flake:2020:VHA

[FMG^{+}20] Peter Flake, Phil Moorby, Steve Golson, Arturo Salz, and Simon Davidmann. Verilog HDL and its ancestors and descendants. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):87:1–87:90, June 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3386337.

Feser:2020:DOR

[FMTSL20] John Feser, Sam Madden, Nan Tang, and Armando Solar-Lezama. Deductive optimization of relational data storage. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):170:1– 170:30, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428238.

Farka:2021:AAC

[FNB^{+}21] František Farka, Aleksandar Nanevski, Anindya Banerjee, Germán Andrés Delbianco, and Ignacio Fábregas. On algebraic abstractions for concurrent separation logics. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 5:1–5:32, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434286.

Fromherz:2021:PSP

[FRS^{+}21] Aymeric Fromherz, Aseem Rastogi, Nikhil Swamy, Sydney Gibson, Guido Martínez, Denis Merigoux, and Tahina Ramananandro. Steel: prooforiented programming in a dependently typed concurrent separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 85:1–85:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473590.

Fiore:2022:FMS

[FS22] Marcelo Fiore and Dmitrij Szamozvancev. Formal metatheory of second-order abstract syntax. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 53:1–53:29, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498715.

Ferles:2021:VCU

[FSD21] Kostas Ferles, Jon Stephens, and Isil Dillig. Verifying correct usage of context-free API protocols. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 17:1–17:30, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434298.

Ferles:2022:SFG

[FSK^{+}22] Kostas Ferles, Benjamin Sepanski, Rahul Krishnan, James Bornholt, and Isil Dillig. Synthesizing fine-grained synchronization protocols for implicit monitors. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):67:1–67:26, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3527311.

Feldman:2021:LBI

[FSSW21] Yotam M. Y. Feldman, Mooly Sagiv, Sharon Shoham, and James R. Wilcox. Learning the boundary of inductive invariants. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 15:1–15:30, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434296.

Feldman:2022:PDR

[FSSW22] Yotam M. Y. Feldman, Mooly Sagiv, Sharon Shoham, and James R. Wilcox. Propertydirected reachability as abstract interpretation in the monotone theory. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 15:1–15:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498676.

Fox:2023:VMA

[FSX^{+}23] Anthony C. J. Fox, Gareth Stockwell, Shale Xiong, Hanno Becker, Dominic P. Mulligan, Gustavo Petri, and Nathan Chong. A verification methodology for the Arm confidential computing architecture: From a secure specification to safe implementations. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):88:1–88:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586040.

Fluckiger:2018:CSO

[FSY^{+}18] Olivier Flückiger, Gabriel Scherer, Ming-Ho Yee, Aviral Goel, Amal Ahmed, and Jan Vitek. Correctness of speculative optimizations with dynamic deoptimization. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):49:1–49:??, January 2018. CODEN ???? ISSN 2475-1421.

Farzan:2020:RSP

[FV20] Azadeh Farzan and Anthony Vandikas. Reductions for safety proofs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 13:1–13:28, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371081.

Hou:2022:LPT

[FW22] Kuen-Bang Hou (Favonia) and Zhuyang Wang. Logarithm and program testing. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):64:1–64:26, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498726.

Foner:2018:KYL

[FZL18] Kenneth Foner, Hengchu Zhang, and Leonidas Lampropoulos. Keep your laziness in check. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):102:1– 102:30, July 2018. URL https:

                                        //dl.acm.org/doi/abs/10.1145/
                                        3236797.

Farooq:2020:LIP

[FZSN20] Umar Farooq, Zhijia Zhao, Manu Sridharan, and Iulian Neamtiu. LiveDroid: identifying and preserving mobile app state in volatile runtime environments. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):160:1–160:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428228.

Grossman:2018:ODE

[GAGG^{+}18] Shelly Grossman, Ittai Abraham, Guy Golan-Gueta, Yan Michalevsky, Noam Rinetzky, Mooly Sagiv, and Yoni Zohar. Online detection of effectively callback free objects with applications to smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 48:1–48:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Greenberg:2020:EFS

[GB20] Michael Greenberg and Austin J.

Blatt. Executable formal semantics for the POSIX shell. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 43:1–43:30, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371111.

Gao:2020:FDS

[GBR^{+}20] Xiang Gao, Shraddha Barke, Arjun Radhakrishna, Gustavo Soares, Sumit Gulwani, Alan Leung, Nachiappan Nagappan, and Ashish Tiwari. Feedback-driven semisupervised synthesis of program transformations. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):219:1–219:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428287.

Gregersen:2021:MLR

[GBTB21] Simon Oddershede Gregersen, Johan Bay, Amin Timany, and Lars Birkedal. Mechanized logical relations for termination-insensitive noninterference. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 10:1–10:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434291.

Gong:2018:ESE

[GCS^{+}18] Zhangxiaowen Gong, Zhi Chen, Justin Szaday, David Wong, Zehra Sura, Neftali Watkinson, Saeed Maleki, David Padua, Alexander Veidenbaum, Alexandru Nicolau, and Josep Torrellas. An empirical study of the effect of source-level loop transformations on compiler stability. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):126:1–126:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276496.

Gilbert:2019:DPI

[GCST19] Gaëtan Gilbert, Jesper Cockx, Matthieu Sozeau, and Nicolas Tabareau. Definitional proofirrelevance without K. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):3:1–3:28, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290316.

Guo:2020:ECT

[GCY^{+}20] Shengjian Guo, Yueqi Chen, Jiyong Yu, Meng Wu, Zhiqiang Zuo, Peng Li, Yueqiang Cheng, and Huibo Wang. Exposing cache timing side-channel leaks through out-of-order symbolic execution. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):147:1–147:32, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428215.

Gavazzo:2023:EQR

[GD23] Francesco Gavazzo and Cecilia Di Florio. Elements of quantitative rewriting. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):63:1–63:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571256.

Goel:2021:WWE

[GDBK^{+}21] Aviral Goel, Pierre Donat-Bouillud, Filip Krikava, Christoph M. Kirsch, and Jan Vitek. What we eval in the shadows: a large-scale study of eval in R programs. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):125:1–125:23, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485502.

Gavran:2020:IST

[GDM20] Ivan Gavran, Eva Darulova, and Rupak Majumdar. Interactive synthesis of temporal specifications from examples and natural language. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):201:1–201:26, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428269.

Greenman:2018:STS

[GF18] Ben Greenman and Matthias Felleisen. A spectrum of type soundness and performance. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 71:1–71:32, July 2018. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3236766.

Greenman:2019:CMG

[GFD19] Ben Greenman, Matthias Felleisen, and Christos Dimoulas. Complete monitors for gradual types. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):122:1– 122:29, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360548.

Grech:2017:HDL

[GFFS17] Neville Grech, George Fourtounis, Adrian Francalanza, and Yannis Smaragdakis. Heaps don’t lie: countering unsoundness with heap snapshots. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):68:1– 68:??, October 2017. CODEN ???? ISSN 2475-1421.

Gondelman:2021:DCM

[GGN^{+}21] Léon Gondelman, Simon Oddershede Gregersen, Abel Nieto, Amin Timany, and Lars Birkedal. Distributed causal memory: modular specification and verification in higherorder distributed separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 42:1–42:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434323.

Georges:2021:EPL

[GGV^{+}21] Aïna Linn Georges, Armaël Guéneau, Thomas Van Strydonck, Amin Timany, Alix Trieu, Sander Huyghebaert, Dominique Devriese, and Lars Birkedal. Efficient and provable local capability revocation using uninitialized capabilities. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 6:1–6:30, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434287.

Gibbons:2018:RAW

[GHHW18] Jeremy Gibbons, Fritz Henglein, Ralf Hinze, and Nicolas Wu. Relational algebra by way of adjunctions. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):86:1–86:28, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236781.

Griesemer:2020:FG

[GHK^{+}20] Robert Griesemer, Raymond Hu, Wen Kokke, Julien Lange, Ian Lance Taylor, Bernardo Toninho, Philip Wadler, and Nobuko Yoshida. Featherweight Go. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):149:1–149:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428217.

Ghosh:2020:KLP

[GHMM20] Ritwika Ghosh, Chiao Hsieh, Sasa Misailovic, and Sayan Mitra. Koord: a language for programming and verifying distributed robotics application. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):232:1– 232:30, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428300.

Goharshady:2023:AGA

[GHMM23] Amir Kafshdar Goharshady, S. Hitarth, Fatemeh Mohammadi, and Harshit Jitendra Motwani. Algebro-geometric algorithms for template-based synthesis of polynomial programs. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):100:1– 100:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586052.

Guo:2020:PST

[GJJ^{+}20] Zheng Guo, Michael James, David Justo, Jiaxiao Zhou, Ziteng Wang, Ranjit Jhala, and Nadia Polikarpova. Program synthesis by type-guided abstraction refinement. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):12:1–12:28, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371080.

Goel:2021:PMB

[GJK^{+}21] Aviral Goel, Jan Jecmen, Sebastián Krynski, Olivier Flückiger, and Jan Vitek. Promises are made to be broken: migrating R to strict semantics. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):101:1–101:20, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485478.

Graf:2020:LYG

[GJS20] Sebastian Graf, Simon Peyton Jones, and Ryan G. Scott. Lower your guards: a compositional pattern-match coverage checker. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):107:1–107:30, August

2020. URL https://dl.acm.
org/doi/10.1145/3408989.

Grech:2018:MSG

[GKJ^{+}18] Neville Grech, Michael Kong, Anton Jurisevic, Lexi Brent, Bernhard Scholz, and Yannis Smaragdakis. MadMax: surviving out-of-gas conditions in Ethereum smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):116:1– 116:27, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276486.

Gomes:2017:VSE

[GKMB17] Victor B. F. Gomes, Martin Kleppmann, Dominic P. Mulligan, and Alastair R. Beresford. Verifying strong eventual consistency in distributed systems. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):109:1–109:??, October 2017. CODEN ???? ISSN 2475-1421.

Ghica:2022:HLE

[GLBP22] Dan Ghica, Sam Lindley, Marcos Maroñas Bravo, and Maciej Piróg. High-level effect handlers in C++. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):183:1–183:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563445.

Gomez-Londono:2020:DYS

[GLPS^{+}20] Alejandro Gómez-Londoño, Johannes Åman Pohjola, Hira Taqdees Syeda, Magnus O. Myreen, and Yong Kiam Tan. Do you have space for dessert? A verified space cost semantics for CakeML programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):204:1–204:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428272.

Gulwani:2020:SIT

[GLR^{+}20] Sumit Gulwani, Vu Le, Arjun Radhakrishna, Ivan Radicek, and Mohammad Raza. Structure interpretation of text formats. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):212:1– 212:29, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428280.

Griffin:2020:TTL

[GLSY20] Jeremiah Griffin, Mohsen Lesani, Narges Shadab, and Xizhe Yin. TLC: temporal logic of distributed components. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):123:1–123:30, August

2020. URL https://dl.acm.
org/doi/10.1145/3409005.

Grech:2022:EAD

[GLTS22] Neville Grech, Sifis Lagouvardos, Ilias Tsatiris, and Yannis Smaragdakis. Elipmoc: advanced decompilation of Ethereum smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):77:1– 77:27, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527321.

Germane:2021:NSL

[GM21] Kimball Germane and Jay Mc-Carthy. Newly-single and loving it: improving higher-order must-alias analysis with heap fragments. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 96:1–96:28, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473601.

Gutsfeld:2021:AFA

[GMOO21] Jens Oliver Gutsfeld, Markus Müller-Olm, and Christoph Ohrem. Automata and fixpoints for asynchronous hyperproperties. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 38:1–38:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434319.

Goncharov:2023:THO

[GMS^{+}23] Sergey Goncharov, Stefan Milius, Lutz Schröder, Stelios Tsampas, and Henning Urbat. Towards a higher-order mathematical operational semantics. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 22:1–22:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571215.

Goldstein:2022:PR

[GP22] Harrison Goldstein and Benjamin C. Pierce. Parsing randomness. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):128:1–128:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563291.

Ghilezan:2021:PSA

[GPP^{+}21] Silvia Ghilezan, Jovanka Pantović, Ivan Prokić, Alceste Scalas, and Nobuko Yoshida. Precise subtyping for asynchronous multiparty sessions. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):16:1–16:28, January

2021. URL https://dl.acm.
org/doi/10.1145/3434297.

Gupta:2020:CGC

[GRB20] Shubhani Gupta, Abhishek Rose, and Sorav Bansal. Counterexample-guided correlation algorithm for translation validation. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):221:1–221:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428289.

Gao:2021:PAO

[GRS^{+}21] Xiang Gao, Arjun Radhakrishna, Gustavo Soares, Ridwan Shariffdeen, Sumit Gulwani, and Abhik Roychoudhury. APIfix: outputoriented program synthesis for combating breaking changes in libraries. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):161:1–161:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485538.

Genc:2019:DAU

[GRXB19] Kaan Genç, Jake Roemer, Yufan Xu, and Michael D. Bond. Dependence-aware, unbounded sound predictive race detection. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):179:1–179:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360605.

Grech:2017:PTU

[GS17] Neville Grech and Yannis Smaragdakis. P/Taint: unified points-to and taint analysis. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):102:1– 102:??, October 2017. CODEN ???? ISSN 2475-1421.

Garg:2022:SCQ

[GS22] Pranav Garg and Srinivasan H. Sengamedu. Synthesizing code quality rules from examples. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):187:1– 187:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563350.

Gratzer:2019:IMD

[GSB19] Daniel Gratzer, Jonathan Sterling, and Lars Birkedal. Implementing a modal dependent type theory. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):107:1–107:29, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341711.

Gancher:2023:CCE

[GSF^{+}23] Joshua Gancher, Kristina Sojakova, Xiong Fan, Elaine Shi, and Greg Morrisett. A core calculus for equational proofs of cryptographic protocols. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 30:1–30:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571223.

Gaher:2022:SSL

[GSS^{+}22] Lennard Gäher, Michael Sammler, Simon Spies, Ralf Jung, Hoang-Hai Dang, Robbert Krebbers, Jeehoon Kang, and Derek Dreyer. Simuliris: a separation logic framework for verifying concurrent program optimizations. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 28:1–28:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498689.

Giarrusso:2020:SSS

[GST^{+}20] Paolo G. Giarrusso, Léo Stefanesco, Amin Timany, Lars Birkedal, and Robbert Krebbers. Scala step-by-step: soundness for DOT with step-indexed logical relations in Iris. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):114:1–114:29, August

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3408996.

Giannarakis:2021:PPP

[GSW21] Nick Giannarakis, Alexandra Silva, and David Walker. ProbNV: probabilistic verification of network control planes. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 90:1–90:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473595.

Georges:2022:TCE

[GTB22] Aïna Linn Georges, Alix Trieu, and Lars Birkedal. Le temps des cérises: efficient temporal stack safety on capability machines using directed capabilities. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):74:1–74:30, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3527318.

Grosser:2020:FLP

[GTF^{+}20] Tobias Grosser, Theodoros Theodoridis, Maximilian Falkenstein, Arjun Pitchanathan, Michael Kruse, Manuel Rigger, Zhendong Su, and Torsten Hoefler. Fast linear programming through transprecision computing on small and sparse data. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):195:1–195:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428263.

Gokhale:2021:AMS

[GTT21] Satyajit Gokhale, Alexi Turcotte, and Frank Tip. Automatic migration from synchronous to asynchronous JavaScript APIs. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):160:1– 160:27, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485537.

Gu:2023:OCS

[GTU23] Yu Gu, Takeshi Tsukada, and Hiroshi Unno. Optimal CHC solving via termination proofs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 21:1–21:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571214.

Goel:2019:DIU

[GV19] Aviral Goel and Jan Vitek. On the design, implementation, and use of laziness in R. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):153:1–153:27, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360579.

Gorjiara:2020:SSE

[GXD20] Hamed Gorjiara, Guoqing Harry Xu, and Brian Demsky. Satune: synthesizing efficient SAT encoders. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):146:1–146:32, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428214.

Gheri:2023:HMS

[GY23] Lorenzo Gheri and Nobuko Yoshida. Hybrid multiparty session types: Compositionality for protocol specification through endpoint projection. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):79:1– 79:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586031.

Geisler:2020:GTG

[GYK^{+}20] Dietrich Geisler, Irene Yoon, Aditi Kabra, Horace He, Yinnon Sanders, and Adrian Sampson. Geometry types for graphics programming. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):173:1–173:25, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428241.

Hamana:2017:HPY

[Ham17]  Makoto Hamana.  How to

prove your calculus is decidable: practical applications of second-order algebraic theories and computation. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):22:1–22:??, September 2017. CODEN ???? ISSN 2475-1421.

Hinrichsen:2020:AST

[HBK20] Jonas Kastberg Hinrichsen, Jesper Bengtson, and Robbert Krebbers. Actris: sessiontype based reasoning in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):6:1–6:30, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371074.

Henriksen:2019:DGS

[HBP19] Ian Henriksen, Gianfranco Bilardi, and Keshav Pingali. Derivative grammars: a symbolic approach to parsing with derivatives. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):127:1–127:28, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360553.

Hirsch:2021:GSP

[HC21] Andrew K. Hirsch and Ethan Cecchetti. Giving semantics to program-counter labels via secure effects. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 35:1–35:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434316.

Huang:2021:DIU

[HdSO21] Xuejing Huang and Bruno C. d. S. Oliveira. Distributing intersection and union types with splits and duality (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 89:1–89:24, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473594.

Huang:2019:MVA

[HFCG19] Mingzhang Huang, Hongfei Fu, Krishnendu Chatterjee, and Amir Kafshdar Goharshady. Modular verification for almostsure termination of probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):129:1–129:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360555.

Hoeflich:2022:HIK

[HFS22] Joshua Hoeflich, Robert Bruce Findler, and Manuel Serrano. Highly illogical, Kirk: spotting type mismatches in the large despite broken contracts, unsound types, and too many linters. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):142:1–142:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563305.

Hirsch:2022:PHO

[HG22] Andrew K. Hirsch and Deepak Garg. Pirouette: higherorder typed functional choreographies. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 23:1–23:27, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498684.

Haudebourg:2020:RLT

[HGJ20] Timothée Haudebourg, Thomas Genet, and Thomas Jensen. Regular language type inference with term rewriting. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):112:1–112:29, August

2020. URL https://dl.acm.
org/doi/10.1145/3408994.

Hackett:2018:PPO

[HH18] Jennifer Hackett and Graham Hutton. Parametric polymorphism and operational improvement. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):68:1–68:24, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236763.

Hackett:2019:CNC

[HH19] Jennifer Hackett and Graham Hutton. Call-by-need is clairvoyant call-by-value. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):114:1–114:23, July

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3341718.

He:2020:TAE

[HH20] Fei He and Jitao Han. Termination analysis for evolving programs: an incremental approach by reusing certified modules. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):199:1–199:27, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428267.

Haner:2020:ABO

[HHT20] Thomas Häner, Torsten Hoefler, and Matthias Troyer. Assertion-based optimization of quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):133:1–133:20, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428201.

Henry:2021:CSA

[HHY^{+}21] Rawn Henry, Olivia Hsu, Rohan Yadav, Stephen Chou, Kunle Olukotun, Saman Amarasinghe, and Fredrik Kjolstad. Compilation of sparse array programming models. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):128:1–128:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485505.

Hickey:2020:HC

[Hic20]  Rich Hickey.  A history

of Clojure. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):71:1–71:46, June 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3386321.

Hirschowitz:2019:FMS

[Hir19]  Tom Hirschowitz.  Familial

monads and structural operational semantics. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 21:1–21:28, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290334.

Holk:2018:SCC

[HJL^{+}18] Lukáš Holk, Petr Janku, Anthony W. Lin, Philipp Rümmer, and Tomáš Vojnar. String constraints with concatenation and transducers solved efficiently. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 4:1–4:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Hui:2020:AS

[HK20] Roger K. W. Hui and Morten J.

Kromberg. APL since 1978.

Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):69:1–69:108, June

2020. URL https://dl.acm.
org/doi/abs/10.1145/3386319.

Heunen:2022:QIE

[HK22] Chris Heunen and Robin Kaarsgaard. Quantum information effects. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 2:1–2:27, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498663.

Hark:2020:ALH

[HKGK20] Marcel Hark, Benjamin Lucien Kaminski, Jürgen Giesl, and Joost-Pieter Katoen. Aiming low is harder: induction for lower bounds in probabilistic program verification. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 37:1–37:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371105.

Honore:2021:MAA

[HKSS21] Wolf Honoré, Jieung Kim, Ji-Yong Shin, and Zhong Shao. Much ADO about failures: a fault-aware model for compositional verification of strongly consistent distributed systems. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):97:1– 97:31, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485474.

Handa:2021:OAD

[HKVR21] Shivam Handa, Konstantinos Kallas, Nikos Vasilakis, and Martin C. Rinard. An orderaware dataflow model for parallel Unix pipelines. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):65:1–65:28, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3473570.

Hu:2020:UDF

[HL20] Jason Z. S. Hu and Ondrej Lhoták. Undecidability of d: and its decidable fragments. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):9:1–9:30, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371077.

Hagedorn:2020:AHP

[HLK^{+}20] Bastian Hagedorn, Johannes Lenfers, Thomas Kœhler, Xueying Qin, Sergei Gorlatch, and Michel Steuwer. Achieving high-performance the functional way: a functional pearl on expressing high-performance optimizations as rewrite strategies. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 92:1–92:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408974.

Hillerstrom:2020:EEA

[HLL20] Daniel Hillerström, Sam Lindley, and John Longley. Effects for efficiency: asymptotic speedup with first-class control. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):100:1– 100:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408982.

Houshmand:2021:PGD

[HLV21] Farzin Houshmand, Mohsen Lesani, and Keval Vora. Grafs: declarative graph analytics. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 83:1–83:32, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473588.

Haas:2022:CCT

[HMdL22] Thomas Haas, Roland Meyer, and Hernán Ponce de León. CAAT: consistency as a theory. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):129:1– 129:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563292.

He:2022:PIR

[HMP^{+}22] Wenlei He, Julián Mestre, Sergey Pupyrev, Lei Wang, and Hongtao Yu. Profile inference revisited. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):52:1–52:24, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498714.

Hameer:2019:TAF

[HP19] Aliya Hameer and Brigitte Pientka. Teaching the art of functional programming using automated grading (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):115:1–115:15, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341719.

Ho:2022:ARV

[HP22] Son Ho and Jonathan Protzenko.

Aeneas: Rust verification by functional translation. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):116:1–116:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547647.

Hainry:2023:GNP

[HP23] Emmanuel Hainry and Romain Péchoux. A general noninterference policy for polynomial time. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 28:1–28:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571221.

Horn:2018:IRL

[HPC18] Rudi Horn, Roly Perera, and James Cheney. Incremental relational lenses. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):74:1–74:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236769.

Herklotz:2021:FVH

[HPRW21] Yann Herklotz, James D. Pollard, Nadesh Ramanathan, and John Wickerson. Formal verification of high-level synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):117:1– 117:30, October 2021. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3485494.

Hietala:2021:VOQ

[HRH^{+}21] Kesha Hietala, Robert Rand, Shih-Han Hung, Xiaodi Wu, and Michael Hicks. A verified optimizer for quantum circuits. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 37:1–37:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434318.

Hunt:2023:RSS

[HSS23] Sebastian Hunt, David Sands, and Sandro Stucki. Reconciling Shannon and Scott with a lattice of computable information. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 68:1–68:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571740.

Hirsch:2018:SLS

[HT18] Andrew K. Hirsch and Ross Tate. Strict and lazy semantics for effects: layering monads and comonads. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):88:1–88:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236783.

Hoang:2022:RTH

[HTLS22] Tram Hoang, Anton Trunov, Leonidas Lampropoulos, and Ilya Sergey. Random testing of a higher-order blockchain language (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):122:1–122:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547653.

Handley:2020:LYA

[HVH20] Martin A. T. Handley, Niki Vazou, and Graham Hutton. Liquidate your assets: reasoning about resource usage in liquid Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 24:1–24:27, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371092.

Hamza:2019:SFF

[HVK19] Jad Hamza, Nicolas Voirol, and Viktor Kuncak. System FR: formalized foundations for the Stainless verifier. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):166:1–166:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360592.

Holtzen:2020:SEI

[HVM20] Steven Holtzen, Guy Van den Broeck, and Todd Millstein. Scaling exact inference for discrete probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):140:1–140:31, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428208.

He:2021:TSE

[HWC^{+}21] Paul He, Eddy Westbrook, Brent Carmer, Chris Phifer, Valentin Robert, Karl Smeltzer, Andrei Ştefănescu, Aaron Tomb, Adam Wick, Matthew Yacavone, and Steve Zdancewic. A type system for extracting functional specifications from memory-safe imperative programs. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):135:1–135:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485512.

Iorga:2021:SSM

[IDSW21] Dan Iorga, Alastair F. Donaldson, Tyler Sorensen, and John Wickerson. The semantics of shared memory in Intel CPU/FPGA systems. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):120:1–120:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485497.

Ikebuchi:2022:CDS

[IEC22] Mirai Ikebuchi, Andres Erbsen, and Adam Chlipala. Certifying derivation of state machines from coroutines. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):24:1–24:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498685.

Ingalls:2020:ESS

[Ing20] Daniel Ingalls. The evolution of Smalltalk: from Smalltalk-72 through Squeak. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):85:1–85:101, June

2020. URL https://dl.acm.
org/doi/abs/10.1145/3386335.

Ishimwe:2021:PDA

[INN21] Didier Ishimwe, KimHao Nguyen, and ThanhVu Nguyen. Dynaplex: analyzing program complexity using dynamically inferred recurrence relations. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):138:1–138:23, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485515.

Inala:2018:WIW

[IS18] Jeevana Priya Inala and Rishabh Singh. WebRelate: integrating web data with spreadsheets using examples. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 2:1–2:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Igarashi:2017:PGT

[ISI17] Yuu Igarashi, Taro Sekiyama, and Atsushi Igarashi. On polymorphic gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP):40:1–40:??, September 2017. CODEN ???? ISSN 2475-1421.

Igarashi:2017:GST

[ITVW17] Atsushi Igarashi, Peter Thiemann, Vasco T. Vasconcelos, and Philip Wadler. Gradual session types. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):38:1–38:??, September 2017. CODEN ???? ISSN 2475-1421.

Jaber:2020:SAC

[Jab20] Guilhem Jaber. SyTeCi: automating contextual equivalence for higher-order programs with references. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 59:1–59:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371127.

Jacobs:2021:PPP

[Jac21] Jules Jacobs. Paradoxes of probabilistic programming: and how to condition on events of measure zero with infinitesimal probabilities. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):58:1–58:26, January

2021. URL https://dl.acm.
org/doi/10.1145/3434339.

Jacobs:2023:HOL

[JB23] Jules Jacobs and Stephanie Balzer. Higher-order leak and deadlock free locks. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):36:1–36:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571229.

Jacobs:2022:CGM

[JBK22a] Jules Jacobs, Stephanie Balzer, and Robbert Krebbers. Connectivity graphs: a method for proving deadlock freedom based on separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):1:1–1:33, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498662.

Jacobs:2022:MGF

[JBK22b] Jules Jacobs, Stephanie Balzer, and Robbert Krebbers. Multiparty GV: functional multiparty session types with certified deadlock freedom. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):107:1–107:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3547638.

Jung:2020:SBA

[JDKD20] Ralf Jung, Hoang-Hai Dang, Jeehoon Kang, and Derek Dreyer. Stacked borrows: an aliasing model for Rust. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):41:1–41:32, January

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3371109.

Jacobs:2022:PSM

[JDT22] Koen Jacobs, Dominique Devriese, and Amin Timany. Purity of an ST monad: full abstraction by semantically typed back-translation. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):82:1–82:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3527326.

Jang:2022:MMU

[JGMP22] Junyoung Jang, Samuel Gélineau, Stefan Monnier, and Brigitte Pientka. Moebius: metaprogramming using contextual types: the stage where system f can pattern match on itself. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 39:1–39:27, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498700.

James:2020:DFS

[JGW^{+}20] Michael B. James, Zheng Guo, Ziteng Wang, Shivani Doshi, Hila Peleg, Ranjit Jhala, and Nadia Polikarpova. Digging for fold: synthesis-aided API discovery for Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):205:1–205:27, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428273.

Jeong:2017:DDC

[JJCO17] Sehun Jeong, Minseok Jeon, Sungdeok Cha, and Hakjoo Oh. Data-driven context-sensitivity for points-to analysis. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):100:1–100:??, October 2017. CODEN ???? ISSN 2475-1421.

Jung:2018:RSF

[JJKD18] Ralf Jung, Jacques-Henri Jourdan, Robbert Krebbers, and Derek Dreyer. RustBelt: securing the foundations of the Rust programming language. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 66:1–66:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Jeon:2018:PSP

[JJO18] Minseok Jeon, Sehun Jeong, and Hakjoo Oh. Precise and scalable points-to analysis via data-driven context tunneling. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):140:1–140:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276510.

Jagadeesan:2020:PPS

[JJR20] Radha Jagadeesan, Alan Jeffrey, and James Riely. Pomsets with preconditions: a simple model of relaxed memory. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):194:1– 194:30, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428262.

Jochems:2023:HOM

[JJR23] Jerome Jochems, Eddie Jones, and Steven Ramsay. Higherorder MSL Horn constraints. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):69:1–69:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571262.

Jia:2022:SVQ

[JKL^{+}22] Xiaodong Jia, Andre Kornell, Bert Lindenhovius, Michael Mislove, and Vladimir Zamdzhiev. Semantics for variational quantum programming. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 26:1–26:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498687.

Jia:2021:DBP

[JKT21] Xiaodong Jia, Ashish Kumar, and Gang Tan. A derivative-based parser generator for visibly pushdown grammars. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):151:1– 151:24, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485528.

Ji:2023:IOG

[JKXH23] Ruyi Ji, Chaozhe Kong, Yingfei Xiong, and Zhenjiang Hu. Improving oracle-guided inductive synthesis by efficient question selection. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):103:1–103:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586055.

Jeon:2020:LGB

[JLO20] Minseok Jeon, Myungho Lee, and Hakjoo Oh. Learning graph-based heuristics for pointer analysis without handcrafting applicationspecific features. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):179:1–179:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428247.

Jung:2020:FOP

[JLP^{+}20] Ralf Jung, Rodolphe Lepigre, Gaurav Parthasarathy, Marianna Rapoport, Amin Timany, Derek Dreyer, and Bart Jacobs. The future is ours: prophecy variables in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 45:1–45:32, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371113.

Jones:2020:PTC

[JME20] Mark P. Jones, J. Garrett Morris, and Richard A. Eisenberg. Partial type constructors: or, making ad hoc datatypes less ad hoc. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 40:1–40:28, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371108.

Jeon:2022:RCC

[JO22] Minseok Jeon and Hakjoo Oh. Return of CFA: call-site sensitivity can be superior to object sensitivity even for objectoriented programs. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):58:1–58:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498720.

Jangda:2019:FFS

[JPBG19] Abhinav Jangda, Donald Pinckney, Yuriy Brun, and Arjun Guha. Formal foundations of serverless computing. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):149:1–149:26, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360575.

Jin:2022:NAS

[JPR22] Charles Jin, Phitchaya Mangpo Phothilimthana, and Sudip Roy. Neural architecture search using property guided synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):166:1– 166:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563329.

Jones:2021:IDR

[JR21] Eddie Jones and Steven Ramsay. Intensional datatype refinement: with application to scalable verification of patternmatch safety. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):

55:1–55:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434336.

Jeffrey:2022:LSC

[JRB^{+}22] Alan Jeffrey, James Riely, Mark Batty, Simon Cooksey, Ilya Kaysin, and Anton Podkopaev. The leaky semicolon: compositional semantic dependencies for relaxed-memory concurrency. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 54:1–54:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498716.

Ji:2020:GDP

[JSXH20] Ruyi Ji, Yican Sun, Yingfei Xiong, and Zhenjiang Hu. Guiding dynamic programming via structural probability for accelerating programming by example. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):224:1–224:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428292.

Jacobs:2021:FAS

[JTD21] Koen Jacobs, Amin Timany, and Dominique Devriese. Fully abstract from static to gradual. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):7:1–7:30, January 2021. URL https:

//dl.acm.org/doi/10.1145/
3434288.

Jacobs:2023:FCB

[JW23] Jules Jacobs and Thorsten Wißmann. Fast coalgebraic bisimilarity minimization. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 52:1–52:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571245.

Jaber:2021:PQM

[JWJ^{+}21] Nouraldin Jaber, Christopher Wagner, Swen Jacobs, Milind Kulkarni, and Roopsha Samanta. QuickSilver: modeling and parameterized verification for distributed agreementbased systems. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):157:1–157:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485534.

Ji:2021:GST

[JXXH21] Ruyi Ji, Jingtao Xia, Yingfei Xiong, and Zhenjiang Hu. Generalizable synthesis through unification. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA): 167:1–167:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485544.

Kavvos:2019:MCI

[Kav19]  G. A. Kavvos.  Modali-

ties, cohesion, and information flow. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 20:1–20:29, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290333.

Klimis:2023:TBC

[KCB^{+}23] Vasileios Klimis, Jack Clark, Alan Baker, David Neto, John Wickerson, and Alastair F. Donaldson. Taking back control in an intermediate representation for GPU computing. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 60:1–60:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571253.

Kincaid:2018:NLR

[KCBR18] Zachary Kincaid, John Cyphert, Jason Breck, and Thomas Reps. Non-linear reasoning for invariant synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 54:1–54:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Ko:2022:DGP

[KCL22] Hsiang-Shang Ko, Liang-Ting Chen, and Tzu-Chi Lin. Datatype-generic programming meets elaborator reflection. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):98:1–98:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547629.

Koutris:2023:FGC

[KD23] Paraschos Koutris and Shaleen Deep. The fine-grained complexity of CFL reachability. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):59:1– 59:??, January 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571252.

Kim:2023:UL

[KDR23] Jinwoo Kim, Loris DAntoni, and Thomas Reps. Unrealizability logic. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 23:1–23:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571216.

Keidel:2019:SRC

[KE19] Sven Keidel and Sebastian Erdweg. Sound and reusable components for abstract interpretation. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):176:1–176:28, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360602.

Kesner:2022:FGC

[Kes22] Delia Kesner. A fine-grained computational interpretation of Girard’s intuitionistic proofnets. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):8:1–8:28, January 2022. CODEN ???? ISSN 2475-1421 (electronic).

URL https://dl.acm.org/
doi/10.1145/3498669.

Kaki:2018:SRT

[KESJ18] Gowtham Kaki, Kapil Earanky, KC Sivaramakrishnan, and Suresh Jagannathan. Safe replication through bounded concurrency verification. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):164:1–164:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276534.

Kohler:2020:RSC

[KEW^{+}20] Mirko Köhler, Nafise Eskandani, Pascal Weisenburger, Alessandro Margara, and Guido Salvaneschi. Rethinking safe consistency in distributed objectoriented programming. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):188:1–188:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428256.

Kiss:2019:HOT

[KFEJ19] Csongor Kiss, Tony Field, Susan Eisenbach, and Simon Peyton Jones. Higherorder type-level programming in Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):102:1–102:26, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341706.

Kazerounian:2021:PSS

[KFM21] Milod Kazerounian, Jeffrey S.

Foster, and Bonan Min. Sim-Typer: sound type inference for Ruby using type equality prediction. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):106:1–106:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485483.

Koppel:2022:SEP

[KGdV^{+}22] James Koppel, Zheng Guo, Edsko de Vries, Armando Solar-Lezama, and Nadia Polikarpova. Searching entangled program spaces. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 91:1–91:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547622.

Ko:2018:ABB

[KH18] Hsiang-Shang Ko and Zhenjiang Hu. An axiomatic basis for bidirectional programming. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 41:1–41:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Kahn:2021:AAR

[KH21] David M. Kahn and Jan Hoffmann. Automatic amortized resource analysis with the quantum physicist’s method. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):76:1–76:29, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3473581.

Kim:2021:SGS

[KHDR21] Jinwoo Kim, Qinheping Hu, Loris D’Antoni, and Thomas Reps. Semantics-guided synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 30:1–30:32, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434311.

Keuchel:2022:VSE

[KHLD22] Steven Keuchel, Sander Huyghebaert, Georgy Lukyanov, and Dominique Devriese. Verified symbolic execution with Kripke specification monads (and no meta-programming). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):97:1–97:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547628.

King:2020:HGP

[Kin20]  Paul King.  A history of

the Groovy programming language. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):76:1– 76:53, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/
3386326.

Krebbers:2018:MGE

[KJJ^{+}18] Robbert Krebbers, Jacques-Henri Jourdan, Ralf Jung, Joseph Tassarotti, Jan-Oliver Kaiser, Amin Timany, Arthur Charguéraud, and Derek Dreyer. MoSeL: a general, extensible modal framework for interactive proofs in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):77:1–77:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236772.

Krawiec:2022:PCA

[KJK^{+}22] Faustyna Krawiec, Simon Peyton Jones, Neel Krishnaswami, Tom Ellis, Richard A. Eisenberg, and Andrew Fitzgibbon. Provably correct, asymptotically efficient, higher-order reverse-mode automatic differentiation. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 48:1–48:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498710.

Kaposi:2019:CQI

[KKA19] Ambrus Kaposi, András Kovács, and Thorsten Altenkirch. Constructing quotient inductiveinductive types. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):2:1–2:24, January

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3290315.

Kjolstad:2017:TAC

[KKC^{+}17] Fredrik Kjolstad, Shoaib Kamil, Stephen Chou, David Lugato, and Saman Amarasinghe. The tensor algebra compiler. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):77:1– 77:??, October 2017. CODEN ???? ISSN 2475-1421.

Kaminski:2017:RAC

[KKCV17] Ted Kaminski, Lucas Kramer, Travis Carlson, and Eric Van Wyk. Reliable and automatic composition of language extensions to C: the ableC extensible language framework. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):98:1–98:??, October 2017. CODEN ???? ISSN 2475-1421.

Karachalias:2021:ECA

[KKPS21] Georgios Karachalias, Filip Koprivec, Matija Pretnar, and Tom Schrijvers. Efficient compilation of algebraic effect handlers. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):102:1–102:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485479.

Kokologiannakis:2021:PPS

[KKRV21] Michalis Kokologiannakis, Ilya Kaysin, Azalea Raad, and Viktor Vafeiadis. PerSeVerE: persistency semantics for verification under ext4. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):43:1–43:29, January

2021. URL https://dl.acm.
org/doi/10.1145/3434324.

Koppel:2022:ADC

[KKSL22] James Koppel, Jackson Kearl, and Armando Solar-Lezama. Automatically deriving controlflow graph generators from operational semantics. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):117:1–117:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547648.

Konnov:2019:TMC

[KKT19] Igor Konnov, Jure Kukovec, and Thanh-Hai Tran. TLA+ model checking made symbolic. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):123:1– 123:30, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360549.

Kincaid:2023:WLM

[KKZ23] Zachary Kincaid, Nicolas Koh, and Shaowei Zhu. When less is more: Consequence-finding in a weak theory of arithmetic. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 44:1–44:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571237.

Khyzha:2021:TXT

[KL21] Artem Khyzha and Ori Lahav. Taming x86-TSO persistency. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 47:1–47:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434328.

Kabir:2020:DCO

[KLL20] Ifaz Kabir, Yufeng Li, and Ondrej Lhoták. ιDOT: a DOT calculus with object initialization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):208:1–208:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428276.

Krogmeier:2022:SAU

[KLMM22] Paul Krogmeier, Zhengyao Lin, Adithya Murali, and P. Madhusudan. Synthesizing axiomatizations using logic learning. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):185:1–185:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563348.

Karimov:2022:WDA

[KLO^{+}22] Toghrul Karimov, Engel Lefaucheux, Joël Ouaknine, David Purser, Anton Varonka, Markus A.

Whiteland, and James Worrell. What’s decidable about linear loops? Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 65:1–65:25, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498727.

Kokologiannakis:2018:ESM

[KLSV18] Michalis Kokologiannakis, Ori Lahav, Konstantinos Sagonas, and Viktor Vafeiadis. Effective stateless model checking for C/C++ concurrency. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):17:1–17:??, January 2018. CODEN ???? ISSN 2475-1421.

Kokologiannakis:2023:KAW

[KLV23] Michalis Kokologiannakis, Ori Lahav, and Viktor Vafeiadis. Kater: Automating weak memory model metatheory and consistency checking. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):19:1–19:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571212.

Kristensen:2017:TTS

[KM17] Erik Krogh Kristensen and Anders Møller. Type test scripts for TypeScript testing. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):90:1–90:??, October 2017. CODEN ???? ISSN 2475-1421.

Kong:2020:SQL

[KM20] Lingkun Kong and Konstantinos Mamouras. StreamQL: a query language for processing streaming time series. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):183:1–183:32, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428251.

Krogmeier:2022:LFF

[KM22] Paul Krogmeier and P. Madhusudan. Learning formulas in finite variable logics. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):10:1–10:28, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498671.

Krogmeier:2023:LDL

[KM23] Paul Krogmeier and P. Madhusudan. Languages with decidable learning: a metatheorem. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1): 80:1–80:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586032.

Kalita:2022:SAT

[KMD^{+}22] Pankaj Kumar Kalita, Sujit Kumar Muduli, Loris D’Antoni, Thomas Reps, and Subhajit Roy. Synthesizing abstract transformers. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):171:1–171:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563334.

Kokologiannakis:2022:TSO

[KMGV22] Michalis Kokologiannakis, Iason Marmanis, Vladimir Gladstein, and Viktor Vafeiadis. Truly stateless, optimal dynamic partial order reduction. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 49:1–49:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498711.

Kavvos:2020:REF

[KMLD20] G. A. Kavvos, Edward Morehouse, Daniel R. Licata, and Norman Danner. Recurrence extraction for functional programs through call-by-push-value. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 15:1–15:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371083.

Kerinec:2023:WPR

[KMO23] Axel Kerinec, Giulio Manzonetto, and Federico Olimpieri. Why are proofs relevant in proof-relevant models? Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):8:1–8:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3571201.

Kokke:2019:BLT

[KMP19] Wen Kokke, Fabrizio Montesi, and Marco Peressotti. Better late than never: a fullyabstract semantics for classical processes. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 24:1–24:29, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290337.

Krikava:2019:SIE

[KMV19] Filip Krikava, Heather Miller, and Jan Vitek. Scala implicits are everywhere: a large-scale study of the use of Scala implicits in the wild. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):163:1–163:28, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360589.

Kolosick:2022:ITN

[KNJ^{+}22] Matthew Kolosick, Shravan Narayan, Evan Johnson, Conrad Watt, Michael LeMay, Deepak Garg, Ranjit Jhala, and Deian Stefan. Isolation without taxation: near-zero-cost transitions for WebAssembly and SFI. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 27:1–27:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498688.

Kaki:2018:ATC

[KNNJ18] Gowtham Kaki, Kartik Nagar, Mahsa Najafzadeh, and Suresh Jagannathan. Alone together: compositional reasoning and inference for weak isolation. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 27:1–27:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Khairunnesa:2017:EIB

[KNNR17] Samantha Syeda Khairunnesa, Hoan Anh Nguyen, Tien N. Nguyen, and Hridesh Rajan. Exploiting implicit beliefs to resolve sparse usage problem in usage-based specification mining. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):83:1– 83:??, October 2017. CODEN ???? ISSN 2475-1421.

Kallas:2020:DDO

[KNSA20] Konstantinos Kallas, Filip Niksic, Caleb Stanford, and Rajeev Alur. DiffStream: differential output testing for stream processing programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):153:1–153:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428221.

Kang:2023:MCB

[KO23] Chan Gu Kang and Hakjoo Oh. Modular componentbased quantum circuit synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):87:1– 87:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586039.

Kodosky:2020:L

[Kod20] Jeffrey Kodosky. LabVIEW. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):78:1–78:54, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386328.

Kovacs:2020:EFC

[Kov20] András Kovács. Elaboration with first-class implicit function types. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):101:1– 101:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408983.

Kovacs:2022:SCT

[Kov22] András Kovács. Staged compilation with two-level type theory. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 110:1–110:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547641.

Kalhauge:2018:SDP

[KP18a] Christian Gram Kalhauge and Jens Palsberg. Sound deadlock prediction. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):146:1–146:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276516.

Kuncar:2018:SCD

[KP18b] Ondřej Kunčar and Andrei Popescu. Safety and conservativity of definitions in HOL and Isabelle/HOL. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 24:1–24:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Kjelstrom:2022:DCI

[KP22] Adam Husted Kjelstrøm and Andreas Pavlogiannis. The decidability and complexity of interleaved bidirected Dyck reachability. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 12:1–12:26, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498673.

Keidel:2018:CSP

[KPE18] Sven Keidel, Casper Bach Poulsen, and Sebastian Erdweg. Compositional soundness proofs of abstract interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 72:1–72:26, July 2018. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3236767.

Kolesar:2022:CEN

[KPH22] John C. Kolesar, Ruzica Piskac, and William T. Hallahan. Checking equivalence in a nonstrict language. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):177:1–177:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563340.

Kuperberg:2021:CPS

[KPP21] Denis Kuperberg, Laureline Pinault, and Damien Pous. Cyclic proofs, system t, and the power of contraction. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):1:1–1:28, January

                                        2021. URL https://dl.acm.
                                        org/doi/10.1145/3434282.

Kaki:2019:MRD

[KPSJ19] Gowtham Kaki, Swarn Priya, KC Sivaramakrishnan, and Suresh Jagannathan. Mergeable replicated data types. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):154:1–154:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360580.

Koppel:2018:OTM

[KPSL18] James Koppel, Varot Premtoon, and Armando Solar-Lezama. One tool, many languages: language-parametric transformation with incremental parametric syntax. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):122:1–122:28, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276492.

Kiss:2018:GDG

[KPW18] Csongor Kiss, Matthew Pickering, and Nicolas Wu. Generic deriving of generic traversals. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 85:1–85:30, July 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3236780.

Krauter:2021:PST

[KRB^{+}21] Nicolas Krauter, Patrick Raaf, Peter Braam, Reza Salkhordeh, Sebastian Erdweg, and André Brinkmann. Persistent software transactional memory in Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 63:1–63:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473568.

Kokologiannakis:2019:ELH

[KRV19] Michalis Kokologiannakis, Azalea Raad, and Viktor Vafeiadis. Effective lock handling in stateless model checking. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):173:1–173:26, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360599.

Koparkar:2021:ETT

[KRV^{+}21] Chaitanya Koparkar, Mike Rainey, Michael Vollmer, Milind Kulkarni, and Ryan R. Newton. Efficient tree-traversals: reconciling parallelism and dense data representations. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):91:1–91:29, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3473596.

K:2022:SCH

[KSG22] Hari Govind V. K., Sharon Shoham, and Arie Gurfinkel. Solving constrained Horn clauses modulo algebraic data types and recursive functions. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):60:1–60:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498722.

Kirisame:2022:OHL

[KSP22] Marisa Kirisame, Pranav Shenoy, and Pavel Panchekha. Optimal heap limits for reducing browser memory use. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):160:1–160:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563323.

Kumar:2020:SRI

[KSS20] Aayan Kumar, Vivek Seshadri, and Rahul Sharma. Shiftry: RNN inference in 2KB of RAM. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):182:1–182:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428250.

Koppel:2018:CFR

[KSSL18] James Koppel, Gabriel Scherer, and Armando Solar-Lezama. Capturing the future by replaying the past (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 76:1–76:29, July 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3236771.

Kirkham:2020:FEM

[KSTM20] Jake Kirkham, Tyler Sorensen, Esin Tureci, and Margaret Martonosi. Foundations of empirical memory consistency testing. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):226:1– 226:29, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428294.

Krishna:2018:GFC

[KSW18] Siddharth Krishna, Dennis Shasha, and Thomas Wies. Go with the flow: compositional abstractions for concurrent data structures. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 37:1–37:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Kobayashi:2023:HZV

[KTST23] Naoki Kobayashi, Kento Tanahashi, Ryosuke Sato, and Takeshi Tsukada. HFL(Z) validity checking for automated program verification. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):6:1–6:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571199.

Kim:2020:DPF

[KVT20] Sung Kook Kim, Arnaud J. Venet, and Aditya V. Thakur. Deterministic parallel fixpoint computation. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 14:1–14:33, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371082.

Kidney:2021:AWS

[KW21] Donnacha Oisín Kidney and Nicolas Wu. Algebras for weighted search. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 72:1–72:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473577.

Knoth:2020:LRT

[KWR^{+}20] Tristan Knoth, Di Wang, Adam Reynolds, Jan Hoffmann, and Nadia Polikarpova. Liquid resource types. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 106:1–106:29, August 2020.

URL https://dl.acm.org/
doi/10.1145/3408988.

Kammar:2022:FAM

[KyKS22] Ohad Kammar, Shin ya Katsumata, and Philip Saville. Fully abstract models for effectful λ-calculi via category-theoretic logical relations. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 44:1–44:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498705.

Kallas:2023:EMA

[KZA^{+}23] Konstantinos Kallas, Haoran Zhang, Rajeev Alur, Sebastian Angel, and Vincent Liu. Executing microservice applications on serverless, correctly. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):13:1–13:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571206.

Kaiser:2018:MTT

[KZK^{+}18] Jan-Oliver Kaiser, Beta Ziliani, Robbert Krebbers, Yann Régis-Gianas, and Derek Dreyer. Mtac2: typed tactics for backward reasoning in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):78:1–78:31, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236773.

Li:2021:DEP

[LA21] John M. Li and Andrew W. Appel. Deriving efficient program transformations from rewrite rules. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 74:1–74:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473579.

Le:2020:DDT

[LAF^{+}20] Ton Chanh Le, Timos Antonopoulos, Parisa Fathololumi, Eric Koskinen, and ThanhVu Nguyen. DynamiTe: dynamic termination and non-termination proofs. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):189:1– 189:30, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428257.

Lim:2021:AGC

[LAGN21] Jay P. Lim, Mridul Aanjaneya, John Gustafson, and Santosh Nagarakatte. An approach to generate correctly rounded math libraries for new floating point variants. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):29:1–29:30, January

2021. URL https://dl.acm.
org/doi/10.1145/3434310.

Lu:2023:GSC

[LB23] Sirui Lu and Rastislav Bodík.

Grisette: Symbolic compilation as a functional programming library. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 16:1–16:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571209.

Liu:2022:VTP

[LBCRK22] Amanda Liu, Gilbert Louis Bernstein, Adam Chlipala, and Jonathan Ragan-Kelley. Verified tensor-program optimization via high-level scheduling rewrites. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 55:1–55:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498717.

Leissa:2018:APE

[LBH^{+}18] Roland Leißa, Klaas Boesche, Sebastian Hack, Arsène Pérard-Gayot, Richard Membarth, Philipp Slusallek, André Müller, and Bertil Schmidt. AnyDSL: a partial evaluation framework for programming highperformance libraries. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):119:1–119:30, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276489.

Livinskii:2020:RTC

[LBR20] Vsevolod Livinskii, Dmitry Babokin, and John Regehr. Random testing for C and C++ compilers with YARP-Gen. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):196:1– 196:25, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428264.

Lubin:2021:HST

[LC21] Justin Lubin and Sarah E. Chasins. How staticallytyped functional programmers write code. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):155:1–155:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485532.

Lee:2023:ISS

[LC23] Woosuk Lee and Hangyeol Cho. Inductive synthesis of structurally recursive functional programs from non-recursive expressions. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 70:1–70:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571263.

Lubin:2020:PSL

[LCOC20] Justin Lubin, Nick Collins, Cyrus Omar, and Ravi Chugh. Program sketching with live bidirectional evaluation. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):109:1–109:29, August

2020. URL https://dl.acm.
org/doi/10.1145/3408991.

Lin:2023:GPC

[LCT^{+}23] Zhengyao Lin, Xiaohong Chen, Minh-Thai Trinh, John Wang, and Grigore Roşu. Generating proof certificates for a language-agnostic deductive program verifier. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):77:1–77:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586029.

Lew:2020:TTD

[LCTS^{+}20] Alexander K. Lew, Marco F.

Cusumano-Towner, Benjamin Sherman, Michael Carbin, and Vikash K. Mansinghka. Trace types and denotational semantics for sound programmable inference in probabilistic languages. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 19:1–19:32, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371087.

Lee:2021:CTP

[Lee21]  Woosuk Lee.  Combining

the top-down propagation and bottom-up enumeration for inductive program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 54:1–54:28, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434335.

Lemerre:2023:STA

[Lem23] Matthieu Lemerre. SSA translation is an abstract interpretation. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 65:1–65:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571258.

Liang:2018:PCO

[LF18] Hongjin Liang and Xinyu Feng. Progress of concurrent objects with partial methods. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):20:1–20:??, January 2018. CODEN ???? ISSN 2475-1421.

Liu:2022:SLI

[LFY^{+}22] Hongming Liu, Hongfei Fu, Zhiyong Yu, Jiaxin Song, and Guoqiang Li. Scalable linear invariant generation with Farkas’ lemma. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):132:1–132:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563295.

Lazarek:2021:HEB

[LGFD21] Lukas Lazarek, Ben Greenman, Matthias Felleisen, and Christos Dimoulas. How to evaluate blame for gradual types. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 68:1–68:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473573.

Lagouvardos:2020:PSM

[LGTS20] Sifis Lagouvardos, Neville Grech, Ilias Tsatiris, and Yannis Smaragdakis. Precise static modeling of Ethereum “memory”. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):190:1– 190:26, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428258.

Liu:2022:SFI

[LH22] Bozhen Liu and Jeff Huang.

SHARP: fast incremental contextsensitive pointer analysis for Java. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):88:1– 88:28, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527332.

Lattuada:2023:VVR

[LHC^{+}23] Andrea Lattuada, Travis Hance, Chanhee Cho, Matthias Brun, Isitha Subasinghe, Yi Zhou, Jon Howell, Bryan Parno, and Chris Hawblitzel. Verus: Verifying Rust programs using linear ghost types. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):85:1–85:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3586037.

Lee:2018:RHL

[LHJ^{+}18] Juneyoung Lee, Chung-Kil Hur, Ralf Jung, Zhengyang Liu, John Regehr, and Nuno P. Lopes. Reconciling highlevel optimizations and lowlevel code in LLVM. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):125:1–125:28, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276495.

Lampropoulos:2019:CGP

[LHP19] Leonidas Lampropoulos, Michael Hicks, and Benjamin C. Pierce. Coverage guided, property based testing. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):181:1–181:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360607.

Lew:2023:ASA

[LHSM23] Alexander K. Lew, Mathieu Huot, Sam Staton, and Vikash K. Mansinghka. ADEV: Sound automatic differentiation of expected values of probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 5:1–5:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571198.

Licker:2020:DFO

[LJ20] Nandor Licker and Timothy M.

Jones. Duplo: a framework for OCaml post-link optimisation. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 98:1–98:29, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408980.

LeGlaunec:2023:REM

[LKM23] Alexis Le Glaunec, Lingkun Kong, and Konstantinos Mamouras. Regular expression matching using bit vector automata. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1): 92:1–92:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3586044.

Lazarek:2020:DBS

[LKS^{+}20] Lukas Lazarek, Alexis King, Samanvitha Sundar, Robert Bruce Findler, and Christos Dimoulas. Does blame shifting work? Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):

65:1–65:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371133.

Lorenzen:2022:RCF

[LL22] Anton Lorenzen and Daan Leijen. Reference counting with frame limited reuse. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):103:1–103:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3547634.

Leijen:2023:TRM

[LL23] Daan Leijen and Anton Lorenzen. Tail recursion modulo context: an equational approach. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 40:1–40:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571233.

Liu:2020:TES

[LLB^{+}20] Fengyun Liu, Ondrej Lhoták, Aggelos Biboudis, Paolo G. Giarrusso, and Martin Odersky. A type-and-effect system for object initialization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):175:1–175:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428243.

Le:2022:QIS

[LLSS22] Xuan-Bach Le, Shang-Wei Lin, Jun Sun, and David Sanan. A quantum interpretation of separating conjunction for local reasoning of quantum programs based on separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 36:1–36:27, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498697.

Li:2022:LTL

[LLZ^{+}22] Jialin Li, Andrea Lattuada, Yi Zhou, Jonathan Cameron, Jon Howell, Bryan Parno, and Chris Hawblitzel. Linear types for large-scale systems verification. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):69:1–69:28, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527313.

Liu:2017:VDJ

[LMM17a] Lun Liu, Todd Millstein, and Madanlal Musuvathi. A volatile-by-default JVM for server applications. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):49:1–49:??, October 2017. CODEN ???? ISSN 2475-1421.

Lopes:2017:DMC

[LMM^{+}17b] Cristina V. Lopes, Petr Maj, Pedro Martins, Vaibhav Saini, Di Yang, Jakub Zitny, Hitesh Sajnani, and Jan Vitek. DéjàVu: a map of code duplicates on GitHub. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):84:1–84:??, October 2017. CODEN ???? ISSN 2475-1421.

Liu:2022:BSG

[LMÖ^{+}22] Si Liu, Jose Meseguer, Peter Csaba Ölveczky, Min Zhang, and David Basin. Bridging the semantic gap between qualitative and quantitative models of distributed systems. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):136:1–136:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563299.

Loding:2018:FNP

[LMP18] Christof Löding, P. Madhusudan, and Lucas Peña. Foundations for natural proofs and quantifier instantiation. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):10:1–10:??, January 2018. CODEN ???? ISSN 2475-1421.

Lindenhovius:2019:MLN

[LMZ19] Bert Lindenhovius, Michael Mislove, and Vladimir Zamdzhiev. Mixed linear and nonlinear recursive types. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):111:1–111:29, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341715.

Lim:2022:OPA

[LN22] Jay P. Lim and Santosh Nagarakatte. One polynomial approximation to produce correctly rounded results of an elementary function for multiple representations and rounding modes. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 3:1–3:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498664.

Lahav:2021:MWM

[LNO^{+}21] Ori Lahav, Egor Namakonov, Jonas Oberhauser, Anton Podkopaev, and Viktor Vafeiadis. Making weak memory models fair. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):98:1– 98:27, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485475.

Leobas:2020:SOD

[LP20] Guilherme Vieira Leobas and Fernando Magno Quintão Pereira. Semiring optimizations: dynamic elision of expressions with identity and absorbing elements. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):131:1–131:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428199.

Laddad:2022:KSC

[LPM^{+}22] Shadaj Laddad, Conor Power, Mae Milano, Alvin Cheung, and Joseph M. Hellerstein. Katara: synthesizing CRDTs with verified lifting. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):173:1–173:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563336.

Lampropoulos:2018:GGG

[LPP18] Leonidas Lampropoulos, Zoe Paraskevopoulou, and Benjamin C. Pierce. Generating good generators for inductive relations. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 45:1–45:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Liu:2020:VRD

[LPR^{+}20] Yiyun Liu, James Parker, Patrick Redmond, Lindsey Kuper, Michael Hicks, and Niki Vazou. Verifying replicated data types with typeclass refinements in Liquid Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA): 216:1–216:30, November 2020.

URL https://dl.acm.org/
doi/10.1145/3428284.

Lewchenko:2019:SPR

[LRGC19] Nicholas V. Lewchenko, Arjun Radhakrishna, Akash Gaonkar, and Pavol Cerný. Sequential programming for replicated data stores. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):106:1–106:28, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341710.

Liu:2020:VTF

[LRS^{+}20] Mengqi Liu, Lionel Rieg, Zhong Shao, Ronghui Gu, David Costanzo, Jung-Eun Kim, and Man-Ki Yoon. Virtual timeline: a formal abstraction for verifying preemptive schedulers with temporal isolation. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 20:1–20:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371088.

Le:2022:FRB

[LRV^{+}22] Quang Loc Le, Azalea Raad, Jules Villard, Josh Berdine, Derek Dreyer, and Peter W. O’Hearn. Finding real bugs in big programs with incorrectness logic. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):81:1–81:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527325.

Lee:2023:SAP

[LRY23] Wonyeol Lee, Xavier Rival, and Hongseok Yang. Smoothness analysis for probabilistic programs with application to optimised variational inference. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 12:1–12:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571205.

Li:2023:AFT

[LS23] Shaohua Li and Zhendong Su. Accelerating fuzzing through prefix-guided execution. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):75:1– 75:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586027.

Lee:2018:APC

[LSA18] Wonyeol Lee, Rahul Sharma, and Alex Aiken. On automatically proving the correctness of math.h implementations. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 47:1–47:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Liu:2022:CVT

[LSC^{+}22] Mengqi Liu, Zhong Shao, Hao Chen, Man-Ki Yoon, and Jung-Eun Kim. Compositional virtual timelines: verifying dynamic-priority partitions with algorithmic temporal isolation. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):127:1–127:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563290.

Lei:2022:TTR

[LSDZ22] Yuxiang Lei, Yulei Sui, Shuo Ding, and Qirun Zhang. Taming transitive redundancy for context-free language reachability. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):180:1– 180:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563343.

Lepigre:2022:VVR

[LSM^{+}22] Rodolphe Lepigre, Michael Sammler, Kayvan Memarian, Robbert Krebbers, Derek Dreyer, and Peter Sewell. VIP: verifying real-world C idioms with integer-pointer casts. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 20:1–20:32, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498681.

Lee:2018:ADC

[LSSO18] Junho Lee, Dowon Song, Sunbeom So, and Hakjoo Oh. Automatic diagnosis and correction of logical errors for functional programming assignments. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):158:1–158:30, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276528.

Liu:2020:WSF

[LSZ^{+}20] Hongyu Liu, Sam Silvestro, Xiangyu Zhang, Jian Huang, and Tongping Liu. WATCHER: insitu failure diagnosis. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):143:1–143:27, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428211.

Li:2022:EAD

[LSZ22] Yuanbo Li, Kris Satya, and Qirun Zhang. Efficient algorithms for dynamic bidirected Dyck-reachability. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):62:1–62:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498724.

Li:2018:PGC

[LTMS18] Yue Li, Tian Tan, Anders Møller, and Yannis Smaragdakis. Precision-guided context sensitivity for pointer analysis. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):141:1– 141:29, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276511.

Labrada:2022:PSG

[LTTD22] Elizabeth Labrada, Matías Toro, Éric Tanter, and Dominique Devriese. Plausible sealing for gradual parametricity. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):70:1– 70:28, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527314.

Li:2022:VCQ

[LVH^{+}22] Liyi Li, Finn Voichick, Kesha Hietala, Yuxiang Peng, Xiaodi Wu, and Michael Hicks. Verified compilation of quantum oracles. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):146:1–146:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563309.

Li:2023:TPD

[LVSZ23] Jianlin Li, Leni Ven, Pengyuan Shi, and Yizhou Zhang. Typepreserving, dependence-aware guide generation for sound, effective amortized probabilistic inference. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 50:1–50:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571243.

Li:2022:PAT

[LW22a] Yao Li and Stephanie Weirich. Program adverbs and Tlön embeddings. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 101:1–101:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547632.

Loehr:2022:SMP

[LW22b] Devon Loehr and David Walker. Safe, modular packet pipeline programming. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 38:1–38:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498699.

Li:2019:IBD

[LWNN19] Yi Li, Shaohua Wang, Tien N. Nguyen, and Son Van Nguyen. Improving bug detection via context-based code representation learning and attentionbased neural networks. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):162:1–162:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360588.

Lanzinger:2021:SPC

[LWUD21] Florian Lanzinger, Alexander Weigl, Mattias Ulbrich, and Werner Dietl. Scalability and precision by combining expressive type systems and deductive verification. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):143:1–143:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485520.

Lam:2020:LSL

[LWW^{+}20] Wing Lam, Stefan Winter, Anjiang Wei, Tao Xie, Darko Marinov, and Jonathan Bell. A large-scale longitudinal study of flaky tests. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):202:1–202:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428270.

Liu:2022:CGT

[LWY^{+}22] Jiawei Liu, Yuxiang Wei, Sen Yang, Yinlin Deng, and Lingming Zhang. Coverageguided tensor compiler fuzzing with joint IR-pass mutation. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):73:1– 73:26, April 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527317.

Lu:2019:PPY

[LX19] Jingbo Lu and Jingling Xue.

Precision-preserving yet fast object-sensitive pointer analysis with partial context sensitivity. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):148:1– 148:29, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360574.

Li:2018:AAT

[LY18] Yangjia Li and Mingsheng Ying. Algorithmic analysis of termination problems for quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 35:1–35:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Luan:2019:ACR

[LYB^{+}19] Sifei Luan, Di Yang, Celeste Barnaby, Koushik Sen, and Satish Chandra. Aroma: code recommendation via structural code search. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):152:1–152:28, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360578.

Lee:2020:TVS

[LYRY20] Wonyeol Lee, Hangyeol Yu, Xavier Rival, and Hongseok Yang. Towards verified stochastic variational inference for probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 16:1–16:33, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371084.

Laurel:2022:DNA

[LYSM22] Jacob Laurel, Rem Yang, Gagandeep Singh, and Sasa Misailovic. A dual number abstraction for static analysis of Clarke Jacobians. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):56:1–56:30, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498718.

Laurel:2022:GCA

[LYU^{+}22] Jacob Laurel, Rem Yang, Shubham Ugare, Robert Nagel, Gagandeep Singh, and Sasa Misailovic. A general construction for abstract interpretation of higher-order automatic differentiation. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):161:1–161:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563324.

Lesani:2022:CVT

[LyXK^{+}22] Mohsen Lesani, Li yao Xia, Anders Kaseorg, Christian J. Bell, Adam Chlipala, Benjamin C. Pierce, and Steve Zdancewic. C4: verified transactional objects. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):80:1–80:31, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527324.

Li:2021:RAG

[LyXW21] Yao Li, Li yao Xia, and Stephanie Weirich. Reasoning about the garden of forking paths. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 80:1–80:28, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473585.

Li:2017:TPT

[LZ17] Xia Li and Lingming Zhang. Transforming programs and tests in tandem for fault localization. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):92:1– 92:??, October 2017. CODEN ???? ISSN 2475-1421.

Li:2018:DFD

[LZ18] Peixuan Li and Danfeng Zhang. A derivation framework for dependent security label inference. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):115:1–115:26, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276485.

Li:2021:CBI

[LZR21] Yuanbo Li, Qirun Zhang, and Thomas Reps. On the complexity of bidirected interleaved Dyck-reachability. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):59:1–59:28, January

                                        2021. URL https://dl.acm.
                                        org/doi/10.1145/3434340.

Li:2023:SSS

[LZR23] Yuanbo Li, Qirun Zhang, and Thomas Reps. Single-source-single-target interleaved-Dyck reachability via integer linear programming. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):35:1–35:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571228.

Li:2020:PBR

[LZY^{+}20] Gushu Li, Li Zhou, Nengkun Yu, Yufei Ding, Mingsheng Ying, and Yuan Xie. Projectionbased runtime assertions for testing and debugging quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):150:1–150:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428218.

Mendis:2018:GGO

[MA18] Charith Mendis and Saman Amarasinghe. goSLP: globally optimized superword level parallelism framework. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):110:1–110:28, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276480.

Maillard:2019:DMA

[MAA^{+}19] Kenji Maillard, Danel Ahman, Robert Atkey, Guido Martínez, Cătălin Hriţcu, Exequiel Rivas, and Éric Tanter. Dijkstra monads for all. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):104:1–104:29, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341708.

Muller:2018:CPG

[MAH18] Stefan K. Muller, Umut A. Acar, and Robert Harper. Competitive parallelism: getting your priorities right. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):95:1–95:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236790.

Mainland:2017:BLT

[Mai17] Geoffrey Mainland. Better living through operational semantics: an optimizing compiler for radio protocols. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):19:1–19:??, September 2017. CODEN ???? ISSN 2475-1421.

Mansky:2017:VMS

[MAN17] William Mansky, Andrew W. Appel, and Aleksey Nogin. A verified messaging system. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):87:1– 87:??, October 2017. CODEN ???? ISSN 2475-1421.

Materzok:2022:GCG

[Mat22]  Marek Materzok.  Gener-

ating circuits with generators. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 92:1–92:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549821.

Mariano:2022:ATI

[MCF^{+}22] Benjamin Mariano, Yanju Chen, Yu Feng, Greg Durrett, and Isil Dillig. Automated transpilation of imperative to functional code using neural-guided program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):71:1–71:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527315.

Maeng:2017:AIE

[MCL17] Kiwan Maeng, Alexei Colin, and Brandon Lucia. Alpaca: intermittent execution without checkpoints. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):96:1–96:??, October 2017. CODEN ???? ISSN 2475-1421.

Merigoux:2021:PCP

[MCP21] Denis Merigoux, Nicolas Chataing, and Jonathan Protzenko. Catala: a programming language for the law. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 77:1–77:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473582.

Moine:2023:HLS

[MCP23] Alexandre Moine, Arthur Charguéraud, and François Pottier. A high-level separation logic for heap space under garbage collection. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 25:1–25:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571218.

Mukherjee:2020:LBC

[MDBL20] Suvam Mukherjee, Pantazis Deligiannis, Arpita Biswas, and Akash Lal. Learning-based controlled concurrency testing. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):230:1– 230:31, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428298.

Mogk:2019:FTP

[MDSM19] Ragnar Mogk, Joscha Drechsler, Guido Salvaneschi, and Mira Mezini. A fault-tolerant programming model for distributed interactive applications. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):144:1– 144:29, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360570.

Morris:2017:CTF

[ME17] J. Garrett Morris and Richard A.

Eisenberg. Constrained type families. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 42:1–42:??, September 2017. CODEN ???? ISSN 2475-1421.

Mellies:2019:CCS

[Mel19] Paul-André Mellies. Categorical combinatorics of scheduling and synchronization in game semantics. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 23:1–23:30, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290336.

Mackay:2022:NSR

[MEND22] Julian Mackay, Susan Eisenbach, James Noble, and Sophia Drossopoulou. Necessity specifications for robustness. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):154:1–154:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563317.

Miltner:2018:SBL

[MFP^{+}18] Anders Miltner, Kathleen Fisher, Benjamin C. Pierce, David Walker, and Steve Zdancewic. Synthesizing bijective lenses. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 1:1–1:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Michael:2023:MSE

[MGB^{+}23] Alexandra E. Michael, Anitha Gollamudi, Jay Bosamiya, Evan Johnson, Aidan Denlinger, Craig Disselkoen, Conrad Watt, Bryan Parno, Marco Patrignani, Marco Vassena, and Deian Stefan. MSWasm: Soundly enforcing memory-safe execution of unsafe code. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):15:1–15:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571208.

Miltner:2019:FSE

[MGL^{+}19] Anders Miltner, Sumit Gulwani, Vu Le, Alan Leung, Arjun Radhakrishna, Gustavo Soares, Ashish Tiwari, and Abhishek Udupa. On the fly synthesis of edit suggestions. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):143:1–143:29, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360569.

Malewski:2021:GSD

[MGT21] Stefan Malewski, Michael Greenberg, and Éric Tanter. Gradually structured data. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):126:1– 126:29, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485503.

Muller:2021:MAE

[MH21] Stefan K. Muller and Jan Hoffmann. Modeling and analyzing evaluation cost of CUDA kernels. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 25:1–25:31, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434306.

Menz:2023:CSD

[MHLG23] Jan Menz, Andrew K. Hirsch, Peixuan Li, and Deepak Garg. Compositional security definitions for higher-order where declassification. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):89:1–89:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586041.

Mastrangelo:2019:CAD

[MHN19] Luis Mastrangelo, Matthias Hauswirth, and Nathaniel Nystrom. Casting about in the dark: an empirical study of cast operations in Java programs. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):158:1–158:31, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360584.

MacQueen:2020:HSM

[MHR20] David MacQueen, Robert Harper, and John Reppy. The history of Standard ML. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):86:1–86:100, June

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3386336.

Maillard:2020:NRP

[MHRV20] Kenji Maillard, Catalin Hritcu, Exequiel Rivas, and Antoine Van Muylder. The next 700 relational program logics. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):4:1–4:33, January 2020. URL https://

                                        dl.acm.org/doi/abs/10.1145/
                                        3371072.

Milanova:2020:FGT

[Mil20]  Ana Milanova.  FlowCFL:

generalized type-based reachability analysis: graph reduction and equivalence of CFL-based and type-based reachability. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):178:1–178:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428246.

Montagu:2020:SRA

[MJ20] Benoît Montagu and Thomas Jensen. Stable relations and abstract interpretation of higherorder programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):119:1–119:30, August

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3409001.

Mevel:2021:FVC

[MJ21] Glen Mével and Jacques-Henri Jourdan. Formal verification of a concurrent bounded queue in a weak memory model. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):66:1–66:29, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3473571.

Mishra:2022:SGC

[MJ22] Ashish Mishra and Suresh Jagannathan. Specificationguided component-based synthesis from effectful libraries. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):147:1–147:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563310.

Midtgaard:2017:EDQ

[MJK^{+}17] Jan Midtgaard, Mathias Nygaard Justesen, Patrick Kasting, Flemming Nielson, and Hanne Riis Nielson. Effectdriven QuickChecking of compilers. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 15:1–15:??, September 2017. CODEN ???? ISSN 2475-1421.

Mevel:2020:CCS

[MJP20] Glen Mével, Jacques-Henri Jourdan, and François Pottier. Cosmo: a concurrent separation logic for multicore OCaml. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 96:1–96:29, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408978.

Mulder:2023:PAL

[MK23] Ike Mulder and Robbert Krebbers. Proof automation for linearizability in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):91:1–91:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586043.

Mayer:2018:BED

[MKC18] Mikaël Mayer, Viktor Kuncak, and Ravi Chugh. Bidirectional evaluation with direct manipulation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):127:1–127:28, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276497.

Mukherjee:2020:DBP

[MKLR20] Manasij Mukherjee, Pranav Kant, Zhengyang Liu, and John Regehr. Dataflow-based pruning for speeding up superoptimization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):177:1–177:24, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428245.

Machiry:2022:CCC

[MKM^{+}22] Aravind Machiry, John Kastner, Matt McCutchen, Aaron Eline, Kyle Headley, and Michael Hicks. C to checked C by 3C. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):78:1–78:29, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527322.

Mazinanian:2017:UUL

[MKTD17] Davood Mazinanian, Ameya Ketkar, Nikolaos Tsantalis, and Danny Dig. Understanding the use of lambda expressions in Java. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):85:1– 85:??, October 2017. CODEN ???? ISSN 2475-1421.

Mathur:2018:WHA

[MKV18] Umang Mathur, Dileep Kini, and Mahesh Viswanathan. What happens — after the first race? Enhancing the predictive power of happensbefore based dynamic race detection. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):145:1– 145:29, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276515.

Moiseenko:2022:MCM

[MKV22] Evgenii Moiseenko, Michalis Kokologiannakis, and Viktor Vafeiadis. Model checking for a multi-execution memory model. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):152:1–152:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563315.

Madsen:2020:FMP

[ML20a] Magnus Madsen and Ondrej Lhoták. Fixpoints for the masses: programming with first-class Datalog constraints. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA): 125:1–125:28, November 2020.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428193.

Moler:2020:HM

[ML20b] Cleve Moler and Jack Little. A history of MATLAB. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):81:1–81:67, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386331.

Margalit:2021:VOR

[ML21] Roy Margalit and Ori Lahav. Verifying observational robustness against a C11-style memory model. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 4:1–4:33, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434285.

Maillard:2022:RGT

[MLBTT22] Kenji Maillard, Meven Lennon-Bertrand, Nicolas Tabareau, and Éric Tanter. A reasonably gradual type theory. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):124:1–124:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547655.

Mokhov:2019:SAF

[MLMD19] Andrey Mokhov, Georgy Lukyanov, Simon Marlow, and Jeremie Dimino. Selective applicative functors. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):90:1–90:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341694.

Madsen:2017:MRA

[MLT17] Magnus Madsen, Ondrej Lhoták, and Frank Tip. A model for reasoning about JavaScript promises. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA): 86:1–86:??, October 2017. CO-DEN ???? ISSN 2475-1421.

Morris:2019:AED

[MM19] J. Garrett Morris and James McKinna. Abstracting extensible data types: or, rows by any other name. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 12:1–12:28, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290325.

Maina:2018:SQL

[MMF^{+}18] Solomon Maina, Anders Miltner, Kathleen Fisher, Benjamin C. Pierce, David Walker, and Steve Zdancewic. Synthesizing quotient lenses. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):80:1–80:29, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236775.

Miltner:2019:SSL

[MMF^{+}19] Anders Miltner, Solomon Maina, Kathleen Fisher, Benjamin C. Pierce, David Walker, and Steve Zdancewic. Synthesizing symmetric lenses. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):95:1–95:28, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341699.

Mokhov:2018:BSC

[MMJ18] Andrey Mokhov, Neil Mitchell, and Simon Peyton Jones. Build systems à la carte. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):79:1–79:29, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236774.

Mathur:2020:DMS

[MMK^{+}20] Umang Mathur, Adithya Murali, Paul Krogmeier, P. Madhusudan, and Mahesh Viswanathan. Deciding memory safety for single-pass heap-manipulating programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 35:1–35:29, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371103.

McIver:2018:NPR

[MMKK18] Annabelle McIver, Carroll Morgan, Benjamin Lucien Kaminski, and Joost-Pieter Katoen. A new proof rule for almostsure termination. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 33:1–33:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Muller:2022:PGP

[MMS^{+}22] Mark Niklas Müller, Gleb Makarchuk, Gagandeep Singh, Markus Püschel, and Martin Vechev. PRIMA: general and precise neural network certification via scalable convex hull approximations. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):43:1–43:33, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498704.

Majumdar:2018:WRT

[MN18] Rupak Majumdar and Filip Niksic. Why is random testing effective for partition tolerance bugs? Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 46:1–46:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Miltner:2022:BSR

[MNB^{+}22] Anders Miltner, Adrian Trejo Nuñez, Ana Brendel, Swarat Chaudhuri, and Isil Dillig. Bottom-up synthesis of recursive functional programs using angelic execution. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):21:1–21:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498682.

Moller:2020:DLJ

[MNT20] Anders Møller, Benjamin Barslev Nielsen, and Martin Toldam Torp. Detecting locations in JavaScript programs affected by breaking library changes. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA): 187:1–187:25, November 2020.

                                        URL https://dl.acm.org/
                                        doi/10.1145/3428255.

Moy:2021:CRS

[MNTHV21] Cameron Moy, Phúc C. Nguy˜ ên, Sam Tobin-Hochstadt, and David Van Horn. Corpse reviver: sound and efficient gradual typing via contract verification. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 53:1–53:28, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434334.

Moreira:2021:PVS

[MOP21] Angélica Aparecida Moreira, Guilherme Ottoni, and Fernando Magno Quintão Pereira. VESPA: static profiling for binary optimization. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):144:1–144:28, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485521.

Morihata:2019:LCA

[Mor19] Akimasa Morihata. Lambda calculus with algebraic simplification for reduction parallelization by equational reasoning. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP): 80:1–80:25, July 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3341644.

Melgratti:2017:CCH

[MP17] Hernán Melgratti and Luca Padovani. Chaperone contracts for higher-order sessions. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 35:1–35:??, September 2017. CODEN ???? ISSN 2475-1421.

Migeed:2020:WDA

[MP20] Zeina Migeed and Jens Palsberg. What is decidable about gradual types? Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 29:1–29:29, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371097.

Mathiasen:2021:FGP

[MP21a] Anders Alnor Mathiasen and Andreas Pavlogiannis. The fine-grained and parallel complexity of Andersen’s pointer analysis. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 34:1–34:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434315.

Mazza:2021:ADP

[MP21b] Damiano Mazza and Michele Pagani. Automatic differentiation in PCF. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):28:1–28:27, January

                                        2021. URL https://dl.acm.
                                        org/doi/10.1145/3434309.

Madiot:2022:SLH

[MP22] Jean-Marie Madiot and François Pottier. A separation logic for heap space under garbage collection. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 11:1–11:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498672.

Mackay:2020:DSP

[MPAG20] Julian Mackay, Alex Potanin, Jonathan Aldrich, and Lindsay Groves. Decidable subtyping for path dependent types. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):

66:1–66:27, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371134.

Murali:2022:MGS

[MPB^{+}22] Adithya Murali, Lucas Peña, Eion Blanchard, Christof Löding, and P. Madhusudan. Modelguided synthesis of inductive lemmas for FOL with least fixpoints. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):191:1–191:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563354.

Mazza:2018:PAF

[MPV18] Damiano Mazza, Luc Pellissier, and Pierre Vial. Polyadic approximations, fibrations and intersection types. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 6:1–6:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Mathur:2021:OPS

[MPV21] Umang Mathur, Andreas Pavlogiannis, and Mahesh Viswanathan. Optimal prediction of synchronizationpreserving races. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 36:1–36:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434317.

Muduli:2022:SMF

[MR22] Sujit Kumar Muduli and Subhajit Roy. Satisfiability modulo fuzzing: a synergistic combination of SMT solving and fuzzing. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):169:1–169:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563332.

Melo:2018:ISS

[MRdAP18] Leandro T. C. Melo, Rodrigo G. Ribeiro, Marcus R. de Araújo, and Fernando Magno Quintão Pereira. Inference of static semantics for incomplete C programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 29:1–29:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Meier:2018:VMD

[MRG18] Remigius Meier, Armin Rigo, and Thomas R. Gross. Virtual machine design for parallel dynamic programming languages. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):109:1– 109:25, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276479.

Mariano:2019:PSA

[MRX^{+}19] Benjamin Mariano, Josh Reese, Siyuan Xu, ThanhVu Nguyen, Xiaokang Qiu, Jeffrey S. Foster, and Armando Solar-Lezama. Program synthesis with algebraic library specifications. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):132:1–132:25, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360558.

Miraldo:2019:EAT

[MS19] Victor Cacciari Miraldo and Wouter Swierstra. An efficient algorithm for type-safe structural diffing. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):113:1–113:29, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341717.

Monnier:2020:EEL

[MS20] Stefan Monnier and Michael Sperber. Evolution of Emacs Lisp. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):74:1– 74:55, June 2020. URL https:/

/dl.acm.org/doi/abs/10.1145/
3386324.

Moosbrugger:2022:MPL

[MSBK22] Marcel Moosbrugger, Miroslav Stankovič, Ezio Bartocci, and Laura Kovács. This is the moment for probabilistic loops. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):178:1– 178:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563341.

Muller:2023:BDS

[MSBO23] Marius Müller, Philipp Schuster, Jonathan Immanuel Brachthäuser, and Klaus Ostermann. Back to direct style: Typed and tight. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):104:1– 104:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3586056.

Marntirosian:2020:RIS

[MSdSOK20] Koar Marntirosian, Tom Schrijvers, Bruno C. d. S. Oliveira, and Georgios Karachalias. Resolution as intersection subtyping via modus ponens. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):206:1–206:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428274.

Miyazaki:2019:DTI

[MSI19] Yusuke Miyazaki, Taro Sekiyama, and Atsushi Igarashi. Dynamic type inference for gradual Hindley–Milner typing. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):18:1–18:29, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290331.

Meuli:2020:EAA

[MSRH20] Giulia Meuli, Mathias Soeken, Martin Roetteler, and Thomas Häner. Enabling accuracyaware quantum compilers using symbolic resource estimation. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):130:1– 130:26, November 2020. URL

https://dl.acm.org/doi/10.
1145/3428198.

Muehlboeck:2017:SGT

[MT17] Fabian Muehlboeck and Ross Tate. Sound gradual typing is nominally alive and well. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):56:1– 56:??, October 2017. CODEN ???? ISSN 2475-1421.

Muehlboeck:2018:EUI

[MT18] Fabian Muehlboeck and Ross Tate. Empowering union and intersection types with integrated subtyping. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):112:1–112:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276482.

Muehlboeck:2021:TSN

[MT21] Fabian Muehlboeck and Ross Tate. Transitioning from structural to nominal code with efficient gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):127:1–127:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485504.

Marcozzi:2019:CFH

[MTDC19] Michaël Marcozzi, Qiyi Tang, Alastair F. Donaldson, and Cristian Cadar. Compiler fuzzing: how much does it matter? Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):155:1–155:29, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360581.

Muller:2022:SPP

[Mul22] Stefan K. Muller. Static prediction of parallel computation graphs. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 46:1–46:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498708.

Mogelberg:2019:BPT

[MV19] Rasmus Ejlers Møgelberg and Niccolò Veltri. Bisimulation as path type for guarded recursive types. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):4:1–4:29, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290317.

Moller:2020:EAO

[MV20] Anders Møller and Oskar Haarklou Veileborg. Eliminating abstraction overhead of Java stream pipelines using ahead-of-time program optimization. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):168:1–168:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428236.

Madsen:2020:PTE

[MvdP20] Magnus Madsen and Jaco van de Pol. Polymorphic types and effects with Boolean unification. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):154:1–154:29, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428222.

Madsen:2021:RNT

[MvdP21] Magnus Madsen and Jaco van de Pol. Relational nullable types with Boolean unification. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):110:1– 110:28, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485487.

Matsuda:2020:SLP

[MW20a] Kazutaka Matsuda and Meng Wang. Sparcl: a language for partially-invertible computation. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):118:1– 118:31, August 2020. URL

https://dl.acm.org/doi/10.
1145/3409000.

Meyer:2020:PLC

[MW20b] Roland Meyer and Sebastian Wolff. Pointer life cycle types for lock-free data structures with memory reclamation. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 68:1–68:36, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371136.

Muller:2019:FRP

[MWA19] Stefan K. Muller, Sam Westrick, and Umut A. Acar. Fairness in responsive parallelism. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):81:1–81:30, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341685.

Meyer:2022:CPL

[MWW22] Roland Meyer, Thomas Wies, and Sebastian Wolff. A concurrent program logic with a future and history. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):174:1–174:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563337.

Majumdar:2020:MMC

[MYZ20] Rupak Majumdar, Nobuko Yoshida, and Damien Zufferey. Multiparty motion coordination: from choreographies to robotics programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):134:1–134:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428202.

New:2018:GEP

[NA18] Max S. New and Amal Ahmed. Graduality from embeddingprojection pairs. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):73:1–73:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236768.

Newcomb:2020:VIH

[NAJ^{+}20] Julie L. Newcomb, Andrew Adams, Steven Johnson, Rastislav Bodik, and Shoaib Kamil. Verifying and improving Halide’s term rewriting system with program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):166:1–166:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428234.

Nanevski:2019:SCP

[NBDF19] Aleksandar Nanevski, Anindya Banerjee, Germán Andrés Delbianco, and Ignacio Fábregas. Specifying concurrent programs in separation logic: morphisms and simulations. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):161:1–161:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360587.

Nardelli:2018:JSR

[NBP^{+}18] Francesco Zappa Nardelli, Julia Belyakova, Artem Pelenitsyn, Benjamin Chung, Jeff Bezanson, and Jan Vitek. Julia subtyping: a rational reconstruction. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):113:1–113:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276483.

Narayanan:2017:SCA

[NcS17] Praveen Narayanan and Chung chieh Shan. Symbolic conditioning of arrays in probabilistic programs. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):11:1–11:??, September 2017. CODEN ???? ISSN 2475-1421.

Near:2019:DEH

[NDA^{+}19] Joseph P. Near, David Darais, Chike Abuah, Tim Stevens, Pranav Gaddamadugu, Lun Wang, Neel Somani, Mu Zhang, Nikhil Sharma, Alex Shan, and Dawn Song. Duet: an expressive higher-order language and linear type system for statically enforcing differential privacy. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):172:1–172:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360598.

Nieto:2022:MVO

[NGR^{+}22] Abel Nieto, Léon Gondelman, Alban Reynaud, Amin Timany, and Lars Birkedal. Modular verification of op-based CRDTs in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):188:1–188:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563351.

Nguyen:2018:SCV

[NGTHV18] Phúc C. Nguy˜ ên, Thomas Gilray, Sam Tobin-Hochstadt, and David Van Horn. Soft contract verification for higherorder stateful programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):51:1–51:??, January 2018. CODEN ???? ISSN 2475-1421.

New:2020:GPT

[NJA20] Max S. New, Dustin Jamner, and Amal Ahmed. Graduality and parametricity: together again for the first time. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 46:1–46:32, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371114.

New:2019:GTT

[NLA19] Max S. New, Daniel R. Licata, and Amal Ahmed. Gradual type theory. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 15:1–15:31, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290328.

Nguyen:2022:MPM

[NPWW22] Minh Nguyen, Roly Perera, Meng Wang, and Nicolas Wu. Modular probabilistic models via algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):104:1–104:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547635.

Nie:2020:UEI

[NPZ^{+}20] Pengyu Nie, Marinela Parovic, Zhiqiang Zang, Sarfraz Khurshid, Aleksandar Milicevic, and Milos Gligoric. Unifying execution of imperative generators and declarative specifications. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):217:1–217:26, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428285.

Niu:2022:CAL

[NSGH22] Yue Niu, Jonathan Sterling, Harrison Grodin, and Robert Harper. A cost-aware logical framework. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 9:1–9:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498670.

Nuyts:2017:PQD

[NVD17] Andreas Nuyts, Andrea Vezzosi, and Dominique Devriese. Parametric quantifiers for dependent type theory. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):32:1–32:??, September 2017. CODEN ???? ISSN 2475-1421.

Nandi:2018:FPC

[NWP^{+}18] Chandrakana Nandi, James R.

Wilcox, Pavel Panchekha, Taylor Blau, Dan Grossman, and Zachary Tatlock. Functional programming for compiling and decompiling computeraided design. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):99:1–99:31, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236794.

Nandi:2021:RRI

[NWZ^{+}21] Chandrakana Nandi, Max Willsey, Amy Zhu, Yisu Remy Wang, Brett Saiki, Adam Anderson, Adriana Schulz, Dan Grossman, and Zachary Tatlock. Rewrite rule inference using equality saturation. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):119:1– 119:28, October 2021. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485496.

Omar:2018:RPL

[OA18] Cyrus Omar and Jonathan Aldrich. Reasonably programmable literal notation. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):106:1–106:32, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236801.

Odersky:2018:SFA

[OBL^{+}18] Martin Odersky, Olivier Blanvillain, Fengyun Liu, Aggelos Biboudis, Heather Miller, and Sandro Stucki. Simplicitly: foundations and applications of implicit function types. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 42:1–42:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Ostermann:2022:IEL

[OBS^{+}22] Klaus Ostermann, David Binder, Ingo Skupin, Tim Süberkrüb, and Paul Downen. Introduction and elimination, left and right. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 106:1–106:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547637.

Orchard:2017:VSP

[OCDR17] Dominic Orchard, Mistral Contrastin, Matthew Danish, and Andrew Rice. Verifying spatial properties of array computations. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):75:1– 75:??, October 2017. CODEN ???? ISSN 2475-1421.

Ou:2018:TUC

[OD18] Peizhao Ou and Brian Demsky. Towards understanding the costs of avoiding outof-thin-air results. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):136:1–136:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276506.

OHearn:2020:IL

[O’H20] Peter W. O’Hearn. Incorrectness logic. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 10:1–10:32, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371078.

Ott:2018:BPS

[OLC^{+}18] Jason Ott, Tyson Loveless, Chris Curtis, Mohsen Lesani, and Philip Brisk. BioScript: programming safe chemistry on laboratories-on-a-chip. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):128:1–128:31, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276498.

Orchard:2019:QPR

[OLE19] Dominic Orchard, Vilem-Benjamin Liepelt, and Harley Eades III. Quantitative program reasoning with graded modal types. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):110:1– 110:30, July 2019. URL https:

//dl.acm.org/doi/abs/10.1145/
3341714.

Ozkan:2018:RTD

[OMN^{+}18] Burcu Kulahcioglu Ozkan, Rupak Majumdar, Filip Niksic, Mitra Tabaei Befrouei, and Georg Weissenbacher. Randomized testing of distributed systems with probabilistic guarantees. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):160:1– 160:28, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276530.

Ozkan:2019:TAR

[OMO19] Burcu Kulahcioglu Ozkan, Rupak Majumdar, and Simin Oraee. Trace aware random testing for distributed systems. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):180:1– 180:29, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360606.

Ohman:2022:VRC

[ÖN22] Joakim Öhman and Aleksandar Nanevski. Visibility reasoning for concurrent snapshot algorithms. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 33:1–33:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498694.

Owens:2017:VEF

[ONK^{+}17] Scott Owens, Michael Norrish, Ramana Kumar, Magnus O. Myreen, and Yong Kiam Tan. Verifying efficient function calls in CakeML. Proceedings of the ACM on Programming Languages (PACMPL), 1 (ICFP):18:1–18:??, September 2017. CODEN ???? ISSN 2475-

Ohori:2018:FPO

[OUM18] Atsushi Ohori, Katsuhiro Ueno, and Hisayuki Mima. Finitary polymorphism for optimizing type-directed compilation. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP): 81:1–81:29, July 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3236776.

Omar:2019:LFP

[OVCH19] Cyrus Omar, Ian Voysey, Ravi Chugh, and Matthew A. Hammer. Live functional programming with typed holes. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):14:1–14:32, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290327.

Paraskevopoulou:2019:CCS

[PA19a] Zoe Paraskevopoulou and Andrew W. Appel. Closure conversion is safe for space. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):83:1–83:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341687.

Patterson:2019:NCC

[PA19b] Daniel Patterson and Amal Ahmed. The next 700 compiler correctness theorems (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):85:1–85:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341689.

Parreaux:2020:SEA

[Par20] Lionel Parreaux. The simple essence of algebraic subtyping: principal type inference with subtyping made easy (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):124:1–124:28, August

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3409006.

Pavlogiannis:2020:FSE

[Pav20] Andreas Pavlogiannis. Fast, sound, and effectively complete dynamic race prediction. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 17:1–17:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371085.

Pelenitsyn:2021:TSJ

[PBC^{+}21] Artem Pelenitsyn, Julia Belyakova, Benjamin Chung, Ross Tate, and Jan Vitek. Type stability in Julia: avoiding performance pathologies in JIT compilation. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):150:1–150:26, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485527.

Papadakis:2017:SPS

[PBS^{+}17] Manolis Papadakis, Gilbert Louis Bernstein, Rahul Sharma, Alex Aiken, and Pat Hanrahan. Seam: provably safe local edits on graphs. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):78:1–78:??, October 2017. CODEN ???? ISSN 2475-1421.

Parreaux:2022:MPT

[PC22] Lionel Parreaux and Chun Yin Chau. MLstruct: principal type inference in a Boolean algebra of structural types. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):141:1–141:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563304.

Phipps-Costin:2021:SBG

[PCAGG21] Luna Phipps-Costin, Carolyn Jane Anderson, Michael Greenberg, and Arjun Guha. Solver-based gradual type migration. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA): 111:1–111:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3485488.

Palmkvist:2023:SRA

[PCHB23] Viktor Palmkvist, Elias Castegren, Philipp Haller, and David Broman. Statically resolvable ambiguity. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 58:1–58:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571251.

Perez:2018:FTF

[Per18] Ivan Perez. Fault tolerant functional reactive programming (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):96:1–96:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236791.

Pacak:2020:SAD

[PES20] André Pacak, Sebastian Erdweg, and Tamás Szabó. A systematic approach to deriving incremental type checkers. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):127:1– 127:28, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428195.

Panchekha:2019:MVW

[PETK19] Pavel Panchekha, Michael D. Ernst, Zachary Tatlock, and Shoaib Kamil. Modular verification of web page layout. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):151:1– 151:26, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360577.

Pulte:2018:SAC

[PFD^{+}18] Christopher Pulte, Shaked Flur, Will Deacon, Jon French, Susmit Sarkar, and Peter Sewell. Simplifying ARM concurrency: multicopy-atomic axiomatic and operational models for ARMv8. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 19:1–19:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Paraskevopoulou:2021:CCC

[PG21] Zoe Paraskevopoulou and Anvay Grover. Compiling with continuations, correctly. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):114:1–114:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485491.

Poesia:2017:SPC

[PGFP17] Gabriel Poesia, Breno Guimarães, Fabrício Ferracioli, and Fernando Magno Quintão Pereira. Static placement of computation on heterogeneous devices. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):50:1– 50:??, October 2017. CODEN ???? ISSN 2475-1421.

Peleg:2020:PRE

[PGIY20] Hila Peleg, Roi Gabay, Shachar Itzhaky, and Eran Yahav. Programming with a read– eval–synth loop. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):159:1–159:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428227.

Pickard:2021:CDT

[PH21] Mitchell Pickard and Graham Hutton. Calculating dependently-typed compilers (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):82:1–82:27, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3473587.

Padon:2018:RLS

[PHL^{+}18] Oded Padon, Jochen Hoenicke, Giuliano Losa, Andreas Podelski, Mooly Sagiv, and Sharon Shoham. Reducing liveness to safety in first-order logic. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 26:1–26:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Park:2019:PSM

[PHSR19] Gyunghee Park, Jaemin Hong, Guy L. Steele Jr., and Sukyoung Ryu. Polymorphic symmetric multiple dispatch with variance. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 11:1–11:28, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290324.

Pan:2019:ARR

[PHXD19] Rong Pan, Qinheping Hu, Gaowei Xu, and Loris D’Antoni. Automatic repair of regular expressions. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):139:1–139:29, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360565.

Pitts:2023:LNS

[Pit23] Andrew M. Pitts. Locally nameless sets. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 17:1–17:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571210.

Paszke:2021:GPI

[PJD^{+}21] Adam Paszke, Daniel D.

Johnson, David Duvenaud, Dimitrios Vytiniotis, Alexey Radul, Matthew J. Johnson, Jonathan Ragan-Kelley, and Dougal Maclaurin. Getting to the point: index sets and parallelism-preserving autodiff for pointful array programming. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 88:1–88:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473593.

Padhi:2018:FFS

[PJP^{+}18] Saswat Padhi, Prateek Jain, Daniel Perelman, Oleksandr Polozov, Sumit Gulwani, and Todd Millstein. FlashProfile: a framework for synthesizing data profiles. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):150:1–150:28, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276520.

Patel:2021:VCM

[PKSW21] Nisarg Patel, Siddharth Krishna, Dennis Shasha, and Thomas Wies. Verifying concurrent multicopy search structures. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):113:1– 113:32, October 2021. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3485490.

Pombrio:2017:IST

[PKW17] Justin Pombrio, Shriram Krishnamurthi, and Mitchell Wand. Inferring scope through syntactic sugar. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 44:1–44:??, September 2017. CODEN ???? ISSN 2475-1421. Paraskevopoulou:2021:COC

[PLA21] Zoe Paraskevopoulou, John M. Li, and Andrew W. Appel. Compositional optimizations for CertiCoq. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 86:1–86:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473591.

Padhye:2019:FDS

[PLS^{+}19] Rohan Padhye, Caroline Lemieux, Koushik Sen, Laurent Simon, and Hayawardh Vijayakumar. FuzzFactory: domainspecific fuzzing with waypoints. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):174:1– 174:29, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360600.

Padon:2017:PME

[PLSS17] Oded Padon, Giuliano Losa, Mooly Sagiv, and Sharon Shoham. Paxos made EPR: decidable reasoning about distributed protocols. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):108:1–108:??, October 2017. CODEN ???? ISSN 2475-1421.

Patrignani:2021:SER

[PMD21] Marco Patrignani, Eric Mark Martin, and Dominique Devriese. On the semantic expressiveness of recursive types. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 21:1–21:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434302.

Pulte:2023:CVS

[PMS^{+}23] Christopher Pulte, Dhruv C.

Makwana, Thomas Sewell, Kayvan Memarian, Peter Sewell, and Neel Krishnaswami. CN: Verifying systems C code with separation-logic refinement types. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):

1:1–1:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571194.

Perez:2017:TDF

[PN17] Ivan Perez and Henrik Nilsson. Testing and debugging functional reactive programming. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 2:1–2:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Perera:2022:LVG

[PNPW22] Roly Perera, Minh Nguyen, Tomas Petricek, and Meng Wang. Linked visualisations via Galois dependencies. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):7:1–7:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498668.

Porncharoenwase:2022:FFS

[PNWT22] Sorawee Porncharoenwase, Luke Nelson, Xi Wang, and Emina Torlak. A formal foundation for symbolic evaluation with merging. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):47:1–47:28, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498709.

Pottier:2017:VU

[Pot17] François Pottier. Visitors unchained. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 28:1–28:??, September 2017. CODEN ???? ISSN 2475-1421.

Poesia:2020:DDC

[PP20] Gabriel Poesia and Fernando Magno Quintão Pereira. Dynamic dispatch of contextsensitive optimizations. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):167:1–167:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428235.

Paltenghi:2022:BQC

[PP22] Matteo Paltenghi and Michael Pradel. Bugs in quantum computing platforms: an empirical study. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):86:1–86:27, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527330.

Palmer:2020:HOD

[PPSW20] Zachary Palmer, Theodore Park, Scott Smith, and Shiwei Weng. Higher-order demand-driven symbolic evaluation. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):102:1– 102:28, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408984.

Poulsen:2018:ITD

[PRT^{+}18] Casper Bach Poulsen, Arjen Rouvoet, Andrew Tolmach, Robbert Krebbers, and Eelco Visser. Intrinsically-typed definitional interpreters for imperative languages. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 16:1–16:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Pradel:2018:DLA

[PS18] Michael Pradel and Koushik Sen. DeepBugs: a learning approach to name-based bug detection. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):147:1–147:25, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276517.

Pavlinovic:2021:DFR

[PSW21] Zvonimir Pavlinovic, Yusen Su, and Thomas Wies. Data flow refinement type inference. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 19:1–19:31, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434300.

Polikarpova:2020:LIF

[PSY^{+}20] Nadia Polikarpova, Deian Stefan, Jean Yang, Shachar Itzhaky, Travis Hance, and Armando Solar-Lezama. Liquid information flow control. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):105:1– 105:30, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408987.

Pedrot:2020:FTH

[PT20] Pierre-Marie Pédrot and Nicolas Tabareau. The fire triangle: how to mix substitution, dependent elimination, and effects. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 58:1–58:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371126.

Pujet:2022:OEN

[PT22] Loïc Pujet and Nicolas Tabareau.

Observational equality: now for good. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 32:1–32:27, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498693.

Popescu:2023:ATP

[PT23a] Andrei Popescu and Dmitriy Traytel. Admissible types-to-PERs relativization in higherorder logic. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 42:1–42:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571235.

Pujet:2023:IOE

[PT23b] Loïc Pujet and Nicolas Tabareau. Impredicative observational equality. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 74:1–74:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571739.

Pedrot:2019:RET

[PTFT19] Pierre-Marie Pédrot, Nicolas Tabareau, Hans Jacob Fehrmann, and Éric Tanter. A reasonably exceptional type theory. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):108:1–108:29, July

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3341712.

Puente:2017:PMR

[Pue17] Juan Pedro Bolívar Puente. Persistence for the masses: RRB-vectors in a systems language. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 16:1–16:??, September 2017. CODEN ???? ISSN 2475-1421.

Pitchanathan:2021:PFF

[PUW^{+}21] Arjun Pitchanathan, Christian Ulmann, Michel Weber, Torsten Hoefler, and Tobias Grosser. FPL: fast Presburger arithmetic through transprecision. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):162:1–162:26, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485539.

Pelsmaeker:2022:LPS

[PvAPV22] Daniel A. A. Pelsmaeker, Hendrik van Antwerpen, Casper Bach Poulsen, and Eelco Visser. Language-parametric static semantic code completion. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):85:1–85:30, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527329.

Poulsen:2023:HAM

[PvdR23] Casper Bach Poulsen and Cas van der Rest. Hefty algebras: Modular elaboration of higherorder algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):62:1–62:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571255.

Parreaux:2018:UAS

[PVSK18] Lionel Parreaux, Antoine Voizard, Amir Shaikhha, and Christoph E. Koch. Unifying analytic and statically-typed quasiquotes. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):13:1–13:??, January 2018. CODEN ???? ISSN 2475-

Parkinson:2017:PSN

[PVV^{+}17] Matthew Parkinson, Dimitrios Vytiniotis, Kapil Vaswani, Manuel Costa, Pantazis Deligiannis, Dylan McDermott, Aaron Blankstein, and Jonathan Balkind. Project snowflake: non-blocking safe manual memory management in .NET. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):95:1–95:??, October 2017. CODEN ???? ISSN 2475-1421.

Padon:2022:IDP

[PWK^{+}22] Oded Padon, James R. Wilcox, Jason R. Koenig, Kenneth L. McMillan, and Alex Aiken. Induction duality: primal-dual search for invariants. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):50:1–50:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498712.

Pailoor:2020:APS

[PWSD20] Shankara Pailoor, Xinyu Wang, Hovav Shacham, and Isil Dillig. Automated policy synthesis for system call sandboxing. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):135:1– 135:26, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428203.

Park:2021:GTA

[PWZS21] Jiwon Park, Dominik Winterer, Chengyu Zhang, and Zhendong Su. Generative type-aware mutation for testing SMT solvers. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):152:1–152:19, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485529.

Popescu:2021:SSA

[PXA^{+}21] Natalie Popescu, Ziyang Xu, Sotiris Apostolakis, David I. August, and Amit Levy. Safer at any speed: automatic context-aware safety enhancement for Rust. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):103:1–103:23, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485480.

Protzenko:2017:VLL

[PZR^{+}17] Jonathan Protzenko, Jean-Karim Zinzindohoué, Aseem Rastogi, Tahina Ramananandro, Peng Wang, Santiago Zanella-Béguelin, Antoine Delignat-Lavaud, Catalin Hritcu, Karthikeyan Bhargavan, Cédric Fournet, and Nikhil Swamy. Verified lowlevel programming embedded in F*. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 17:1–17:??, September 2017. CODEN ???? ISSN 2475-1421.

Qu:2019:RCA

[QGG19] Weihao Qu, Marco Gaboardi, and Deepak Garg. Relational cost analysis for functionalimperative programs. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):92:1–92:29, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341696.

Qian:2021:CSS

[QKB21] Zesen Qian, G. A. Kavvos, and Lars Birkedal. Client-server sessions in linear logic. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):62:1–62:31, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3473567.

Quiring:2022:ABE

[QRS22] Benjamin Quiring, John Reppy, and Olin Shivers. Analyzing binding extent in 3CPS. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 114:1–114:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547645.

Qiu:2017:NSP

[QSL17] Xiaokang Qiu and Armando Solar-Lezama. Natural synthesis of provably-correct datastructure manipulations. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):65:1–65:??, October 2017. CODEN ???? ISSN 2475-1421.

Rice:2017:DAS

[RAJ^{+}17] Andrew Rice, Edward Aftandilian, Ciera Jaspan, Emily Johnston, Michael Pradel, and Yulissa Arroyo-Paredes. Detecting argument selection defects. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):104:1–104:??, October 2017. CODEN ???? ISSN 2475-1421.

Ramsey:2022:BRR

[Ram22]  Norman Ramsey.  Beyond

relooper: recursive translation of unstructured control flow to structured control flow (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 90:1–90:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547621.

Richards:2017:VAK

[RAT17] Gregor Richards, Ellen Arteca, and Alexi Turcotte. The VM already knew that: leveraging compile-time knowledge to optimize gradual typing. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):55:1–55:??, October 2017. CODEN ???? ISSN 2475-

Raad:2022:CIS

[RBDO22] Azalea Raad, Josh Berdine, Derek Dreyer, and Peter W. O’Hearn. Concurrent incorrectness separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):34:1–34:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498695.

Radicek:2018:MRR

[RBG^{+}18] Ivan Radiček, Gilles Barthe, Marco Gaboardi, Deepak Garg, and Florian Zuleger. Monadic refinements for relational cost analysis. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 36:1–36:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Rocha:2021:PTS

[RC21] Pedro Rocha and Luís Caires. Propositions-as-types and shared state. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 79:1–79:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473584.

Reid:2017:WGG

[Rei17] Alastair Reid. Who guards the guards? Formal validation of the ARM v8-m architecture specification. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):88:1–88:??, October 2017. CODEN ???? ISSN 2475-1421.

Ritschel:2022:CGD

[RFH^{+}22] Nico Ritschel, Felipe Fronchetti, Reid Holmes, Ronald Garcia, and David C. Shepherd. Can guided decomposition help end-users write larger blockbased programs? A mobile robot experiment. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):133:1–133:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563296.

Roth:2023:FAF

[RG23] Ori Roth and Yossi Gil. Fluent APIs in functional languages. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):105:1– 105:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3586057.

Rajani:2021:UTT

[RGGH21] Vineet Rajani, Marco Gaboardi, Deepak Garg, and Jan Hoffmann. A unifying type-theory for higher-order (amortized) cost analysis. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 27:1–27:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434308.

Ringer:2017:SAL

[RGSNT17] Talia Ringer, Dan Grossman, Daniel Schwartz-Narbonne, and Serdar Tasiran. A solver-aided language for test input generation. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):91:1– 91:??, October 2017. CODEN ???? ISSN 2475-1421.

Ritter:2022:AAI

[RH22] Fabian Ritter and Sebastian Hack. AnICA: analyzing inconsistencies in microarchitectural code analyzers. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):125:1–125:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563288.

Rioux:2023:BBO

[RHdSOZ23] Nick Rioux, Xuejing Huang, Bruno C. d. S. Oliveira, and Steve Zdancewic. A bowtie for a beast: Overloading, eta expansion, and extensible data types in F▷◁. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 18:1–18:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571211.

Rivera:2022:SVP

[RK22] Elijah Rivera and Shriram Krishnamurthi. Structural versus pipeline composition of higher-order functions (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 102:1–102:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547633.

Rapoport:2017:SSP

[RKHL17] Marianna Rapoport, Ifaz Kabir, Paul He, and Ondrej Lhoták. A simple soundness proof for dependent object types. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):46:1–46:??, October 2017. CODEN ???? ISSN 2475-1421.

Rama:2018:ROS

[RKS18] Girish Maskeri Rama, Raghavan Komondoor, and Himanshu Sharma. Refinement in object-sensitivity points-to analysis via slicing. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):142:1–142:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276512.

Rouvoet:2021:ITC

[RKV21] Arjen Rouvoet, Robbert Krebbers, and Eelco Visser. Intrinsically typed compilation with nameless labels. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):22:1–22:28, January

2021. URL https://dl.acm.
org/doi/10.1145/3434303.

Rapoport:2019:PDF

[RL19] Marianna Rapoport and Ondrej Lhoták. A path to DOT: formalizing fully pathdependent types. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):145:1–145:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360571.

Reid:2020:HCS

[RLS20] John Reid, Bill Long, and Jon Steidel. History of coarrays and SPMD parallelism in Fortran. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):72:1–72:30, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386322.

Raad:2020:POG

[RLV20] Azalea Raad, Ori Lahav, and Viktor Vafeiadis. Persistent Owicki—Gries reasoning: a program logic for reasoning about persistent programs on Intel-x86. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):151:1–151:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428219.

Rosemann:2021:AIS

[RMH21] Julian Rosemann, Simon Moll, and Sebastian Hack. An abstract interpretation for SPMD divergence on reducible control flow graphs. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):31:1–31:31, January

2021. URL https://dl.acm.
org/doi/10.1145/3434312.

Raad:2022:EIX

[RMV22] Azalea Raad, Luc Maranget, and Viktor Vafeiadis. Extending Intel-x86 consistency and persistency: formalising the semantics of Intel-x86 memory types and non-temporal stores. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 22:1–22:31, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498683.

Raghothaman:2020:PGS

[RMZ^{+}20] Mukund Raghothaman, Jonathan Mendelson, David Zhao, Mayur Naik, and Bernhard Scholz. Provenance-guided synthesis of Datalog programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):62:1–62:27, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371130.

Rahmani:2019:CDT

[RNDJ19] Kia Rahmani, Kartik Nagar, Benjamin Delaware, and Suresh Jagannathan. CLOTHO: directed test generation for weakly consistent database systems. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):117:1– 117:28, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360543.

Roth:2021:SSM

[Rot21] Ori Roth. Study of the subtyping machine of nominal subtyping with variance. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):137:1–137:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485514.

Radul:2023:YOL

[RPF^{+}23] Alexey Radul, Adam Paszke, Roy Frostig, Matthew J. Johnson, and Dougal Maclaurin. You only linearize once: Tangents transpose to gradients. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 43:1–43:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571236.

Rahmani:2021:MMP

[RRG^{+}21] Kia Rahmani, Mohammad Raza, Sumit Gulwani, Vu Le, Daniel Morris, Arjun Radhakrishna, Gustavo Soares, and Ashish Tiwari. Multimodal program inference: a marriage of pre-trained language models and componentbased synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):158:1–158:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485535.

Rigger:2020:FBD

[RS20] Manuel Rigger and Zhendong Su. Finding bugs in database systems via query partitioning. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):211:1–211:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428279.

Ricciotti:2017:IFP

[RSPC17] Wilmer Ricciotti, Jan Stolarek, Roly Perera, and James Cheney. Imperative functional programs that explain their work. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 14:1–14:??, September 2017. CODEN ???? ISSN 2475-1421.

Radanne:2020:KBF

[RST20] Gabriel Radanne, Hannes Saffrich, and Peter Thiemann. Kindly bent to free us. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):103:1–103:29, August

2020. URL https://dl.acm.
org/doi/10.1145/3408985.

Reynaud:2021:PMS

[RSY21] Alban Reynaud, Gabriel Scherer, and Jeremy Yallop. A practical mode system for recursive definitions. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):

45:1–45:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434326.

Raad:2018:PSW

[RV18] Azalea Raad and Viktor Vafeiadis. Persistence semantics for weak memory: integrating epoch persistency with the TSO memory model. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):137:1–137:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276507.

Rouvoet:2020:KWA

[RvAP^{+}20] Arjen Rouvoet, Hendrik van Antwerpen, Casper Bach Poulsen, Robbert Krebbers, and Eelco Visser. Knowing when to ask: sound scheduling of name resolution in type checkers derived from declarative specifications. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):180:1–180:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428248.

Raad:2020:PSI

[RWNV20] Azalea Raad, John Wickerson, Gil Neiger, and Viktor Vafeiadis. Persistency semantics of the Intel-x86 architecture. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 11:1–11:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371079.

Raad:2019:WPS

[RWV19] Azalea Raad, John Wickerson, and Viktor Vafeiadis. Weak persistency semantics from the ground up: formalising the persistency semantics of ARMv8 and transactional models. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):135:1–135:27, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360561.

Rioux:2020:CF

[RZ20] Nick Rioux and Steve Zdancewic.

Computation focusing. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):95:1–95:27, August

2020. URL https://dl.acm.
org/doi/10.1145/3408977.

Smith:2019:SDP

[SA19] Calvin Smith and Aws Albarghouthi. Synthesizing differentially private programs. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):94:1–94:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341698.

Spath:2017:IEP

[SAB17] Johannes Späth, Karim Ali, and Eric Bodden. IDEal: efficient and precise alias-aware dataflow analysis. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):99:1–99:??, October 2017. CODEN ???? ISSN 2475-

Shapira:2023:DLR

[SADC23] Yuval Shapira, Eran Avneri, and Dana Drachsler-Cohen. Deep learning robustness verification for few-pixel attacks. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):90:1– 90:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586042.

St-Amour:2017:HRS

[SAFF^{+}17] Vincent St-Amour, Daniel Feltey, Spencer P. Florence, Shu-Hung You, and Robert Bruce Findler. Herbarium Racketensis: a stroll through the woods (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP):1:1–1:??, September 2017. CODEN ???? ISSN 2475-1421.

Sangiorgi:2022:ECT

[San22] Davide Sangiorgi. From enhanced coinduction towards enhanced induction. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):18:1–18:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498679.

Swierstra:2019:PTS

[SB19] Wouter Swierstra and Tim Baanen. A predicate transformer semantics for effects (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):103:1–103:26, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341707.

Szabo:2018:ILB

[SBEV18] Tamás Szabó, Gábor Bergmann, Sebastian Erdweg, and Markus Voelter. Incrementalizing lattice-based program analyses in Datalog. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):139:1–139:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276509.

Sozeau:2020:CCC

[SBF^{+}20] Matthieu Sozeau, Simon Boulier, Yannick Forster, Nicolas Tabareau, and Théo Winterhalter. Coq Coq correct! Verification of type checking and erasure for Coq, in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):8:1–8:28, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371076.

Six:2020:CEI

[SBM20] Cyril Six, Sylvain Boulmé, and David Monniaux. Certified and efficient instruction scheduling: application to interlocked VLIW processors. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):129:1–129:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428197.

Schuster:2020:CEH

[SBO20] Philipp Schuster, Jonathan Immanuel Brachthäuser, and Klaus Ostermann. Compiling effect handlers in capabilitypassing style. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 93:1–93:28, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408975.

Stampoulis:2018:PFL

[SC18] Antonis Stampoulis and Adam Chlipala. Prototyping a functional language using higherorder logic programming: a functional pearl on learning the ways of λ Prolog Makam. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):93:1–93:30, July 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3236788.

Shahin:2020:AEV

[SC20] Ramy Shahin and Marsha Chechik. Automatic and efficient variability-aware lifting of functional programs. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):157:1–157:27, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428225.

Schwerter:2021:AGT

[SCJG21] Felipe Bañados Schwerter, Alison M. Clark, Khurram A.

Jafery, and Ronald Garcia. Abstracting gradual typing moving forward: precise and spaceefficient. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 61:1–61:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434342.

Song:2020:CCC

[SCK^{+}20] Youngju Song, Minki Cho, Dongjoo Kim, Yonghyun Kim, Jeehoon Kang, and Chung-Kil Hur. CompCertM: CompCert with C-assembly linking and lightweight modular verification. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 23:1–23:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371091.

Song:2023:CCR

[SCL^{+}23] Youngju Song, Minki Cho, Dongjae Lee, Chung-Kil Hur, Michael Sammler, and Derek Dreyer. Conditional contextual refinement. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 39:1–39:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571232.

Sotiropoulos:2020:MDF

[SCMS20] Thodoris Sotiropoulos, Stefanos Chaliasos, Dimitris Mitropoulos, and Diomidis Spinellis. A model for detecting faults in build specifications. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):144:1–144:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428212.

Sui:2020:FVF

[SCZW20] Yulei Sui, Xiao Cheng, Guanqin Zhang, and Haoyu Wang. Flow2Vec: value-flow-based precise code embedding. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):233:1–233:27, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428301.

Skorstengaard:2019:SEW

[SDB19] Lau Skorstengaard, Dominique Devriese, and Lars Birkedal. StkTokens: enforcing wellbracketed control flow and stack encapsulation using linear capabilities. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 19:1–19:28, January 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3290332.

Sun:2022:CED

[SDdSO22] Yaozhu Sun, Utkarsh Dhandhania, and Bruno C. d. S. Oliveira. Compositional embeddings of domain-specific languages. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):131:1–131:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563294.

Sousa:2018:VTW

[SDL18] Marcelo Sousa, Isil Dillig, and Shuvendu K. Lahiri. Verified three-way program merge. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):165:1– 165:29, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276535.

Siraichi:2019:QAC

[SdSCQ19] Marcos Yukio Siraichi, Vinícius Fernandes dos Santos, Caroline Collange, and Fernando Magno Quintão Pereira. Qubit allocation as a combination of subgraph isomorphism and token swapping. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):120:1–120:29, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360546.

Sivaramakrishnan:2020:RPO

[SDW^{+}20] KC Sivaramakrishnan, Stephen Dolan, Leo White, Sadiq Jaffer, Tom Kelly, Anmol Sahoo, Sudha Parimala, Atul Dhiman, and Anil Madhavapeddy. Retrofitting parallelism onto OCaml. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 113:1–113:30, August 2020.

URL https://dl.acm.org/
doi/10.1145/3408995.

Sakkas:2022:SNP

[SEG^{+}22] Georgios Sakkas, Madeline Endres, Philip J. Guo, Westley Weimer, and Ranjit Jhala. Seq2Parse: neurosymbolic parse error repair. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):167:1–167:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563330.

Sang:2020:SSN

[SEP^{+}20] Bo Sang, Patrick Eugster, Gustavo Petri, Srivatsan Ravi, and Pierre-Louis Roman. Scalable and serializable networked multi-actor programming. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):198:1– 198:30, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428266.

Serrano:2021:JAC

[Ser21] Manuel Serrano. Of JavaScript AOT compilation performance. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):70:1–70:30, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3473575.

Smolka:2020:GKA

[SFH^{+}20] Steffen Smolka, Nate Foster, Justin Hsu, Tobias Kappé, Dexter Kozen, and Alexandra Silva. Guarded Kleene algebra with tests: verification of uninterpreted programs in nearly linear time. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 61:1–61:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371129.

Sun:2022:CPP

[SFH22] Zhihang Sun, Hongyu Fan, and Fei He. Consistencypreserving propagation for SMT solving of concurrent program verification. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):158:1–158:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563321.

Saad:2020:OAS

[SFRM20] Feras A. Saad, Cameron E.

Freer, Martin C. Rinard, and Vikash K. Mansinghka. Optimal approximate sampling from discrete probability distributions. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 36:1–36:31, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371104.

Shaikhha:2019:EDP

[SFVJ19] Amir Shaikhha, Andrew Fitzgibbon, Dimitrios Vytiniotis, and Simon Peyton Jones. Efficient differentiable programming in a functional arrayprocessing language. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):97:1–97:30, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341701.

Stucki:2021:THO

[SG21] Sandro Stucki and Paolo G. Giarrusso. A theory of higherorder subtyping with type intervals. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 69:1–69:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473574.

Swasey:2017:RCV

[SGD17] David Swasey, Deepak Garg, and Derek Dreyer. Robust and compositional verification of object capability patterns. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):89:1– 89:??, October 2017. CODEN ???? ISSN 2475-1421.

Sammler:2020:HLB

[SGDL20] Michael Sammler, Deepak Garg, Derek Dreyer, and Tadeusz Litak. The highlevel benefits of low-level sandboxing. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 32:1–32:32, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371100.

Smaragdakis:2021:SVF

[SGL^{+}21] Yannis Smaragdakis, Neville Grech, Sifis Lagouvardos, Konstantinos Triantafyllou, and Ilias Tsatiris. Symbolic valueflow static analysis: deep, precise, complete modeling of Ethereum smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):163:1–163:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485540.

Spies:2022:LCR

[SGT^{+}22] Simon Spies, Lennard Gäher, Joseph Tassarotti, Ralf Jung, Robbert Krebbers, Lars Birkedal, and Derek Dreyer. Later credits: resourceful reasoning for the later modality. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):100:1–100:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547631.

Selsam:2020:SPB

[SHdM20] Daniel Selsam, Simon Hudon, and Leonardo de Moura. Sealing pointer-based optimizations behind pure functions. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):115:1–115:20, August

2020. URL https://dl.acm.
org/doi/10.1145/3408997.

Serrano:2020:QLI

[SHJV20] Alejandro Serrano, Jurriaan Hage, Simon Peyton Jones, and Dimitrios Vytiniotis. A quick look at impredicativity. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 89:1–89:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408971.

Solomon:2020:HL

[SHK^{+}20] Cynthia Solomon, Brian Harvey, Ken Kahn, Henry Lieberman, Mark L. Miller, Artemis Papert, and Brian Silverman. History of Logo. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):79:1–79:66, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386329.

Shaikhha:2022:FCP

[SHSO22] Amir Shaikhha, Mathieu Huot, Jaclyn Smith, and Dan Olteanu. Functional collection programming with semi-ring dictionaries. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):89:1– 89:33, April 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3527333.

Shi:2019:RAS

[SHTZ^{+}19] August Shi, Milica Hadzi-Tanovic, Lingming Zhang, Darko Marinov, and Owolabi Legunsen. Reflection-aware static regression test selection. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):187:1– 187:29, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360613.

Senanayake:2020:SIS

[SHW^{+}20] Ryan Senanayake, Changwan Hong, Ziheng Wang, Amalee Wilson, Stephen Chou, Shoaib Kamil, Saman Amarasinghe, and Fredrik Kjolstad. A sparse iteration space transformation framework for sparse tensor algebra. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):158:1–158:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428226.

Surbatovich:2019:DIB

[SJL19] Milijana Surbatovich, Limin Jia, and Brandon Lucia. I/O dependent idempotence bugs in intermittent systems. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):183:1–183:31, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360609.

Stump:2020:SFP

[SJSM20] Aaron Stump, Christopher Jenkins, Stephan Spahn, and Colin McDonald. Strong functional pearl: Harper’s regular-expression matcher in Cedille. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):122:1– 122:25, August 2020. URL

https://dl.acm.org/doi/10.
1145/3409004.

Spiwack:2022:LQT

[SKB^{+}22] Arnaud Spiwack, Csongor Kiss, Jean-Philippe Bernardy, Nicolas Wu, and Richard A. Eisenberg. Linearly qualified types: generic inference for capabilities and uniqueness. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 95:1–95:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547626.

Spies:2021:TSI

[SKD21] Simon Spies, Neel Krishnaswami, and Derek Dreyer. Transfinite step-indexing for termination. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 13:1–13:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434294.

Sprenger:2020:ISL

[SKE^{+}20] Christoph Sprenger, Tobias Klenze, Marco Eilers, Felix A. Wolf, Peter Müller, Martin Clochard, and David Basin. Igloo: soundly linking compositional refinement and separation logic for distributed system verification. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):152:1–152:31, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428220.

Scibior:2018:FPM

[ŚKG18] Adam Ścibior, Ohad Kammar, and Zoubin Ghahramani. Functional programming for modular Bayesian inference. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):83:1–83:29, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236778.

Samak:2020:SRC

[SKR20] Malavika Samak, Deokhwan Kim, and Martin C. Rinard. Synthesizing replacement classes. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 52:1–52:33, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371120.

Salvaneschi:2019:LIP

[SKS^{+}19] Guido Salvaneschi, Mirko Köhler, Daniel Sokolowski, Philipp Haller, Sebastian Erdweg, and Mira Mezini. Languageintegrated privacy-aware distributed queries. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):167:1–167:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360593.

Scibior:2018:DVH

[ŚKV^{+}18] Adam Ścibior, Ohad Kammar, Matthijs Vákár, Sam Staton, Hongseok Yang, Yufei Cai, Klaus Ostermann, Sean K. Moss, Chris Heunen, and Zoubin Ghahramani. Denotational validation of higher-order Bayesian inference. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):60:1–60:??, January 2018. CODEN ???? ISSN 2475-1421.

Singh:2023:OAL

[SL23] Abhishek Kr Singh and Ori Lahav. An operational approach to library abstraction under relaxed memory concurrency. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 53:1–53:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571246.

Scott:2017:MCD

[SLDN17] Ryan G. Scott, Omar S. Navarro Leija, Joseph Devietti, and Ryan R. Newton. Monadic composition for deterministic, parallel batch processing. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):73:1–73:??, October 2017. CODEN ???? ISSN 2475-1421.

Susag:2022:SER

[SLHR22] Zachary Susag, Sumit Lahiri, Justin Hsu, and Subhajit Roy. Symbolic execution for randomized programs. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):181:1–181:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563344.

Surbatovich:2020:TFF

[SLJ20] Milijana Surbatovich, Brandon Lucia, and Limin Jia. Towards a formal foundation of intermittent computing. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):163:1–163:31, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428231.

Song:2019:ASD

[SLO19] Dowon Song, Myungho Lee, and Hakjoo Oh. Automatic and scalable detection of logical errors in functional programming assignments. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):188:1–188:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360614.

Sozeau:2019:ERH

[SM19] Matthieu Sozeau and Cyprien Mangin. Equations reloaded: high-level dependently-typed functional programming and proving in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):86:1–86:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341690.

Stievenart:2020:FCB

[SM20] Quentin Stiévenart and Magnus Madsen. Fuzzing channelbased concurrency runtimes using types and effects. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):186:1–186:27, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428254.

Sherman:2019:SRS

[SMC19] Benjamin Sherman, Jesse Michel, and Michael Carbin. Sound and robust solid modeling via exact real arithmetic and continuity. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):99:1–99:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341703.

Sherman:2021:CSD

[SMC21] Benjamin Sherman, Jesse Michel, and Michael Carbin. λ_{s}: computable semantics for differentiable programming with higher-order functions and datatypes. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 3:1–3:31, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434284.

Sampson:2017:SSH

[SMM17] Adrian Sampson, Kathryn S. McKinley, and Todd Mytkowicz. Static stages for heterogeneous programming. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):71:1–71:??, October 2017. CODEN ???? ISSN 2475-1421.

Santos:2018:JJV

[SMN^{+}18] José Fragoso Santos, Petar Maksimović, Daiva Naudžiūnienė, Thomas Wood, and Philippa Gardner. JaVerT: JavaScript verification toolchain. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):50:1–50:??, January 2018. CODEN ???? ISSN 2475-1421.

Spall:2020:BSP

[SMTH20] Sarah Spall, Neil Mitchell, and Sam Tobin-Hochstadt. Build scripts with perfect dependencies. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):169:1–169:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428237.

Shajii:2019:SHP

[SNB^{+}19] Ariya Shajii, Ibrahim Numanagić, Riyadh Baghdadi, Bonnie Berger, and Saman Amarasinghe. Seq: a highperformance language for bioinformatics. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 125:1–125:29, October 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3360551.

Stein:2019:SAD

[SNCM19] Benno Stein, Benjamin Barslev Nielsen, Bor-Yuh Evan Chang, and Anders Møller. Static analysis with demand-driven value refinement. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):140:1–140:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360566.

Sergey:2019:SSC

[SNJ^{+}19] Ilya Sergey, Vaivaswatha Nagaraj, Jacob Johannsen, Amrit Kumar, Anton Trunov, and Ken Chan Guan Hao. Safer smart contract programming with Scilla. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):185:1–185:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360611.

Sela:2022:CS

[SP22] Gal Sela and Erez Petrank. Concurrent size. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):137:1–137:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563300.

Spivey:2017:FCP

[Spi17] Michael Spivey. Faster coroutine pipelines. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 5:1–5:??, September 2017. CO-DEN ???? ISSN 2475-1421.

Selakovic:2018:TGH

[SPKT18] Marija Selakovic, Michael Pradel, Rezwana Karim, and Frank Tip. Test generation for higher-order functions in dynamic languages. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):161:1–161:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276531.

Singh:2018:PCD

[SPV18] Gagandeep Singh, Markus Püschel, and Martin Vechev. A practical construction for decomposing numerical abstract domains. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 55:1–55:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Swamy:2020:SEC

[SRF^{+}20] Nikhil Swamy, Aseem Rastogi, Aymeric Fromherz, Denis Merigoux, Danel Ahman, and Guido Martínez. Steel-Core: an extensible concurrent separation logic for effectful dependently typed programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):121:1– 121:30, August 2020. URL

https://dl.acm.org/doi/10.
1145/3409003.

Seidel:2017:LBL

[SSC^{+}17] Eric L. Seidel, Huma Sibghat, Kamalika Chaudhuri, Westley Weimer, and Ranjit Jhala. Learning to blame: localizing novice type errors with data-driven diagnosis. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):60:1–60:??, October 2017. CODEN ???? ISSN 2475-1421.

Sjoberg:2019:DLC

[SScWS19] Vilhelm Sjöberg, Yuyang Sang, Shu chun Weng, and Zhong Shao. DeepSEA: a language for certified system software. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):136:1– 136:27, October 2019. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3360562.

Sharif:2019:APC

[SSH^{+}19] Hashim Sharif, Prakalp Srivastava, Muhammad Huzaifa, Maria Kotsifakou, Keyur Joshi, Yasmin Sarita, Nathan Zhao, Vikram S. Adve, Sasa Misailovic, and Sarita Adve. ApproxHPVM: a portable compiler IR for accuracy-aware optimizations. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):186:1–186:30, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360612.

Sakka:2017:TFA

[SSK17] Laith Sakka, Kirshanthan Sundararajah, and Milind Kulkarni. TreeFuser: a framework for analyzing and fusing general recursive tree traversals. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):76:1–76:??, October 2017. CODEN ???? ISSN 2475-1421.

Sundararajah:2022:UUL

[SSK22]  Kirshanthan  Sundararajah,

Charitha Saumya, and Milind Kulkarni. UniRec: a unimodularlike framework for nested recursions and loops. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):170:1–170:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563333.

Sorensen:2021:STG

[SSR^{+}21] Tyler Sorensen, Lucas F. Salvador, Harmit Raval, Hugues Evrard, John Wickerson, Margaret Martonosi, and Alastair F. Donaldson. Specifying and testing GPU workgroup progress models. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):131:1–131:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485508.

Sammler:2023:DDA

[SSS^{+}23] Michael Sammler, Simon Spies, Youngju Song, Emanuele DOsualdo, Robbert Krebbers, Deepak Garg, and Derek Dreyer. DimSum: a decentralized approach to multilanguage semantics and verification. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 27:1–27:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571220.

Sivaraman:2022:DDL

[SSSC^{+}22] Aishwarya Sivaraman, Alex Sanchez-Stern, Bretton Chen, Sorin Lerner, and Todd Millstein. Data-driven lemma synthesis for interactive proofs. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):143:1–143:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563306.

Sabok:2021:PPS

[SSSW21] Marcin Sabok, Sam Staton, Dario Stein, and Michael Wolman. Probabilistic programming semantics for name generation. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 11:1–11:29, January 2021. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3434292.

Sekiyama:2021:CTA

[ST21] Taro Sekiyama and Takeshi Tsukada. CPS transformation with affine types for call-by-value implicit polymorphism. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 95:1–95:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473600.

Sekiyama:2020:SRP

[STI20] Taro Sekiyama, Takeshi Tsukada, and Atsushi Igarashi. Signature restriction for polymorphic algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):117:1–117:30, August

2020. URL https://dl.acm.
org/doi/10.1145/3408999.

Stroustrup:2020:TCC

[Str20] Bjarne Stroustrup. Thriving in a crowded and changing world: C++ 2006–2020. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):70:1–70:168, June

2020. URL https://dl.acm.
org/doi/abs/10.1145/3386320.

Sekiyama:2023:TVA

[SU23] Taro Sekiyama and Hiroshi Unno. Temporal verification with answer-effect modification: Dependent temporal typeand-effect system with delimited continuations. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):71:1–71:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571264.

Steele:2021:PLB

[SV21] Guy L. Steele Jr. and Sebastiano Vigna. LXM: better splittable pseudorandom number generators (and almost as fast). Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):148:1–148:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485525.

Smeding:2023:EDN

[SV23] Tom J. Smeding and Matthijs I. L. Vákár. Efficient dualnumbers reverse AD via wellknown program transformations. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 54:1–54:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571247.

Srinivasan:2017:MAM

[SVR17] Venkatesh Srinivasan, Ara Vartanian, and Thomas Reps. Model-assisted machine-code synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA): 61:1–61:??, October 2017. CO-DEN ???? ISSN 2475-1421.

Sergey:2018:PPD

[SWT18] Ilya Sergey, James R. Wilcox, and Zachary Tatlock. Programming and proving with distributed protocols. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):28:1–28:??, January 2018. CODEN ???? ISSN 2475-

Shi:2022:IED

[SWYZ22] Qingkai Shi, Yongchao Wang, Peisen Yao, and Charles Zhang. Indexing the extended Dyck-CFL reachability for context-sensitive program analysis. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):176:1–176:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563339.

Scalas:2019:LMM

[SY19] Alceste Scalas and Nobuko Yoshida. Less is more: multiparty session types revisited. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 30:1–30:29, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290343.

Syme:2020:EHF

[Sym20] Don Syme. The early history of F#. Proceedings of the ACM on Programming Languages (PACMPL), 4 (HOPL):75:1–75:58, June 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3386325.

Su:2021:FAF

[SYW^{+}21] Ting Su, Yichen Yan, Jue Wang, Jingling Sun, Yiheng Xiong, Geguang Pu, Ke Wang, and Zhendong Su. Fully automated functional fuzzing of Android apps for detecting noncrashing logic bugs. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):156:1–156:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485533.

Silver:2021:DMF

[SZ21] Lucas Silver and Steve Zdancewic. Dijkstra monads forever: terminationsensitive specifications for interaction trees. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 26:1–26:28, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434307.

Santolucito:2017:SCF

[SZD^{+}17] Mark Santolucito, Ennan Zhai, Rahul Dhodapkar, Aaron Shim, and Ruzica Piskac. Synthesizing configuration file specifications with association rule learning. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):64:1– 64:??, October 2017. CODEN ???? ISSN 2475-1421.

Shen:2021:COD

[SZD^{+}21] Xipeng Shen, Guoqiang Zhang, Irene Dea, Samantha Andow, Emilio Arroyo-Fang, Neal Gafter, Johann George, Melissa Grueter, Erik Meijer, Olin Grigsby Shivers, Steffi Stumpos, Alanna Tempest, Christy Warden, and Shannon Yang. Coarsening optimization for differentiable programming. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):130:1–130:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485507.

Shen:2019:IRA

[SZZ^{+}19] Bo Shen, Wei Zhang, Haiyan Zhao, Guangtai Liang, Zhi Jin, and Qianxiang Wang. IntelliMerge: a refactoringaware software merging technique. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):170:1– 170:28, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360596.

Timany:2019:MRV

[TB19] Amin Timany and Lars Birkedal.

Mechanized relational verification of concurrent programs with continuations. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):105:1–105:28, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341709.

Tosch:2019:PAT

[TBB^{+}19] Emma Tosch, Eytan Bakshy, Emery D. Berger, David D. Jensen, and J. Eliot B. Moss. PlanAlyzer: assessing threats to the validity of online experiments. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):182:1–182:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360608.

Trinh:2020:ITD

[TCJ20] Minh-Thai Trinh, Duc-Hiep Chu, and Joxan Jaffar. Intertheory dependency analysis for SMT string solvers. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):192:1–192:27, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428260.

Tejiscak:2020:DTC

[Tej20] Matús Tejiscák. A dependently typed calculus with pattern matching and erasure inference. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 91:1–91:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408973.

Turcotte:2020:DTR

[TGKV20] Alexi Turcotte, Aviral Goel, Filip Krikava, and Jan Vitek. Designing types for R, empirically. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):181:1– 181:25, November 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3428249.

Ter-Gabrielyan:2019:MVH

[TGSM19] Arshavir Ter-Gabrielyan, Alexander J. Summers, and Peter Müller. Modular verification of heap reachability properties in separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):121:1–121:28, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360547.

Turonova:2020:RMC

[THL^{+}20] Lenka Turoňová, Lukáš Holík, Ondřej Lengál, Olli Saarikivi, Margus Veanes, and Tomáš Vojnar. Regex matching with counting-set automata. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):218:1–218:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428286.

Titzer:2022:FPI

[Tit22] Ben L. Titzer. A fast in-place interpreter for WebAssembly. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):148:1– 148:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563311.

Ta:2018:ALS

[TLKC18] Quang-Trung Ta, Ton Chanh Le, Siau-Cheng Khoo, and Wei-Ngan Chin. Automated lemma synthesis in symbolic-heap separation logic. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 9:1–9:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Tan:2021:MPA

[TLM^{+}21] Tian Tan, Yue Li, Xiaoxing Ma, Chang Xu, and Yannis Smaragdakis. Making pointer analysis more precise by unleashing the power of selective context sensitivity. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):147:1–147:27, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485524.

Toro:2019:GPR

[TLT19] Matías Toro, Elizabeth Labrada, and Éric Tanter. Gradual parametricity, revisited. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):17:1–17:30, January

2019. URL https://dl.acm.
org/doi/abs/10.1145/3290330.

Tan:2022:SRT

[TML^{+}22] Bryan Tan, Benjamin Mariano, Shuvendu K. Lahiri, Isil Dillig, and Yu Feng. SolType: refinement types for arithmetic overflow in Solidity. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 4:1–4:29, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498665.

Thomson:2022:FIA

[TRWS22] Patrick Thomson, Rob Rix, Nicolas Wu, and Tom Schrijvers. Fusing industry and academia at GitHub (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):

108:1–108:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547639.

Timany:2018:LRM

[TSKJB18] Amin Timany, Léo Stefanesco, Morten Krogh-Jespersen, and Lars Birkedal. A logical relation for monadic encapsulation of state: proving contextual equivalences in the presence of runST. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 64:1–64:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Tabareau:2018:EFU

[TTS18] Nicolas Tabareau, Éric Tanter, and Matthieu Sozeau. Equivalences for free: univalent parametricity for effective transport. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):92:1–92:29, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236787.

Tsukada:2022:SMC

[TU22] Takeshi Tsukada and Hiroshi Unno. Software model-checking as cyclic-proof search. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):63:1–63:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498725.

Thiemann:2020:LDS

[TV20] Peter Thiemann and Vasco T. Vasconcelos. Label-dependent session types. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 67:1–67:29, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371135.

Ugawa:2017:MCC

[UAM17] Tomoharu Ugawa, Tatsuya Abe, and Toshiyuki Maeda. Model checking copy phases of concurrent copying garbage collection with various memory models. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):53:1– 53:??, October 2017. CODEN ???? ISSN 2475-1421.

Urban:2020:PPF

[UCWZ20] Caterina Urban, Maria Christakis, Valentin Wüstholz, and Fuyuan Zhang. Perfectly parallel fairness certification of neural networks. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):185:1–185:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428253.

Ullrich:2022:DUE

[UdM22] Sebastian Ullrich and Leonardo de Moura. ‘do’ unchained: embracing local imperativity in a purely functional language (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):109:1–109:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547640.

Ugare:2022:PTF

[USM22] Shubham Ugare, Gagandeep Singh, and Sasa Misailovic. Proof transfer for fast certification of multiple approximate neural networks. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):75:1–75:29, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527319.

Unno:2018:RCR

[UST18] Hiroshi Unno, Yuki Satake, and Tachio Terauchi. Relatively complete refinement type system for verification of higherorder non-deterministic programs. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 12:1–12:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Unno:2023:MPD

[UTGK23] Hiroshi Unno, Tachio Terauchi, Yu Gu, and Eric Koskinen. Modular primal-dual fixpoint logic solving for temporal verification. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 72:1–72:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571265.

vanAntwerpen:2018:ST

[vAPRV18] Hendrik van Antwerpen, Casper Bach Poulsen, Arjen Rouvoet, and Eelco Visser. Scopes as types. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):114:1– 114:30, October 2018. URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3276484.

Vu:2021:ROS

[VCD^{+}21] Son Tuan Vu, Albert Cohen, Arnaud De Grandmaison, Christophe Guillon, and Karine Heydemann. Reconciling optimization with secure compilation. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):142:1–142:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485519.

vanderRest:2022:ITD

[vdRPR^{+}22] Cas van der Rest, Casper Bach Poulsen, Arjen Rouvoet, Eelco Visser, and Peter Mosses. Intrinsically-typed definitional interpreters à la carte. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):192:1–192:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563355.

vanderRest:2022:CUA

[vdRS22] Cas van der Rest and Wouter Swierstra. A completely unique account of enumeration. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):105:1– 105:??, August 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3547636.

Vassena:2021:AES

[VDvG^{+}21] Marco Vassena, Craig Disselkoen, Klaus von Gleissenthall, Sunjay Cauligi, Rami Gökhan Kıcı, Ranjit Jhala, Dean Tullsen, and Deian Stefan. Automatically eliminating speculative leaks from cryptographic code with Blade. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 49:1–49:30, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434330.

Viering:2021:MST

[VHEZ21] Malte Viering, Raymond Hu, Patrick Eugster, and Lukasz Ziarek. A multiparty session typing discipline for fault-tolerant event-driven distributed programming. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):124:1–124:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485501.

VanRoy:2020:HOM

[VHSS20] Peter Van Roy, Seif Haridi, Christian Schulte, and Gert Smolka. A history of the Oz multiparadigm language. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):83:1–83:56, June 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386333.

Verbruggen:2021:SPE

[VLG21] Gust Verbruggen, Vu Le, and Sumit Gulwani. Semantic programming by example with pre-trained models. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):100:1–100:25, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485477.

Voichick:2023:QUL

[VLRH23] Finn Voichick, Liyi Li, Robert Rand, and Michael Hicks. Qunity: a unified language for quantum and classical computing. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 32:1–32:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571225.

Vezzosi:2019:CAD

[VMA19] Andrea Vezzosi, Anders Mörtberg, and Andreas Abel. Cubical Agda: a dependently typed programming language with univalence and higher inductive types. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):87:1–87:29, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341691.

Vale:2022:LOB

[VMS^{+}22] Arthur Oliveira Vale, Paul-André Melliès, Zhong Shao, Jérémie Koenig, and Léo Stefanesco. Layered and objectbased game semantics. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):42:1–42:32, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498703.

Vedurada:2018:IRO

[VN18] Jyothi Vedurada and V. Krishna Nandivada. Identifying refactoring opportunities for replacing type code with subclass and state. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):138:1–138:28, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276508.

VanStrydonck:2019:LCF

[VPD19] Thomas Van Strydonck, Frank Piessens, and Dominique Devriese. Linear capabilities for fully abstract compilation of separation-logic-verified code. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):84:1–84:29, July 2019.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3341688.

Valliappan:2022:NFS

[VRC22] Nachiappan Valliappan, Fabian Ruch, and Carlos Tomé Cortiñas. Normalization for fitch-style modal calculi. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 118:1–118:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547649.

Vukotic:2019:ATS

[VREV19] Ivana Vukotic, Vincent Rahli, and Paulo Esteves-Veríssimo. Asphalion: trustworthy shielding against Byzantine faults. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):138:1–138:32, October 2019. URL https://dl.

                                        acm.org/doi/abs/10.1145/3360564.

Vandenbroucke:2020:PFP

[VS20] Alexander Vandenbroucke and Tom Schrijvers. PλωNK: functional probabilistic NetKAT. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):39:1–39:27, January

                                        2020. URL https://dl.acm.
                                        org/doi/abs/10.1145/3371107.

Vale:2023:CTL

[VSC23] Arthur Oliveira Vale, Zhong Shao, and Yixuan Chen. A compositional theory of linearizability. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 38:1–38:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571231.

Vazou:2018:RRC

[VTC^{+}18] Niki Vazou, Anish Tondwalkar, Vikraman Choudhury, Ryan G. Scott, Ryan R. Newton, Philip Wadler, and Ranjit Jhala. Refinement reflection: complete verification with SMT. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):53:1–53:??, January 2018. CODEN ???? ISSN 2475-1421.

Vazou:2018:GLT

[VTV18] Niki Vazou, Éric Tanter, and David Van Horn. Gradual liquid type inference. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):132:1–132:25, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276502.

Vasilenko:2022:SCC

[VVB22] Elizaveta Vasilenko, Niki Vazou, and Gilles Barthe. Safe couplings: coupled refinement types. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 112:1–112:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3547643.

Westrick:2022:EDN

[WAA22] Sam Westrick, Jatin Arora, and Umut A. Acar. Entanglement detection with nearzero cost. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP): 115:1–115:??, August 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3547646.

Wadler:2017:EM

[Wad17] Philip Wadler. Editorial message. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP):1:1–1:??, September 2017. CODEN ???? ISSN 2475-1421.

Wang:2022:IUA

[WAPJ22] Zi Wang, Aws Albarghouthi, Gautam Prakriya, and Somesh Jha. Interval universal approximation for neural networks. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 14:1–14:29, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498675.

Winter:2023:RTB

[WBdG^{+}23] Levin N. Winter, Florena Buse, Daan de Graaf, Klaus von Gleissenthall, and Burcu Kulahcioglu Ozkan. Randomized testing of Byzantine fault tolerant algorithms. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):101:1–101:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586053.

Wirfs-Brock:2020:JFY

[WBE20] Allen Wirfs-Brock and Brendan Eich. JavaScript: the first 20 years. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):77:1–77:189, June

2020. URL https://dl.acm.
org/doi/abs/10.1145/3386327.

Wolff:2021:MSV

[WBM^{+}21] Fabian Wolff, Aurel Bílý, Christoph Matheja, Peter Müller, and Alexander J. Summers. Modular specification and verification of closures in Rust. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):145:1–145:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485522.

Wu:2022:SVG

[WBS^{+}22] Haoze Wu, Clark Barrett, Mahmood Sharif, Nina Narodytska, and Gagandeep Singh. Scalable verification of GNNbased job schedulers. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):162:1–162:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3563325.

Wei:2020:CSE

[WBTR20] Guannan Wei, Oliver Bracevac, Shangyin Tan, and Tiark Rompf. Compiling symbolic execution with staging and algebraic effects. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):164:1–164:33, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428232.

Wise:2020:GVR

[WBW^{+}20] Jenna Wise, Johannes Bader, Cameron Wong, Jonathan Aldrich, Éric Tanter, and Joshua Sunshine. Gradual verification of recursive heap data structures. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):228:1–228:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428296.

Wu:2017:HTE

[WC17] Baijun Wu and Sheng Chen. How type errors were fixed and what students did? Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):105:1–105:??, October 2017. CODEN ???? ISSN 2475-1421.

Wang:2018:SSR

[WCB18] Chenglong Wang, Alvin Cheung, and Rastislav Bodik. Speeding up symbolic reasoning for relational queries. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):157:1–157:25, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276527.

Wood:2017:IBD

[WCBG17] Benjamin P. Wood, Man Cao, Michael D. Bond, and Dan Grossman. Instrumentation bias for dynamic data race detection. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):69:1– 69:??, October 2017. CODEN ???? ISSN 2475-1421.

Wu:2017:LUF

[WCC17] Baijun Wu, John Peter Campora III, and Sheng Chen. Learning user friendly typeerror messages. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):106:1–106:??, October 2017. CODEN ???? ISSN 2475-1421.

Waye:2017:WHO

[WCD17] Lucas Waye, Stephen Chong, and Christos Dimoulas. Whip: higher-order contracts for modern services. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 36:1–36:??, September 2017. CODEN ???? ISSN 2475-1421.

Wand:2018:CEP

[WCGC18] Mitchell Wand, Ryan Culpepper, Theophilos Giannakopoulos, and Andrew Cobb. Contextual equivalence for a probabilistic language with continuous random variables and recursion. Proceedings of the ACM on Programming Languages (PACMPL), 2 (ICFP):87:1–87:30, July 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3236782.

Wu:2019:GPE

[WCH^{+}19] Baijun Wu, John Peter Campora III, Yi He, Alexander Schlecht, and Sheng Chen. Generating precise error specifications for C: a zero shot learning approach. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):160:1–160:30, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360586.

Winograd-Cort:2017:FAD

[WCHRP17] Daniel Winograd-Cort, Andreas Haeberlen, Aaron Roth, and Benjamin C. Pierce. A framework for adaptive differential privacy. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 10:1–10:??, September 2017. CODEN ???? ISSN 2475-1421.

Wang:2019:CGM

[WCMH19] Shengyi Wang, Qinxiang Cao, Anshuman Mohan, and Aquinas Hobor. Certifying graphmanipulating C programs via localizations within data structures. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):171:1– 171:30, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360597.

Wei:2019:SAI

[WCR19] Guannan Wei, Yuxuan Chen, and Tiark Rompf. Staged abstract interpreters: fast and modular whole-program analysis via meta-programming. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):126:1–126:32, October 2019. URL https://dl.

acm.org/doi/abs/10.1145/3360552.

Weirich:2019:RDT

[WCVE19] Stephanie Weirich, Pritam Choudhury, Antoine Voizard, and Richard A. Eisenberg. A role for dependent types in Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):101:1–101:29, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3341705.

Wang:2018:VED

[WDLC18] Yuepeng Wang, Isil Dillig, Shuvendu K. Lahiri, and William R. Cook. Verifying equivalence of database-driven applications. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 56:1–56:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Wei:2018:RAA

[WDR18] Guannan Wei, James Decker, and Tiark Rompf. Refunctionalization of abstract abstract machines: bridging the gap between abstract abstract machines and abstract definitional interpreters (functional pearl). Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):105:1–105:28, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236800.

Wang:2017:SDC

[WDS17] Xinyu Wang, Isil Dillig, and Rishabh Singh. Synthesis of data completion scripts using finite tree automata. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):62:1–62:??, October 2017. CODEN ???? ISSN 2475-1421.

Wang:2018:PSU

[WDS18] Xinyu Wang, Isil Dillig, and Rishabh Singh. Program synthesis using abstraction refinement. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL): 63:1–63:??, January 2018. CO-DEN ???? ISSN 2475-1421.

Wang:2020:VE

[WFB^{+}20] Chenglong Wang, Yu Feng, Rastislav Bodik, Alvin Cheung, and Isil Dillig. Visualization by example. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 49:1–49:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371117.

Wang:2020:PES

[WFC^{+}20] Peixin Wang, Hongfei Fu, Krishnendu Chatterjee, Yuxin Deng, and Ming Xu. Proving expected sensitivity of probabilistic programs with randomized variable-dependent termination time. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 25:1–25:30, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371093.

Watanabe:2021:CSH

[WGP^{+}21] Yasunari Watanabe, Kiran Gopinathan, George Pîrlea, Nadia Polikarpova, and Ilya Sergey. Certifying the synthesis of heap-manipulating programs. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 84:1–84:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473589.

Wang:2019:TGW

[WH19] Di Wang and Jan Hoffmann.

Type-guided worst-case input generation. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL): 13:1–13:30, January 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3290326.

Wang:2019:IEM

[WHZ^{+}19] Yu-Ping Wang, Xu-Qiang Hu, Zi-Xin Zou, Wende Tan, and Gang Tan. IVT: an efficient method for sharing subtype polymorphic objects. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):130:1–130:22, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360556.

Wagner:2023:EBV

[WJS23] Christopher Wagner, Nouraldin Jaber, and Roopsha Samanta. Enabling bounded verification of doubly-unbounded distributed agreement-based systems via bounded regions. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):81:1–81:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586033.

Wang:2020:REA

[WKH20] Di Wang, David M. Kahn, and Jan Hoffmann. Raising expectations: automating expected cost analysis with types. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):110:1– 110:31, August 2020. URL

https://dl.acm.org/doi/10.
1145/3408992.

Wang:2023:UCT

[WKN^{+}23] Bo Wang, Aashish Kolluri, Ivica Nikolić, Teodora Baluta, and Prateek Saxena. Usercustomizable transpilation of scripting languages. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):82:1–82:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586034.

Weisenburger:2018:DSD

[WKS18] Pascal Weisenburger, Mirko Köhler, and Guido Salvaneschi. Distributed system development with ScalaLoci. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):129:1–129:30, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276499.

Wang:2023:SCL

[WL23] Chenglin Wang and Fangzhen Lin. Solving conditional linear recurrences for program verification: The periodic case. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):76:1–76:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586028.

Weitz:2017:SLB

[WLH^{+}17] Konstantin Weitz, Steven Lyubomirsky, Stefan Heule, Emina Torlak, Michael D. Ernst, and Zachary Tatlock. SpaceSearch: a library for building and verifying solveraided tools. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 25:1–25:??, September 2017. CODEN ???? ISSN 2475-1421.

Wang:2023:CSL

[WLJ^{+}23] Yanjun Wang, Zixuan Li, Chuan Jiang, Xiaokang Qiu, and Sanjay Rao. Comparative synthesis: Learning nearoptimal network designs by query. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 4:1–4:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571197.

Wong:2018:FVE

[WMLK18] Chu-Pan Wong, Jens Meinicke, Lukas Lazarek, and Christian Kästner. Faster variational execution with transparent bytecode transformation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):117:1–117:30, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276487.

Weidner:2020:CDO

[WMM20] Matthew Weidner, Heather Miller, and Christopher Meiklejohn. Composing and decomposing op-based CRDTs with semidirect products. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):94:1–94:27, August

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3408976.

Williams:2018:RCB

[WMW18] Jack Williams, J. Garrett Morris, and Philip Wadler. The root cause of blame: contracts for intersection and union types. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):134:1–134:29, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276504.

Walia:2019:HLI

[WNC^{+}19] Rajan Walia, Praveen Narayanan, Jacques Carette, Sam Tobin-Hochstadt, and Chung chieh Shan. From high-level inference algorithms to efficient code. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):98:1–98:30, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341702.

Willsey:2021:EFE

[WNW^{+}21] Max Willsey, Chandrakana Nandi, Yisu Remy Wang, Oliver Flatt, Zachary Tatlock, and Pavel Panchekha. egg: Fast and extensible equality saturation. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL): 23:1–23:29, January 2021. URL

https://dl.acm.org/doi/10.
1145/3434304.

Williams:2018:PAO

[WR18] Thomas Williams and Didier Rémy. A principled approach to ornamentation in ML. Proceedings of the ACM on Programming Languages (PACMPL), 2(POPL):21:1–21:??, January 2018. CODEN ???? ISSN 2475-

Watt:2019:WW

[WRPP19] Conrad Watt, Andreas Rossberg, and Jean Pichon-Pharabod. Weakening WebAssembly. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):133:1–133:28, October 2019. URL https://dl.

                                        acm.org/doi/abs/10.1145/3360559.

Wu:2020:HII

[WSG^{+}20] Zhefeng Wu, Zhe Sun, Kai Gong, Lingyun Chen, Bin Liao, and Yihua Jin. Hidden inheritance: an inline caching design for TypeScript performance. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):174:1–174:29, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428242.

Wimmer:2019:IOS

[WSH^{+}19] Christian Wimmer, Codrut Stancu, Peter Hofer, Vojin Jovanovic, Paul Wögerer, Peter B. Kessler, Oleg Pliss, and Thomas Würthinger. Initialize once, start fast: application initialization at build time. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):184:1–184:29, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360610.

Weirich:2017:SDT

[WVdAE17] Stephanie Weirich, Antoine Voizard, Pedro Henrique Azevedo de Amorim, and Richard A.

Eisenberg. A specification for dependent types in Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 31:1–31:??, September 2017. CODEN ???? ISSN 2475-1421.

Wang:2017:TFL

[WWC17] Peng Wang, Di Wang, and Adam Chlipala. TiML: a functional language for practical complexity analysis with invariants. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):79:1– 79:??, October 2017. CODEN ???? ISSN 2475-1421.

Wang:2018:RPS

[WWD18] Yuepeng Wang, Xinyu Wang, and Isil Dillig. Relational program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):155:1–155:27, Octo-

ber 2018. URL https://dl.
acm.org/doi/abs/10.1145/3276525.

Wang:2020:LSP

[WWGW20] Yu Wang, Ke Wang, Fengjuan Gao, and Linzhang Wang. Learning semantic program embeddings with graph interval neural network. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):137:1–137:27, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428205.

Willis:2020:SSP

[WWP20] Jamie Willis, Nicolas Wu, and Matthew Pickering. Staged selective parser combinators. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP):120:1–120:30, August

                                        2020. URL https://dl.acm.
                                        org/doi/10.1145/3409002.

Wang:2020:CVS

[WXWS20] Yuting Wang, Xiangzhe Xu, Pierre Wilke, and Zhong Shao. CompCertELF: verified separate compilation of C programs into ELF object files. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):197:1–197:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428265.

Westrick:2020:DNP

[WYFA20] Sam Westrick, Rohan Yadav, Matthew Fluet, and Umut A. Acar. Disentanglement in nested-parallel programs. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 47:1–47:32, January 2020.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3371115.

Wang:2022:CGC

[WYT^{+}22] Chengpeng Wang, Peisen Yao, Wensheng Tang, Qingkai Shi, and Charles Zhang. Complexityguided container replacement synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):

68:1–68:31, April 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527312.

Wang:2019:DDP

[WZD^{+}19] Fei Wang, Daniel Zheng, James Decker, Xilun Wu, Grégory M. Essertel, and Tiark Rompf. Demystifying differentiable programming: shift/reset the penultimate backpropagator. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):96:1–96:31, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341700.

Wang:2019:DNP

[WZS19] Shuai Wang, Chengyu Zhang, and Zhendong Su. Detecting nondeterministic payment bugs in Ethereum smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA):189:1– 189:29, October 2019. URL

https://dl.acm.org/doi/abs/
10.1145/3360615.

Winterer:2020:UET

[WZS20] Dominik Winterer, Chengyu Zhang, and Zhendong Su. On the unusual effectiveness of type-aware operator mutations for testing SMT solvers. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):193:1–193:25, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428261.

Wang:2022:VCC

[WZSK22] Yuting Wang, Ling Zhang, Zhong Shao, and Jérémie Koenig. Verified compilation of C programs with a nominal memory model. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):25:1–25:31, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498686.

Xie:2020:EHE

[XBH^{+}20] Ningning Xie, Jonathan Immanuel Brachthäuser, Daniel Hillerström, Philipp Schuster, and Daan Leijen. Effect handlers, evidently. Proceedings of the ACM on Programming Languages (PACMPL), 4(ICFP): 99:1–99:29, August 2020. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3408981.

Xu:2018:SJC

[XCB^{+}18] Xiaoran Xu, Keith Cooper, Jacob Brock, Yan Zhang, and Handong Ye. Share-JIT: JIT code cache sharing across processes and its practical implementation. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):124:1–124:23, Octo-

                                        ber 2018. URL https://dl.
                                        acm.org/doi/abs/10.1145/3276494.

Xie:2022:FCN

[XCIL22] Ningning Xie, Youyou Cong, Kazuki Ikemori, and Daan Leijen. First-class names for effect handlers. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):126:1–126:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563289.

Xie:2020:KID

[XEdSO20] Ningning Xie, Richard A.

Eisenberg, and Bruno C. d. S. Oliveira. Kind inference for datatypes. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 53:1–53:28, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371121.

Xu:2023:MTD

[XHdSO23] Han Xu, Xuejing Huang, and Bruno C. d. S. Oliveira. Making a type difference: Subtraction on intersection types as generalized record operations. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):31:1–31:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571224.

Xu:2021:CPC

[XK21] Haoran Xu and Fredrik Kjolstad. Copy-and-patch compilation: a fast compilation algorithm for high-level languages and bytecode. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):136:1–136:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485513.

Xie:2021:GEP

[XL21] Ningning Xie and Daan Leijen. Generalized evidence passing for effect handlers: efficient compilation of effect handlers to C. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 71:1–71:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473576.

Xiang:2020:PIE

[XLD20] Tongtong Xiang, Jeff Y. Luo, and Werner Dietl. Precise inference of expressive units of measurement types. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):142:1–142:28, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428210.

Xie:2022:SCS

[XPL^{+}22] Ningning Xie, Matthew Pickering, Andres Löh, Nicolas Wu, Jeremy Yallop, and Meng Wang. Staging with class: a specification for typed template Haskell. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 61:1–61:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498723.

Xu:2023:PLM

[XTZ^{+}23] Zhenyang Xu, Yongqiang Tian, Mengxiao Zhang, Gaosen Zhao, Yu Jiang, and Chengnian Sun. Pushing the limit of 1minimality of language-agnostic program reduction. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):97:1–97:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586049.

Yang:2021:SDR

[YAC21] Cambridge Yang, Eric Atkinson, and Michael Carbin. Simplifying dependent reductions in the polyhedral model. Proceedings of the ACM on Programming Languages (PACMPL), 5(POPL):20:1–20:33, January

2021. URL https://dl.acm.
org/doi/10.1145/3434301.

Yallop:2017:SGP

[Yal17] Jeremy Yallop. Staged generic programming. Proceedings of the ACM on Programming Languages (PACMPL), 1(ICFP): 29:1–29:??, September 2017. CODEN ???? ISSN 2475-1421.

Yefet:2020:AEM

[YAY20] Noam Yefet, Uri Alon, and Eran Yahav. Adversarial examples for models of code. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):162:1–162:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428230.

Yuan:2022:TDS

[YC22] Charles Yuan and Michael Carbin. Tower: data structures in quantum superposition. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):134:1– 134:??, October 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3563297.

Ye:2022:OAD

[YD22] Qianchuan Ye and Benjamin Delaware. Oblivious algebraic data types. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):51:1–51:29, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3498713.

Yanovski:2021:PGS

[YDJD21] Joshua Yanovski, Hoang-Hai Dang, Ralf Jung, and Derek Dreyer. GhostCell: separating permissions from data in Rust. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):

92:1–92:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473597.

Yang:2017:UTS

[YdSO17] Yanpeng Yang and Bruno C.

d. S. Oliveira. Unifying typing and subtyping. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):47:1–47:??, October 2017. CODEN ???? ISSN 2475-1421.

Yorgey:2018:WDF

[YF18] Brent A. Yorgey and Kenneth Foner. What’s the difference? A functional pearl on subtracting bijections. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):101:1–101:21, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236796.

Yuan:2023:LPM

[YGG^{+}23] Yongwei Yuan, Scott Guest, Eric Griffis, Hannah Potter, David Moon, and Cyrus Omar. Live pattern matching with typed holes. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):96:1–96:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586048.

Yu:2020:IPA

[YHW20] Qianshan Yu, Fei He, and Bow-Yaw Wang. Incremental predicate analysis for regression verification. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):184:1–184:25, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428252.

Yan:2022:ILQ

[YJY22] Peng Yan, Hanru Jiang, and Nengkun Yu. On incorrectness logic for quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA1):72:1–72:28, April 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3527316.

Katsumata:2022:FPG

[yKMUW22] Shin ya Katsumata, Dylan McDermott, Tarmo Uustalu, and Nicolas Wu. Flexible presentations of graded monads. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):123:1– 123:??, August 2022. CODEN ???? ISSN 2475-1421 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3547654.

Yan:2021:PSS

[YM21] Pengbo Yan and Toby Murray. SecRSL: security separation logic for C11 releaseacquire concurrency. Proceedings of the ACM on Programming Languages (PACMPL), 5(OOPSLA):99:1–99:26, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485476.

Yuan:2022:TSR

[YMC22] Charles Yuan, Christopher Mc-Nally, and Michael Carbin. Twist: sound reasoning for purity and entanglement in quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):30:1–30:32, January 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3498691.

Yamaguchi:2021:PSS

[YMDW21] Masaomi Yamaguchi, Kazutaka Matsuda, Cristina David, and Meng Wang. Synbit: synthesizing bidirectional programs using unidirectional sketches. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):105:1–105:31, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485482.

Yuan:2017:ASC

[YMJ17] Binhang Yuan, Vijayaraghavan Murali, and Christopher Jermaine. Abridging source code. Proceedings of the ACM on Programming Languages (PACMPL), 1(OOPSLA):58:1– 58:??, October 2017. CODEN ???? ISSN 2475-1421.

Yamazaki:2019:GFA

[YNIC19] Tetsuro Yamazaki, Tomoki Nakamaru, Kazuhiro Ichikawa, and Shigeru Chiba. Generating a fluent API with syntax checking from an LR grammar. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):134:1–134:24, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360560.

Yao:2021:PAE

[YSHZ21] Peisen Yao, Qingkai Shi, Heqing Huang, and Charles Zhang. Program analysis via efficient symbolic abstraction. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):118:1–118:32, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485495.

Ye:2023:GPL

[YTO23] Wenjia Ye, Matías Toro, and Federico Olmedo. A gradual probabilistic Lambda calculus. Proceedings of the ACM on Programming Languages (PACMPL), 7(OOPSLA1):84:1– 84:??, April 2023. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3586036.

Yallop:2018:PSD

[YvGK18] Jeremy Yallop, Tamara von Glehn, and Ohad Kammar. Partially-static data as free extension of algebras. Proceedings of the ACM on Programming Languages (PACMPL), 2(ICFP):100:1–100:30, July

2018. URL https://dl.acm.
org/doi/abs/10.1145/3236795.

Yallop:2019:LUS

[YW19] Jeremy Yallop and Leo White.

Lambda: the ultimate sublanguage (experience report). Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):116:1–116:17, July

2019. URL https://dl.acm.
org/doi/abs/10.1145/3342713.

Yang:2021:RAE

[YW21] Zhixuan Yang and Nicolas Wu.

Reasoning about effect interaction by fusion. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 73:1–73:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473578.

Yaghmazadeh:2017:SQS

[YWDD17] Navid Yaghmazadeh, Yuepeng Wang, Isil Dillig, and Thomas Dillig. SQLizer: query synthesis from natural language. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):63:1–63:??, October 2017. CODEN ???? ISSN 2475-1421.

Xia:2020:ITR

[yXZH^{+}20] Li yao Xia, Yannick Zakowski, Paul He, Chung-Kil Hur, Gregory Malecha, Benjamin C.

Pierce, and Steve Zdancewic. Interaction trees: representing recursive and impure programs in Coq. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL): 51:1–51:32, January 2020.

URL https://dl.acm.org/
doi/abs/10.1145/3371119.

Yoon:2022:FRA

[YZZ22] Irene Yoon, Yannick Zakowski, and Steve Zdancewic. Formal reasoning about layered monadic interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 6(ICFP):99:1–99:??, August 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547630.

Zhang:2022:RAR

[ZA22] Yizhou Zhang and Nada Amin. Reasoning about “reasoning about reasoning”: semantics and contextual equivalence for probabilistic programs with nested queries and recursion. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 16:1–16:28, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498677.

Zhang:2022:OLP

[ZBG^{+}22] Yuhao Zhang, Yasharth Bajpai, Priyanshu Gupta, Ameya Ketkar, Miltiadis Allamanis, Titus Barik, Sumit Gulwani, Arjun Radhakrishna, Mohammad Raza, Gustavo Soares, and Ashish Tiwari. Overwatch: learning patterns in code edit sequences. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):139:1–139:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563302.

Zhao:2022:SEE

[ZBS^{+}22] Zihan Zhao, Sidi Mohamed Beillahi, Ryan Song, Yuxi Cai, Andreas Veneris, and Fan Long. SigVM: enabling event-driven execution for truly decentralized smart contracts. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):149:1–149:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563312.

Zhou:2023:CFV

[ZBS^{+}23] Li Zhou, Gilles Barthe, Pierre-Yves Strub, Junyi Liu, and Mingsheng Ying. CoqQ: Foundational verification of quantum programs. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL): 29:1–29:??, January 2023. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3571222.

Zakowski:2021:MCE

[ZBY^{+}21] Yannick Zakowski, Calvin Beck, Irene Yoon, Ilia Zaichuk, Vadim Zaliva, and Steve Zdancewic. Modular, compositional, and executable formal semantics for LLVM IR. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 67:1–67:30, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473572.

Zhou:2023:FPT

[ZCH23] Jie Zhou, John Criswell, and Michael Hicks. Fat pointers for temporal memory safety of C. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):86:1–86:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3586038.

Zhang:2022:ILK

[ZdAG22] Cheng Zhang, Arthur Azevedo de Amorim, and Marco Gaboardi. On incorrectness logic and Kleene algebra with top and tests. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 29:1–29:30, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498690.

Zhou:2021:DDA

[ZDDJ21] Zhe Zhou, Robert Dickerson, Benjamin Delaware, and Suresh Jagannathan. Datadriven abductive inference of library specifications. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):116:1–116:29, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485493.

Zilberstein:2023:OLU

[ZDS23] Noam Zilberstein, Derek Dreyer, and Alexandra Silva. Outcome logic: a unifying foundation for correctness and incorrectness reasoning. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):93:1–93:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586045.

Zhao:2019:MFH

[ZdSOS19] Jinxu Zhao, Bruno C. d. S. Oliveira, and Tom Schrijvers. A mechanical formalization of higherranked polymorphic type inference. Proceedings of the ACM on Programming Languages (PACMPL), 3(ICFP):112:1– 112:29, July 2019. URL https:

//dl.acm.org/doi/abs/10.1145/
3341716.

Zhou:2020:RIR

[ZdSOZ20] Yaoda Zhou, Bruno C. d. S. Oliveira, and Jinxu Zhao. Revisiting isorecursive subtyping. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):223:1–223:28, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428291.

Zhou:2020:SVR

[ZFH^{+}20] Fangyi Zhou, Francisco Ferreira, Raymond Hu, Rumyana Neykova, and Nobuko Yoshida. Statically verified refinements for multiparty protocols. Proceedings of the ACM on Programming Languages (PACMPL), 4(OOPSLA):148:1–148:30, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428216.

Zuriel:2019:ELF

[ZFS^{+}19] Yoav Zuriel, Michal Friedman, Gali Sheffi, Nachshon Cohen, and Erez Petrank. Efficient lock-free durable sets. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):128:1–128:26, Octo-

                                        ber 2019. URL https://dl.
                                        acm.org/doi/abs/10.1145/3360554.

Zhang:2023:BOO

[ZGHH23] Xing Zhang, Guanchen Guo, Xiao He, and Zhenjiang Hu. Bidirectional object-oriented programming: Towards programmatic and direct manipulation of objects. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):83:1–83:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3586035.

Zou:2022:OFR

[ZGS^{+}22] Daming Zou, Yuchen Gu, Yuanfeng Shi, MingZhe Wang, Yingfei Xiong, and Zhendong Su. Oracle-free repair synthesis for floatingpoint programs. Proceedings of the ACM on Programming Languages (PACMPL), 6 (OOPSLA2):159:1–159:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563322.

Zhang:2017:EIR

[ZGSN17] Xin Zhang, Radu Grigore, Xujie Si, and Mayur Naik. Effective interactive resolution of static analysis alarms. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):57:1–57:??, October 2017. CODEN ???? ISSN 2475-1421.

Zhu:2018:CRS

[ZH18] Fengmin Zhu and Fei He. Conflict resolution for structured merge via version space algebra. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):166:1– 166:25, October 2018. URL

https://dl.acm.org/doi/abs/
10.1145/3276536.

Zhang:2020:DPS

[ZHL^{+}20] Hailong Zhang, Yu Hao, Sufian Latif, Raef Bassily, and Atanas Rountev. Differentially-private software frequency profiling under linear constraints. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):203:1–203:24, November 2020. URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428271.

Zhang:2022:QSP

[ZK22] Linpeng Zhang and Benjamin Lucien Kaminski. Quantitative strongest post: a calculus for reasoning about the flow of quantitative information. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1):87:1– 87:29, April 2022. CODEN ???? ISSN 2475-1421 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3527331.

Zhang:2017:FUI

[ZM17] Yizhou Zhang and Andrew C. Myers. Familia: unifying interfaces, type classes, and family polymorphism. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):70:1–70:??, October 2017. CODEN ???? ISSN 2475-1421.

Zhang:2019:ASE

[ZM19] Yizhou Zhang and Andrew C. Myers. Abstraction-safe effect handlers via tunneling. Proceedings of the ACM on Programming Languages (PACMPL), 3(POPL):5:1–5:29, January

                                        2019. URL https://dl.acm.
                                        org/doi/abs/10.1145/3290318.

Zhang:2023:ATF

[ZMSD23] Guoqiang Zhang, Benjamin Mariano, Xipeng Shen, and Işıl Dillig. Automated translation of functional big data queries to SQL. Proceedings of the ACM on Programming Languages (PACMPL), 7 (OOPSLA1):95:1–95:??, April 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3586047.

Zyuzin:2021:CMT

[ZN21] Nikita Zyuzin and Aleksandar Nanevski. Contextual modal types for algebraic effects and handlers. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP): 75:1–75:29, August 2021. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473580.

Zhai:2017:ALP

[ZPG^{+}17] Ennan Zhai, Ruzica Piskac, Ronghui Gu, Xun Lao, and Xi Wang. An auditing language for preventing correlated failures in the cloud. Proceedings of the ACM on Programming Languages (PACMPL), 1 (OOPSLA):97:1–97:??, October 2017. CODEN ???? ISSN 2475-1421.

Zhang:2021:SDS

[ZPZX21] Jialu Zhang, Ruzica Piskac, Ennan Zhai, and Tianyin Xu. Static detection of silent misconfigurations with deep interaction analysis. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):140:1–140:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485517.

Zhou:2021:INM

[ZR21] Ziqiao Zhou and Michael K. Reiter. Interpretable noninterference measurement and its application to processor designs. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):141:1–141:30, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3485518.

Zhang:2019:FTL

[ZRH^{+}19] Hengchu Zhang, Edo Roth, Andreas Haeberlen, Benjamin C. Pierce, and Aaron Roth. Fuzzi: a three-level logic for differential privacy. Proceedings of the ACM on Programming Languages (PACMPL), 3 (ICFP):93:1–93:28, July 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3341697.

Zhang:2020:TDP

[ZRH^{+}20] Hengchu Zhang, Edo Roth, Andreas Haeberlen, Benjamin C. Pierce, and Aaron Roth. Testing differential privacy with dual interpreters. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):165:1–165:26, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428233.

Zhu:2022:BFA

[ZSL^{+}22] Fengmin Zhu, Michael Sammler, Rodolphe Lepigre, Derek Dreyer, and Deepak Garg. BFF: foundational and automated verification of bitfieldmanipulating programs. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2):182:1–182:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563345.

Zhang:2020:HBC

[ZSM20] Yizhou Zhang, Guido Salvaneschi, and Andrew C. Myers. Handling bidirectional control flow. Proceedings of the ACM on Programming Languages (PACMPL), 4 (OOPSLA):139:1–139:30, November 2020. URL https:

//dl.acm.org/doi/10.1145/
3428207.

Zwaan:2022:ITC

[ZvAV22] Aron Zwaan, Hendrik van Antwerpen, and Eelco Visser. Incremental type-checking for free: using scope graphs to derive incremental typecheckers. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA2): 140:1–140:??, October 2022. CODEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3563303.

Zhang:2022:RM

[ZWWT22] Yihong Zhang, Yisu Remy Wang, Max Willsey, and Zachary Tatlock. Relational ematching. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL): 35:1–35:22, January 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3498696.

Zhang:2021:UST

[ZXSD21] Guoqiang Zhang, Yuanchao Xu, Xipeng Shen, and Isil Dillig. UDF to SQL translation through compositional lazy inductive synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 5 (OOPSLA):112:1–112:26, October 2021. CODEN ???? ISSN 2475-1421 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3485489.

Zhang:2018:GHP

[ZYB^{+}18] Yunming Zhang, Mengjiao Yang, Riyadh Baghdadi, Shoaib Kamil, Julian Shun, and Saman Amarasinghe. GraphIt: a highperformance graph DSL. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):121:1–121:30, October 2018. URL https://dl.

                                        acm.org/doi/abs/10.1145/3276491.

Zhang:2019:BPD

[ZYT^{+}19] Zhuo Zhang, Wei You, Guanhong Tao, Guannan Wei, Yonghwi Kwon, and Xiangyu Zhang. BDA: practical dependence analysis for binary executables by unbiased whole-program path sampling and per-path abstract interpretation. Proceedings of the ACM on Programming Languages (PACMPL), 3 (OOPSLA):137:1–137:31, Octo-

ber 2019. URL https://dl.
acm.org/doi/abs/10.1145/3360563.

Zhou:2023:RSA

[ZZdSO23] Litao Zhou, Yaoda Zhou, and Bruno C. d. S. Oliveira. Recursive subtyping for all. Proceedings of the ACM on Programming Languages (PACMPL), 7(POPL):48:1–48:??, January 2023. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571241.

Zou:2020:DFP

[ZZX^{+}20] Daming Zou, Muhan Zeng, Yingfei Xiong, Zhoulai Fu, Lu Zhang, and Zhendong Su. Detecting floating-point errors via atomic conditions. Proceedings of the ACM on Programming Languages (PACMPL), 4(POPL):60:1–60:27, January

2020. URL https://dl.acm.
org/doi/abs/10.1145/3371128.